A Bibliography of Numerical Analysis Publications, 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

30 November 2023 Version 1.40

Title word cross-reference

2 [MW96]. Q [HM96, Met96]. QR [Pai90, Wat96]. R [Her96, HM96, Met96]. T [Hat97].

-adic [MW96]. -experiments [Hat97]. -like [Wat96]. -order [Her96, HM96, Met96].

1 [Lip94, VK96]. 12th [LMH91]. 13th [GW90]. 14th [GW92]. 15th [GW94b]. 17th [GHW98]. 18 [SD95]. 18th [Ain95]. 1988 [HMRS90]. 1989/90 [PS90]. 1991 [HMRS92]. 1994 [Ano94b].

2 [KV97, LT94, Tho99]. 21 [SD95]. 23rd [LO90]. 27th [BB92, LO90]. 29th [Ain95].

2nd [MK95].

3rd [MK98].

4 [B^{+}98].

5 [BB99, HL98, PES99, RC98, Zho94].

75th [MGW96]. 77

[Bor91, NL95, PTV96, WNO94].

89 [KvSR90].

90 [Du 97, Ort94a, PTM96, PTV96]. ’92 [HZO92]. ’95 [SD95]. ’96 [Des96, VWY97]. ’97 [Nog97]. ’98 [ALM99].

= [Ald92].

A.R. [MGW96]. absorption [FS90]. abstract [Lac95]. academic [Gir92]. acceptance [Coo90]. accounts [NMW96]. Accuracy [Hig96, Jam94a]. accurate [SS96a]. ACM [Ric90]. Acquisition [HN90]. Activity [MDE90, Sme96]. Adaptive [MDE90, Bab95, Cal96, GN97]. Adaptivity [SH97]. adic [MW96]. Advanced [Cra96, GW98, KU94, Ste98a]. Advances [DSV94, GH94, Lig91]. advection [Erb94, VK93]. advective [Elt90]. advective-diffusive [Elt90]. Advisor [BM90]. Advisory [BM90, Boi90]. aerodynamic [Lac95]. aeronautical [Ano92a]. affine [VY96]. after [Fox90b, Ste96]. after-dinner [Fox90b]. Afternotes [Ste96, Ste98a]. against [Cou94]. age [Blo90]. AGNES [DPR90]. agricultural [CS95]. AI [HD90]. Aiken [Coh90]. aircraft [Cha98]. airfoils [Rei92]. Airobic [DDLW90]. Alamos [Met90]. Algebra [Dem97, DDSvdV98, Duf97, Gen98, Hig97c, TB97, BCS91, BV97, Dat95, DH90, GMW91, Hig97b, Kon93, Kub90, Nas90a]. Algebraic [BCSS96, CM96, Sha96, Baj96, DGY96, ER96, Olv90, Roy96, SW96]. algebry [Kon93]. Algorithm [HHRS90, LG90, NA90, Par90b, BEM96, Cal96, Erh96, Fra96, LeB93]. Algorithms [BP94b, EMU96b, GHN^{+}90, GGM90, Gro90, HHKR93, Hig96, Mat92a, Nat97, Sed92, Bar96b, EMU96a, FP92, Hei96, Hig97a, Kul96, MW86, SS96b, Ste98b, Vas96, Wat96, vS99]. Alive [vG97]. alluvial [Hib93]. alternating [Mat92a]. alternative [ZAB95]. alternativnye [ZAB95]. ambient [Cha93a]. American [Eme90, Nog97, SM90, Sar92]. AMS [AGM93, RSS96]. AMS-SIAM [AGM93, RSS96]. Anaheim [Sar92]. Analisis [RRLA90]. analiza [LG95]. Analyse [LT94, Sch91a, RP98]. Analysis [Ain95, ALM99, ACD^{+}93, BC93, BC94, BC95, Bak97, Bha97, GC95, GB97, Got95, HT91, HD90, Hig90, HJ91, IK94, Jam94a, Jam94b, KC93, Lee97, Ric92, SH97, Tho90, Tod62, Uni98c, Uni99, Ueb97, Ago94, ADK91, Aki94, AI98, AGM93, Ano94b, Ano96, Arm97, AA97, AS93b, Asa95, Ath95, Atk93, AP90, Avr93, Bäc99, Bau98, BB92, Bel97, Bet98, BR92, BL98, Blo90, Bon93, Bre92, Bri95, BT92, Bul94, Bul96, BF93, BF97, Car95, Cha93a, Cha98, CCL99, CL90, Cou94, Ded96, DH95b, DH95a, DFS97, Don96, DW97, EZ91, EWK90a, EWK90b, Eva95, FS90, Fau99, FM92, Fox90a, FK94, GZ96, GW99a, Gau97a, Gau97b, GS90, GW94a, GW99b, Gir92, Gok93, GH94, GW90, GW92, GW94b, GW96, GHW98, GhW90, Guc95, Hag90]. analysis [HP90, HR99, Har99, Her94, Hib93, Ho92, Ho94, Hon93, Hon94, HJ90, Ise96, IZ96, KG91, Ket99, KC91a, KC91b, KC96, KW94, Kre98, Lac95, Lam93, Lan99a, LT94, LeB93, Tal90, Lee94, Lig91, Lip94, LG95, LL97, LSR90, ML90, MAA97, McV90, Mei96, Mik91, MGW96, MS95, Nak96, NN96, Naw94, NMW96, Nog97, Oki96, Ort90a, Ort90b, Osh94, Par91, Par90a, Pax92, PW92, Pax95, PT96, Ply92, PS90, PS91a, PS91b, RP98, Ras94, RSS96, Ric93, RD93, Ric96, Rig96, Roc97, SM90, SJ96, Sch91a, Sch90, SCI^{+}96, SM91, Ste96, Ste98a, SB93, SB96a, SH96, Tao94, Tes93, Tur90, Tyr97, VJ95, Vij95, VWY97, Wak90, Wat92, WG96, WSG91, WZ94, Xu99, YCC96a, YCC96b, bZ93a, Zho94, Zho96, ALM99]. analytic [Ste93]. Analytical [GW98, HSDM91, Hud91, Mod98, Uni98a, Guc95, Mer95, Mer97]. analyzer [Bun90]. Anatomy [DPR90]. and/or [FMO96a]. André [LRRL97]. Anfänger [Opf93]. Angeles [Ju95]. angular [Cox90]. anisotropic [Arm97]. Annual [Eme90, SM90, Sar92]. Any [KMRT90]. Aplicadas [Ald92]. Application [ADK91, CB90, Civ98, Cer93, Fer98]. Applications [Gen98, HHRS90, MN99, PLC94, dS97, AS93a, BCF95, BCF97, Bre92, CM96, CG92, Cha93b, Cha92, Cha95, CC98, CM99, Dat95, DSV94, EG91, EG92, GW99a, Hag90, Mar94, Mar92, MK95, MK98, MS95, Mu97, PC97, PH92, PT96, Rei92, Sha96, SD95, TW98a, Van96, VWY97, ZAB95]. Applications/Architecture [HHRS90]. Applied [Aka94, Ald92, CS95, Dem97, Fau99, GW94a, GW99b, JSW93, Ju95, Nak91, Nak93, RSS96, AI98, AGM93, Ano92b, Ano94b, Des96, HPO92, LT94, MO95, RF96, Ply92, Vin91]. appliquée [LT94]. appreciations [PM90]. Approach [HD90, BZ93b, CM96, CSV96, Dor96, GS91a, GS91b, ML90, Mun97, NN96, Nov96, Sme96, Tho90, TW98b, Van97]. approaches [Rei97, ZAB95]. Approximate [Ded96, Deĭ94, AIL92]. approximating [Car96, Her96]. Approximation [Ell97, GGM90, KSS95, Mor97b, Stu97, Wat97, Adž96, BS92a, BV97, BCFH97, Cuy94, Hro93, QV94, QV97, Ras94, Sam96]. Approximations [ER96, Dab98]. April [LO90, SO97]. aquifer [Hib93]. aquifers [Mod98]. arbitrary [Naw94]. Arcachon [Bru98]. Architecture [HHRS90]. Area [Oki96, LSR90]. areas [LP90]. arguments [Sny97]. arising [HM96]. Arithmetic [Ull90b, Bar96b, LT96, Olv90, Ull90a]. Armed [Mar92, MK95, MK98]. Art [PFTV92, PTV^{+}92, PTM96, PTV96, DW97, LT94, Pre94]. artesian [Mer95, Mer97]. Artificial [DDLW90, Jam94a, Jam94b]. ASME [Cro94, Ju95, SJ96, VS94]. Aspects [Wat97, BV90b, Gol99, Pai90]. assembly [Cha93a]. Assessment [Ano97, Boi97]. Assisting [HD90]. assumed [RKA93]. assumed-stress [RKA93]. assured [WSG91]. Asymptotic [Hag90, KG91, KG93, Rig96]. asymptotically [Cal96]. Asynchronous [Pet96]. Atlanta [SJ96]. atmospheric [Erb94, LP90, LRRL97]. August [AGM93, BC93, BC94, BC95, DSV94, HMRS92, RSS96]. Austin [LL97].

Automated [GL90]. Automatic

[AK93, BGGD90, Bis97, DPR90, DFS97, MOF90, TPD90, Kul96]. aux [Euv94]. Available [Boi90]. average [RW96, Roj96a]. axial [Fri97].

B [JB94]. back [Hib93]. Balancing [CZ90]. Bangalore [DDN95]. banks [Str96]. Base [Uni98c]. Based [AG90b, FHI90, KMRT90, MDE90, STL96, AAK96, Cou97, Dob96, Don96, Har99, KSH91, Lyn97, MW86, Sch96, Shi95, Ste93, Woz96]. Basel [Ull90b]. Basic [HHKR93, Ste98b, Tod78, Lae99, PH92, RS92b, RS92a, Vit94]. Bayesian [RF96, Dor97]. Baysian [Nov96]. beam [Cou94, Fri97]. bearings [Vij95]. Beginners [Opf93]. Beginning [LG95]. behavior [Tao94]. behaviour [Bon93, Tes93]. Beispielen [KKN^{+}91b]. Belgium [HPO92, HZO92]. Benard [EZ91]. best [Hro93]. between [GS90, VY96]. Beyond [Ise97]. biasing [Jam94a]. Biennial [GHW98]. Bifurcation [CCL99, Rei91]. bijin [WY90]. Biorthogonality [Bre92]. Biplots [GH96]. Birkhoff [Lor92]. Birmingham [BB92]. birthday [MGW96]. BIT [Frö90]. bituminous [Bon93]. BKK [LWW96]. Block [STL96, Cox90]. block-angular [Cox90]. blowing [Ago94]. body [Cha92]. Boltzmann [Elt90]. Book [Gra90, Her98, RC96]. bound [LWW96]. Boundary [Atk97, Gol99, Kel93, PC97, Ram94, AA96, Fox90c, Hac92, Hal94, IY94, Kyt95, Li90, Nei91]. Boundary-Value [Kel93]. bounded [Wer96]. bounding [Hol97, LT96, Rie96]. Bounds [Sin96, AH96, Die96]. Brent [PS96b]. bridge [MAA97]. bridges [Tes93]. brief [Tyr97]. Brownian [Cal96]. browsing [BHR96]. Brussels [HPO92, HZO92]. buckling [NN96]. building [Du 97]. Bulgaria [BC93, BC94, BC95, DSV94, VWY97]. Bulgarian [KV97, VK96]. buoyancy [GB97]. Bureau [Tod90]. burn [YCC96a].

C [Abe98, Dub97, EMU96a, Flo95, Gla93, KKL^{+}93, Mor97a, Nak93, OG99, PFTV92, Roj96b, Sed92, Zac96, Zac98]. C-XSC [KKL^{+}93]. CA [KFC97]. cable [MAA97]. Calcolo [FT92]. calculation [LS97, Met96, Sie94, Ume97, TJL92]. Calculations [Act96, CGT97, FT92, KV97, Pru93, VK96]. Calculus [MN99, Bul93]. calibration [WZ94]. California [AS93b, BCF95, BCF97, Ju95, Sar92]. calorically [FMO96a]. Cambridge [Whe90]. Canada [Cel90]. Canadian [Cel90]. canonical [BV90b]. Capabilities [Uni99, Cha98]. capacity [McV90]. capricious [Wal96b]. Capturing [BFHR97, Osh94]. carbon [Rig96]. Carlo [Sob91, Hei96, Kel96, Man99, Nak97, Sob91, Sob94]. Case [DHP97, Wat97, RW96]. catastrophic [NMW96]. cavitating [Vij95]. cavities [Ath95]. cavity [Ath95]. Center [Gee90]. centered [Pan97a]. central [Mer95, Mer97]. certain [Roj96a]. cEVM [Pru93]. channels [Ses98]. Chaos [Rei91]. chaotic [NL96]. characteristic [Sam96]. characteristics [Vin91]. Chast [KV97, VK96]. chemical [CM99, CS99, FMO96a]. child [Frö90]. Chile [Ald92]. China [yKaY92, GhW90]. chined [Ago94]. Chinese [TJL92, WY90]. chislennogo [Bur93, LG95]. Chislennye [BV90a, Nur91, Tik90, KLP90]. Choice [GGM90]. Cholesky [Hig90, SS96b]. Cincias [Ald92]. Circuits [Got95]. circular [Rei92]. City [RSS96]. Civil [Nog97]. CLAM [Gro90]. class [Her96, KKL^{+}93]. classes [BCFH97]. Classical [Ise97, Rie96]. clays [Hol97]. closely [Lam93]. Closer [Gv97]. Clustered [Uni98c]. Coastal [Mod98]. codes [Nak97]. Collected [ACD^{+}93]. Collection [WNO94]. collections [BPRB97]. College [Ste96]. collisions [Lam93]. Colloquium [BC93, BC94, BC95, SO97]. Colorado [AGM93, Hib93]. combinations [Li90]. combinatorics [Baj96]. Combined [AAK96]. combustion [YCC96b]. combustor [YCC96a, YCC96b]. comments [Ode90]. Compact [Nas90a]. Companion [GRM97]. Comparison [NL96, RKA93, Uni98b, LM91]. Compiler [AG90b]. Compiler-Based [AG90b]. complement [GMSW90]. Complex [Ton90, Baj96, Cha92, PTŽ96, SS96b, Ver92]. Complexity [Hei96, CL96, DKS96, ER96, GM96, KLRK98, MW96, Pla96, RW96, Vas96, Wer96, Woz96]. composite [Bul96, Ho92, Ho94, SJ96]. compound [Die96]. comprehensive [HL96, HL98]. Compressible [Uni98b, FMO96b, Lee94]. Computability [Sie96]. Computation [Cra94, Cra96, Gla93, HL95, Her95, JSW93, Jen77, KKN^{+}91a, KSS95, The97, Mer95, Mer97, NA90, Rei91, Ueb97, Var90d, Cou97, CH90, DH95a, Elt90, Fiu95, GW99a, Isa90, JM92, Poz98, SCI^{+}96, Tod90]. Computational [BCF97, BV90b, Des96, GW98, HPO92, Ise97, KLRK98, LM97, LP97, Lan99b, MW96, SS97b, Bab90, Bet98, CC97, Dor96, HSDM91, Hro93, Kon93, Pan97b, Pla96, Rya94, TW98b, Zac96, Zac98, KLP90]. Computations [BP94b, CCF96, GHN^{+}90, GV96b, Roy96, Wat91, Bir90, Eat97, KLRK98, Pan96]. Computer [HS93, KPS94, Sal99, SC96, TW98a, Ull90a, Ull90b, Ano94b, Blo90, BJ92, Coh90, Don95, FK94, Frö90, Har99, JB94, Kel96, KV97, KU94, LT96, Pru93, VK96]. computer-based [Har99]. Computers [DDSv91, DDSvdV98, TPD90, LP97, Nas90a]. Computing [Act96, AK93, CK94, Cor95, DGY96, FJSD96a, FJSD96b, GH93, GH95, GH97, Gee90, GO93, GO96b, GO92a, HH90, HHKR93, Hea97, HRV90, KC93, LP98, MR96, MM96, Nas90b, Ort94b, Ort94a, PWF90, PFTV92, PTV^{+}92, PTM96, PTV96, SK93, SK95, Bjö90, BEM96, Dub97, ER96, FM92, GO92b, GO95, GO96a, G^{+}99, Gut90, KC91b, KC96, KKL^{+}93, Köc94, Kul97, Pre94, SS96a, SR97, SAP97, Van97, Van94, Whe90]. Concepcion [Ald92]. concepts [BB92, BP94c, CG92, DKS96, Hon93, Hon94, PH92, VJ95]. concrete [Avr93, Bon93]. Concurrent [Van94]. condition [Ded96]. conditioning [Dem90, Var90b]. Conference [AS93b, Boi97, Bru98, Civ98, DDN95, Des96, EG91, GW90, GW92, GW94b, GHW98, HMRS90, HMRS92, HPO92, HZO92, HRV90, yKaY92, KFC97, LO90, Mor90, NS93, RRV92, SD95, Smi94, BB92, DSV94, LMH91, Mei96, Nei91, Rak94, Wes90]. Configuration [Uni98c]. Congress [Ald92, SJ96]. Conjectures [Var90d]. Conjugacy [Hes90]. conjunction [Cel90, Nog97]. Conservation [Tho99, LeV90, LeV92, Liu93, Pem92, Shu90]. Consiglio [NS93]. consistent [Bun90]. constitutive [DG90a, Ros90]. constrained [GZ96, Naz96]. Constraints [CGT97, Pyt99]. constructions [Roc97]. Consultation [GL90]. containing [TC98]. Context [SSES90]. continuation [AG90a]. continuous [KD92]. contribution [Kub90, Par90a]. Contributions [Ull90b, TC98]. Control [Pyt99, Uni99, Bul93, Cha98, Hig97a, KD92, Lip94, Sch96, SC96, CSV96]. convection [RST96]. convection-diffusion [RST96]. Convergence [Bro92, Stu97, Her96, HM96, Jam94a, Wat96]. convex [Nem96]. Convolution [KMRT90]. cooling [FS90]. coprocessor [SS96a]. couldn’t [Gro97a]. County [Mer95, Mer97]. coupled [HM96, Met96]. Coupling [Ton90]. course [DH95a, Ise96, Ort90b]. Cray [Bun90]. Creation [KUIO90]. Creek [Tho90]. crew [WSG91]. Critical [SSES90, KG93]. Curvature [Ell97, LS97]. curved [NN96]. curves [ER96]. cutter [ADK91]. cutting [Rak94]. CVD [Gok93]. cybertheatre [Sha98]. cylinders [Rei92]. czastkowych [KZSU91]. Czech [Rak94, Vit94]. Czechoslovakia [Bab90].

Dade [Mer97, Mer95]. DAEs [BS92b]. Dallas [Eme90, SM90]. damage [Arm97]. d’analyse [AP90]. Data [BR92, HD90, Pow97, AKM96, BL98, GC95, GB97, Hon93, Hon94, KLRK98, WZ94]. Datenanalyse [BL98]. Davis [AS93b]. days [Gut90]. Debugging [KUIO90]. December [HRV90, JKB99]. decision [Dor97]. decision-making [Dor97]. Decomposition [Hig90, SBG96b, SBG96a]. decompositions [Ste98b]. deep [VJ95]. defective [Dem90]. defects [Dob96]. definite [Hig90]. deformation [McV90]. degenerate [JP96]. degree [Adž96]. delay [EH97, KP99]. Dense [Hig97c, DH90]. Department [Nei91]. Dependent [Ell97]. DEQSOL [KUIO90]. derivatives [Euv94, Wer96]. dérivées [Euv94]. Descriptive [GM96]. design [Aki94, Bis97, BP94c, Byl94, Hud91, Jam94a, Jam94b, McV90, Pan97a, PW92]. designed [Sha98]. Designing [Du 97]. Designs [Ton90]. Determination [MDE90, RR96, Ost97, Yam96]. Developing [SR97]. Development [Cha95, McV90, Nak97, Rei92, Tes93, Bab90, DHP97, FS90, GS90, Kub90, MAA97]. Developments [Hig97c, BS92b, EG92]. Diagnosis [KUIO90, Ume97]. Diego [BCF95, BCF97]. diesel [Zho94]. diferenciālvienādojumu [Kal96]. diferenciálních [Vit94]. difference [Ses98, Tho95, Vin91]. different [Zei95]. Differential [BD90, C^{+}96, Coo97, DG90b, GO96b, GO92a, GM97, HKP^{+}90, HRP90, Ise97, KP92, yKaY92, LB92, MN99, MOF90, Mor97b, SW95, AIL92, Ain95, Ame92, AA97, Bab95, BRP99, B^{+}98, Bau98, Bet98, BEM96, Bun90, CG92, Deĭ94, Dor96, EH97, Fox90c, GV96a, GO92b, GO95, GR92, GR94, HP90, Her94, Hig97a, Ise96, IZ96, Kal96, KG91, KG93, KLP90, KP99, KZSU91, Lam91, Lan96, Lan99b, LP99, MS95, MM94, QV94, QV97, Ree91, RST96, Sch91b, SW97, SBG96a, Tho95, Tho99, TGK91, TW98b, Vit94, Wal96b, Zei95]. Differentialgleichungen [GO95, TGK91, GR92, GR94, SW95]. differentiation [Bis97]. differentsialnykh [AIL92, Deĭ94, KLP90]. difficulty [CC97]. Diffpack [Lan99b]. diffusion [Jam94a, Jam94b, RST96, Szi92, Tsu91, VK93]. diffusive [Elt90]. Digital [JSW93]. dimensional [Bro92, FMO96b, dG93, Rei92, Wat92]. dimensions [BGP96, MOS96]. dinner [Fox90b]. direct [EZ91, Duf97]. discrete [BEM96, Jam94b, Dés98]. discretization [PS96a]. discrets [Dés98]. Discussion [SSES90, Ree91, SCI^{+}96]. disintegration [Rak94]. disk [Ath95, Cha93a, FS90]. disk-cavities [Ath95]. dissipative [Arm97, Hag90]. distance [DGY96, Spä96]. Distributed [HH90]. distributions [Har99]. Division [Cel90, Cro94, VS94]. do [Gen97, MM96]. Documents [GRM97]. Does [SCI^{+}96]. Domain [SBG96b, SBG96a, Vin91]. domains [Giv92, LM91]. double [NL96]. DSP [Mor97a]. Dundee [GW90, GW92, GW94b, GHW98]. Dynamic [DJT99, Nog97, Pax95, Tes93]. Dynamical [Rei91, SH96, Stu97, Hon93, Hon94]. Dynamics [Bru98, DDN95, Des96, HPO92, KFC97, Mor90, NS93, Bai98, Bir90, Cha93a, Dur99, Erb94, Jam94a, Jam94b, MB93, MB95, Tor97, Tor99, VI90, VS94, Xu99].

Early [Fox90a, Tod90, Whe90]. earthquake [DJ91]. ECCOMAS [Des96]. ECONMIG [Smi94]. Econometrics [MN99, Dor97]. economics [Dor97, Jud98]. Edition [HL95, The97, HL97]. effect [Jam94a]. Effective [Pan96]. Effects [Uni98a, Mod98]. Efficiency [GN97]. Efficient [HN90, Nak97, PS96a, Vij95]. Eiffel [Dub97]. Eigenproblems [vG97]. eigenvalue [CB90, Saa92]. Eigenvalues [CCA93, Cha93c, BEM96]. Eighth [Wes90]. Einführung [GO95, GO96a]. electric [Tao94]. electromagnetic [bZ93a]. Element [BJ93, GRT93, HN90, SH97, Bäc99, Bel90, BS94, DFS97, Hac92, Hal94, IY94, Jol90, Kyt95, Lae93, Lae99, PC97, PH92, Ram94, RKA93, Rie96, Ses98, SG98, Sch91d]. Elementary [Atk93, SK93, SK95, Ric96, Ste96, Sto96]. Elemente [Sch91d]. elementov [Lae93, Lae99]. elements [Aki94, Moh92, Ode90, AP90, Jol90, AP90]. Elimination [STL96, GZ96]. Elliptic [BD90, DG90b, Tho99, Li90, Mat92a, Ree91, SBG96a]. Ellpack [HRP90]. Embedded [Mor92]. emissions [YCC96a, YCC96b]. Emphasis [Wat97]. enclosing [Rih96]. Enclosure [Wal96b]. enclosures [Wie96]. Engine [Uni98c]. engineer [LT94]. Engineering [Act96, Ald92, Cel90, Civ98, Cro94, Des96, G^{+}99, GW98, HN90, HPO92, MR96, MM96, SJ96, Ton90, Van96, VS94, CG92, CGB96, Cha93b, CS99, DJ91, Euv94, Fer98, HZO92, HSDM91, JM90, LP99, MK95, Mat92b, Nak97, PM90, Poz98, Smi94]. Engineers [BB99, Bor91, Cel90, Eme90, Jen77, Nog97, NL95, SM90, Sar92, Aka94, AM96, Bor97, CS95, CC98, CM99, Ett92, GS91a, GS91b, Hof92, JR92, JM92, Pal98, Roj96b]. England [Mor90]. English [GO95]. Enhancement [Ano97, Boi97]. Environment [BD90, MM96, DGLR90]. environments [SR97]. Epilogue [Fox90b]. EPSRC [ALM99]. equal [BCSS96]. equality [GZ96]. Equation [DPR90, Cer93, Lee94, LM91, Mei99, Tsu91, Wie96]. Equations [Atk97, Bak97, BD90, DG90b, Dab98, GO96b, GO92a, GM97, HMRS90, HMRS92, HKP^{+}90, HRP90, Ise97, KP92, yKaY92, LB92, MOF90, Mor97b, STL96, SW95, AIL92, Ain95, APS96, Ame92, AA97, Bab95, B^{+}98, Bau98, Bet98, BEM96, CG92, Chu91, C^{+}96, Coo97, Cox90, DKS96, Deĭ94, Dor96, Dur99, EH97, Euv94, Fox90c, Fra96, GV96a, dG93, GO92b, GO95, GR92, GR94, Hac91, Hac94, HP90, HRW94, Hig97a, Ise96, IZ96, Kal96, KG91, KG93, KLP90, KP99, Koe96, KZSU91, Lae93, Lan99b, LP99, Liu90, MS95, MM94, Naz96, Pie91, PS90, PS91a, PS91b, QV94, QV97, Ree91, Rhe98, RW96, RST96, Ros90, Sal98, Sch91b, SW97, Sha96, SK96, SBG96a, Tho95, Tho99, TGK91, TW98b, Vin91, Vit94, Wal96b, Woo92, Zei95, Euv94]. equilibria [Bri95]. Error [Die96, GZ96, LT96, Mik91, Sin96, SH97, AH96, BR92, CB90, Ehr96, Hig97a, MOS96, SK96]. Error-bounding [LT96]. Errors [Act96, Hat97, Olv90, SS96b]. Esercizi [FT92]. estimates [Ehr96, MOS96, SK96]. estimating [McV90]. estimation [Var90b]. etc [Uni98a]. Euler [Chu91]. EUROMECH [SO97]. European [Civ98, HPO92, HZO92, Smi94]. Evaluating [AG90b, HHRS90, Uni98a, Mod98]. Evaluation [Lyn97, Sin96, EH97, Mu97]. Evolution [DH90, Blo90, Xu99]. evolutions [Mat96]. Ewa [Oki96]. exact [RR96]. Example [HN90]. examples [KKN^{+}92]. exception [Rei97]. Exercise [Her98]. exercises [Don95, FT92]. experience [Met90]. Experiences [BK90, JM90]. Experimental [BCF95, Cou94, DJ91, BCF97, FS90, Guc95, Roc97, SO97, Tes93, Zho94]. Experiments [AD90, KPS94, Hat97, Ume97]. Expert [AEP90, BK90, BM90, DG90b, GGM90, Gro90, HRV90, HKP^{+}90, HRP90, SSES90].

Experts [CZ90, HD90]. Exploiting [AGM93]. Exposition [SJ96]. Extended [Sto96, KKL^{+}93, Kul97]. Extension [KKN^{+}91a]. extensions [Blo97].

fact [Ste96]. factorizations [Pai90]. faculty [Gir92]. failure [ADK91, SJ96]. Falling [Uni99, Cha98]. Falling-leaf [Uni99, Cha98]. fangfa [TJL92]. far [WZ94]. far-infrared [WZ94]. Fast [KU96, SB96b, Ano94a, Wal96a]. fastening [SJ96]. fatigue [BB92]. FEA [SJ96]. feasibility [KLRK98]. Federal [LO90, Mar92, MK95, MK98]. FFT [Ano94a, Coo90, KMRT90]. field [Bun90, Tao94, WSG91]. fields [bZ93a]. Fifteenth [KFC97]. Fifth [LO90, SD95]. fifties [Whe90]. filled [Pan96]. film [FS90]. films [Par91]. filter [Str96]. Filtering [EZ91, GM97, GW99a]. Filters [Vel98]. final [WZ94]. finance [RT97]. finding [Nov96, Vas96]. fine [Erb94]. fine-scale [Erb94]. finis [Jol90]. Finite [Aki94, BJ93, CCF96, HN90, Moh92, SH97, Bäc99, Bel90, BS94, Die96, DFS97, JP96, Jol90, Lee94, Ode90, PH92, Rie96, Ses98, SG98, Tho95, Vin91, BJ93, GRT93, Lae93, Lae99, Sch91d]. finite-difference [Vin91]. Finite-Element [BJ93, Jol90, Sch91d, GRT93, Lae93, Lae99]. Finite-Element-Methode [BJ93, GRT93]. finite-part [Die96]. finiten [Sch91d]. Finland [Nei91]. First [BC93, HPO92, HZO92, HRV90, DH95a, Ise96, VWY97]. Fitted [MOS96]. fitting [Spä96]. flames [Rig96]. flexure [Fri97]. Floating [Sin96, SS96b, Cou94]. Floating-point [SS96b]. floodplain [Hib93]. Florida [Mer97, Mer95]. flow [Ath95, LSR90, Mer95, Mer97, Oki96, RST96, Sie94, Ver92, Vre94, WSG91]. Flowfield [Uni98c]. flows [BCH93, Cel90, Cha92, Cro94, FMO96b, Lee94, Rei92, Ses98]. Fluid [Bir90, Bru98, DDN95, Des96, HPO92, KFC97, MO95, Mor90, NS93, RRV92, Wes90, Ano94b, Bai98, Dur99, MB93, MB95, Tor97, Tor99, Vin98, VS94]. Fluids [Cel90, Cro94, VS94, Moh92, Ros90]. flux [Shu90]. Following [STL96]. food [CS95]. Forces [Mar92, MK95, MK98, Cou94]. forebody [Ago94]. forecast [GhW90]. forecasting [Sme96]. form [BV90b]. Forming [SD95, Sar92]. formula [Ehr96]. Formulation [Arm97]. formulations [Mat92a]. Forsythe [Var90a]. Fortran [Ett92, PTV^{+}92, Pre93, Pre94, RS92b, Bor91, Du 97, EMU96b, NL95, Ort94b, Ort94a, PTM96, PTV96, RS92a, Rei97, WNO94]. Forum [Cel90]. fossils [Har99]. Four [Uni98c]. Four-Engine [Uni98c]. Fourier [Ano94a, Wal96a, GW99a]. Fourteenth [DDN95]. fourth [Civ98]. fracture [LO90]. France [Bru98, Des96, JKB99]. free [Nei91]. Freiburg [LO90]. French [AP90, Dés98, Euv94, Jol90, LT94, RP98, Sch91a]. frequency [Guc95]. FRG [HMRS90, Mar92]. fronts [Fin92]. Fry [ACD^{+}93]. Fuel [YCC96a]. function [Nas90a]. Functional [Bak97, IS93, Ras94, LL97, Mun97]. functions [Loz97, Sam96, Sny97, Ste93]. fundamental [CG92, LS97, BP94b]. Fundamentals [Ano94a, FK94, KS94, SAP97, Wat91, KS93, PC97]. Fure [Zhu92]. Fuzzy [BD96].

G [LL97]. gained [Coo90]. galaxies [Lam93]. galaxy [Lam93]. games [BRP99]. GAMM [AH96, Hac92, Mar92, MK95, MK98, RRV92, Wes90]. GAMM-Conference [Wes90]. GAMM-Seminar [Hac92]. GAMM/IFIP [Mar92, MK95, MK98]. GAMM/IFIP-Workshop [MK95, MK98, Mar92]. Gas [Xu99, Ath95, Elt90, Gok93, Jam94a, Jam94b, VI90, YCC96a, YCC96b]. gas-phase [Gok93]. gas-turbine [Ath95]. gaseous [FMO96a]. Gatlinburg [Var90c]. Gauss [Ehr96]. GCD [EGL96]. general [Bro92, Nak97, Pax92]. generalized [Pai90]. generated [GB97]. Generation [BGGD90, PWF90, Bab95, DFS97, KS93, KS94]. Generator [DPR90]. generic [Ago94, Cha98]. gentle [GKT14]. Geo [Nog97]. Geo-Institute [Nog97]. Geo-Logan [Nog97]. Geohydrology [LSR90]. geological [Car95]. Geomechanics [Rak94, DG90a]. Geometric [BP94c, SS97a, Bar96b, Sal99, ZAB95]. Geometrical [Sch80]. geometricheskie [ZAB95]. geometry [Dem90, Roy96, SW96]. geophysical [Dur99]. George [LL97, Var90a]. Georgia [SJ96]. geostatistics [Lee97]. geotechnical [Civ98, Smi94]. German [BL98, BJ93, DH91, DH93, DB94, GRT93, GO95, GO96a, GR92, GR94, Hac91, HH94, Her95, Her98, Mei99, Opf93, Sch91c, Sch91d, Sch93, Sob91, SB90, Sto93, Sto94, SW95, TGK91]. Germany [AH96, HMRS92, LO90, MK95, MK98]. gewöhnlicher [SW95]. given [MW86]. Gleichungssysteme [Hac91, Mei99]. global [GN97, Kel96]. globalization [Naz96]. GMRES [Erh96]. GNU [Eat97]. godografov [Bur93]. Godunov [Xu99]. Godunov-type [Xu99]. goes [Ste98a]. going [Cor96]. gradients [Hes90]. Graduate [ALM99, Ste98a]. graphic [Nak96]. graphical [Fiu95]. Graphics [BM90, GRM97, Kel96, Sal99]. Grenoble [JKB99]. Grid [MDE90, Vel98, Cer93, KS93, KS94, Sho97]. grids [Cha95]. großer [Hac91]. Ground [Uni98a, Hud91, LSR90, Mod98, Oki96]. Ground-Water [Uni98a, LSR90, Mod98, Oki96]. groups [McV90]. guaranteed [Wan91]. Guide [ALM99, BCC^{+}97, For96, HL95, The97, Tur90, Dor97, HL97, MW93, Boi90].

H [Coh90, Par90a]. Hadamard [Die96]. Hadamard-type [Die96]. half [Gir92]. Handbook [CL90, Zwi92, PES99, RC98]. handling [Rei97]. hardware [GS90]. Harten [TC98]. Hartree [Bun90]. Hawaii [Oki96]. head [Cha93a]. head-disk [Cha93a]. health [NMW96]. heat [Eme90]. Heidelberg [HRW94]. held [Ain95, Ald92, AH96, BB92, Bru98, Cel90, DDN95, HRW94, HMRS90, HMRS92, KFC97, LMH91, LO90, Mar92, MK95, MK98, Mor90, NS93, Nei91, Nog97, SO97, TC98, Ull90b]. Hermitian [LG90]. High [DDSvdV98, FJSD96a, FJSD96b, HL95, The97, Cha98, Eat97, Lam93, Liu93]. high-level [Eat97]. High-Performance [DDSvdV98, FJSD96a, FJSD96b, HL95, The97, Cha98]. higher [BEM96, NL96, Rei92]. higher-order [BEM96]. Historical [Ode90, Roc97, You90]. History [Nas90b, Tod90]. Homogenization [Liu90]. homotopy [LWW96]. honour [Mei96]. Howard [Coh90]. Hradec [Rak94]. Hybrid [NA90, RKA93]. hydrogeology [Hib93]. hydrologic [Hib93]. Hyperbolic [EG91, Mor97b, SH97, Koe96, Liu90, Pem92, Ros90]. hypercircle [Rie96]. hypersurface [DGY96]. hypertext [BHR96].

ice [Cou94]. Idaho [LSR90]. identification [SO97]. If [Gen97]. IFAC [HRV90]. IFIP [Boi97]. IFIP-Workshop [MK95, MK98, Mar92]. II [HMRS92, Cuy94, DB94, HRP90, YCC96b]. II. [YCC96a, YCC96b]. ill [BCF95, BCF97, Dem90, PS96a, Tik95]. ill-conditioning [Dem90]. ill-posed [BCF95, BCF97, PS96a, Tik95]. Illinois [Whe90]. illumination [Bor90, Kel96]. Illustrating [GRM97]. im [MR96]. IMACS [AH96, HRV90]. IMACS/IFAC [HRV90]. Image [GM97, Vel98, Blo97, Bor90]. Imaginary [LG90]. imaging [BCF97]. impact [Cou94, GhW90]. impinging [Gok93]. Implementation [Lac95, Sch96, Jol90]. implementations [RHP91]. implementing [Hol97]. Implicit [STL96, Lee94]. improperly [IY94]. Improved [Fri97, PW92]. Improving [Ost97, Pan97a]. INA [HT91]. Inclusion [Koe96, Pet96, PTŽ96]. incompressible [FMO96b, Rei92]. Incorporating [Gro90]. index [LT96, Olv90]. India [DDN95]. Indicators [MDE90, JP96]. Industrial [SD95, DT97, Sar92]. industry [Ano92b]. inequalities [Ste90]. infeasible [JP96]. infeasible-interior-point [JP96]. Inferno [Cou97, Gro97b]. infinite [Giv92]. influence [Var90a]. information [Dor97, Pla96, Woz96]. information-based [Woz96]. infrared [Lam93, WZ94]. Ingenieria [Ald92]. ingénieur [Euv94, LT94]. inhomogeneous [FM92]. initial [Hig97a, Koe96, Lam91]. input [AKM96, Bro92]. instability [Bri95, Xu99]. Institute [HT91, Nog97]. insurance [NMW96]. integers [Pan96]. Integral [Atk97, Bak97, Fra96, Gol99, PS91a, PS91b]. integrals [Die96]. Integrated [Cha92, Cha95]. Integration [DDLW90, SS97a, Zwi92, EG92, Eva93, Gu93, KU94, RKA93, Sch91b, Sch96, Smi93]. Intelligence [DDLW90]. Intelligent [HRV90]. Intensive [HS93]. interacting [Lam93]. interaction [Ath95, Bul94, MO95, Nog97]. interactions [Hib93]. Interactive [BD90, KUIO90, Eat97]. Interactive/Visual [KUIO90]. Interface [Ell97, PWF90, SSES90, FMO96b]. Interior [SS97c, JP96, Shi95]. interior-point [Shi95]. International [Ald92, AH96, AS93b, BC93, BC94, BC95, Bru98, DDN95, EG91, HRV90, KvSR90, KFC97, LO90, Mor90, NS93, SJ96, SD95, DSV94, HRW94, LMH91, Rak94, VWY97]. Interpolation [Lor92, Pow97, BGP96, Rei90]. interpreting [Ano92a]. Interval [Sme96, KLRK98, Sha96, Sto96]. intra [Ath95]. intra-cavity [Ath95]. Introduction [Bel97, FJSD96a, FJSD96b, GO93, GO96b, LP98, Man99, Ort94b, Ort94a, Pal98, RP98, SB93, SB96a, TW98b, Van97, Woo93, Zac96, Zac98, AG90a, EWK90b, Fiu95, Flo95, Gau97b, GO92b, GO95, GO96a, GKT14, Kyt95, MM94, OG99, Pan97b, Ply92, Tor97, Tor99, Tyr97, Kon93, RP98, Rya94]. Introductory [Hea97]. Inverse [AEP90, BCF95, BCF97, Bro92]. investigation [Vin91]. invited [Des96, HPO92]. Iosipescu [Ho94, Ho92]. IRAS [Lam93]. issues [SJ96]. Italian [BBCM92, QSS98]. Italy [Civ98, NS93]. itératifs [Dés98]. iteration [AS93a]. Iterative [Bjö90, Gv97, GM97, Hac94, AAK96, Car96, Dés98, HM96, Mat92a, Ste90, You90, Hac91]. Ithaca [SD95]. Ithaca/New [SD95].

J [LRRL97]. J. [Par90a]. January [Hac92]. JavaScript [Sha98]. Jean [Mei96]. Jersey [Mod98]. jet [Gok93, Rak94]. Jim [Fox90b, Gol90b]. Jisuan [TJL92]. joining [SJ96]. Joint [Ju95, MW86]. Jordan [BV90b]. JP [Zho94]. JP-5 [Zho94]. July [Ain95, AH96, BCF95, BCF97, Boi97, Bru98, DDN95, LMH91, Mor90, NS93, Nei91, Nog97, RSS96, AGM93, GW94b]. June [Cel90, Cro94, EG91, GW90, GW92, GHW98, Ju95, KFC97, MK95, MK98, SD95, VS94, VWY97, GW94b]. June-July [GW94b]. Jyvaskyla [Nei91].

Karlsruhe [Gee90]. Kerkrade [SO97]. Kiel [Hac92]. kinetic [Rig96]. Kingdom [Boi97, Fox90a]. KIVA [YCC96a, YCC96b]. KIVA-II. [YCC96a, YCC96b]. KKT [Shi95]. KKT-based [Shi95]. Knowledge [BGGD90, FHI90, HN90, KMRT90, MDE90, KSH91]. Knowledge-Based [FHI90, KMRT90, MDE90]. Koenig [AAK96]. kompyutrni [KV97, VK96]. konechnykh [Lae93, Lae99]. Kronrod [Ehr96]. Kutta [NL96, Sch96]. kvadratov [ZAB95, Zor95].

lab [RC96]. laboratory [BZ93b, Don95]. Lagrangian [Naz96]. Lake [Cro94, VS94]. Lanczos [LG90, Par90b]. land [WO90]. Language [Gro90, Eat97, KKN^{+}92]. Large [Noc97, STL96, AD90, Hac91, Hac94, Mu97, Saa92]. large-scale [Mu97]. laser [Par91]. late [Poo97]. lattice [Elt90, LeB93]. Latvian [Kal96]. Lausanne [RRV92]. Laws [Tho99, LeV90, LeV92, Liu93, Pem92, Shu90]. layer [Adž96]. layered [VY96]. Layers [Uni98b]. Lazy [Ric96]. LCP [JP96]. leaf [Cha98, Uni99]. lean-combustion [YCC96b]. Least [ZAB95, Bjo96, Cox90, LeB93, VY96, Zor95]. Least-squares [ZAB95, Zor95]. Lecture [ALM99, Ain95, TC98]. Lectures [CCF96, RG90, Des96, HPO92, MW86, Ste96, Ste98a]. Leicester [Ain95]. Leningrad [Kub90, LL97]. level [Eat97, LT96, Olv90]. level-index [LT96, Olv90]. Lewis [ACD^{+}93]. Libraries [Gee90, Gro97a, Köc94]. Library [DHP97, Du 97, KKL^{+}93]. Lie [Her94]. Lie-theoretic [Her94]. like [Pan96, Wat96, Yam96]. limitations [LS97]. limiters [Jam94a]. Linear [BV97, Dem97, DDSv91, DDSvdV98, Duf97, Gen98, Gv97, Gro90, Hig97c, Lan96, SS97c, TB97, AKM96, BCS91, Bri95, Cox90, Dat95, DH90, GMW91, Hen92, Hig97b, Ho94, Kon93, KU96, Kub90, Mei99, MOS96, Nas90a, Ree91, Sha96, Ste90, Woo92]. linearer [Mei99]. linearization [Don96]. lineinoi [Kon93]. lines [Sch91b]. link [Gir92]. liquefaction [AS93b]. liquid [FS90]. Liquifaction [AS93b]. Liverpool [MW86]. loaded [McV90]. Local [Don96, Vel98, Fra96, Hig97a, LP90]. localization [VI90]. localized [Arm97]. location [BEM96]. Logan [Nog97, Nog97]. Logic [HKP^{+}90]. long [Sme96]. longwall [VJ95]. Lorentz [LL97]. losers [NMW96]. Lösung [Hac91, TGK91]. low [Adž96, YCC96a, YCC96b, Lee94]. low-degree [Adž96]. low-Reynolds [Lee94]. LU [Lee94].

machines [Bun90]. Made [Act96]. magnetotail [Bri95]. making [Dor97]. Manchester [MW86, Smi94, TC98]. Mandarin [TJL92, WY90]. Manipulation [Ton90]. manipulator [Gir92]. Manual [KC93, Don95]. MAPLE [GH93, Bar96a, B^{+}98, Bet98, C^{+}96, Coo97, Cor95, GH95, GH97, RC96, Zac96]. Mapper [HHRS90]. mappings [Sto96]. March [Ano94b, Van96]. Market [BPRB97]. Maryland [Ste96]. massive [Lam93]. massively [Bas96]. Mastering [HL96, HL98]. Matematica [QSS98]. matematicheski [KV97, VK96]. Matematicheskie [Pru93]. matematike [Zhu92]. matematiku [Rya94]. Material [SO97, Byl94]. Materials [Ju95, Bul96, Ho92, Ho94, SD95, TW98a]. Mathematica [Bau98, GV96a, SW97, SK93, SK95, Zac98]. Mathematical [BP94a, Boi90, GL90, LP90, Ric90, Var90d, WNO94, BS94, CH53, CH62, Gol99, HR99, HRV90, JJ46, JJ56, JJ99, KV97, KW94, LS97, Sch80, Sho97, VK96, Wak90, Pru93]. mathematicians [Kub90]. Mathematics [AGM93, CK94, GW98, KC93, LL97, Nei91, Tod78, Ano94b, Bab90, BZ93b, BGR90, DT97, DH91, DH93, DB94, HH91, HH94, Her98, Isa90, KC91b, KC96, Mat92b, QSS98, RSS96, RG90, Rya94, Sch91c, Sch93, SB90, Sto93, Sto94]. Mathematik [Opf93, DH91, DH93, DB94, GO95, HH94, Sch91c, Sch93, Sch97, SB90, Sto93, Sto94, Var90c, Her98]. MATLAB [BD96, BB99, Bor97, CSV96, CM99, Don95, Fau99, GH93, GH95, GH97, Got95, HL96, HL97, HL98, MF99, MM96, Nak96, Pal98, PES99, PL95, RC98, ST98, TW98a, Van97, HL95, The97]. Matrices [BP94a, Cha93c, CCA93, Dem90, Erh96, Pan96]. matricielle [LT94]. Matrix [Bha97, BP94b, CZ90, GHN^{+}90, GV96b, Got95, Hig90, HJ90, HJ91, Jen77, JM92, LG90, MN99, SS90, Ste98b, Wat91, Bel97, BHR96, BPRB97, BEM96, Cox90, Van97, BPRB97, LT94]. matrix-vector [Van97]. maxima [Sme96]. maximum [MOS96]. Maxwell [Dab98, Vin91]. May [BB92, yKaY92, Mar92, Sie94]. measure [Gen97]. mécanique [Euv94]. Mechanical [Cel90, Eme90, SM90, SJ96, Sar92, Bon93, Rak94]. Mechanics [RRV92, Wes90, Ano94b, BS96, CG97, Euv94, Her94, Hro93, Ju95, LO90, PBW90, Rak94, TW98a, Vin98]. mechanism [Arm97]. media [FM92]. medical [BCF95, BCF97, NMW96]. Meeting [Cel90, Cro94, Eme90, Ju95, SM90, Sar92, VS94]. Meinguet [Mei96]. melting [Par91]. memorial [LRRL97, TC98]. Memory [DDSv91]. meromorphic [Sam96]. mesh [Bab95]. metal [Sar92]. Meteorology [RD93, Ano92a, ACD^{+}93]. Method [BJ93, GRT93, Sob91, TJL92, Dan90, Don96, DGLR90, GMSW90, Hal94, Hro93, IY94, Jol90, Lac95, Lae93, Lae99, Man99, PC97, PH92, PTŽ96, PS96b, Sch91b, Sch91d, SG98, Sob94, Ste90, ZAB95, Zor95]. méthode [Jol90, BJ93, GRT93, Sch91d, Sob91]. Methoden [BL98]. methodology [BCFH97]. Methods [Ald92, Ano92a, Bor91, Bru98, CGT97, CH53, CH62, DDN95, DS90, Duf97, For96, GO96b, Gra90, GW98, HS93, HMRS90, HMRS92, JSW93, JJ46, JJ56, JJ99, Kel93, yKaY92, KFC97, LB92, MDE90, MK95, Mor92, Mor90, NS93, NL95, Opf93, Pow97, Pyt99, RS92a, Rhe98, Ric92, RRV92, SS97c, SH97, Uni98a, Uni98b, Ueb97, Ull90b, VI90, Wes90, dS97, Abe98, Act90, AIL92, Aka94, AH96, AG90a, Ame92, Ano92b, AS93a, AAK96, AM96, Bab95, BV90a, Bai98, BCF95, BCF97, BRP99, Bas96, BCH93, Bel90, BBCM92, BS96, Bjo96, BL98, Blo97, BP93, Bor90, Bor97, BL94, BS92a, BJ92, BS94, Bri95, Bro92, BT92, Bul93, Byl94, BS92b, Car96, CG92, Cel90, Cer93, CGB96, Cha93b, Cha92, Cha95, CS95, CC98, Chu91, Civ98, CG97]. methods [CM99, Cro94, CS99, Cuy94, DT97, DG90a, Des96, DSV94, Dob96, DJ91, Dor97, Dor96, DJT99, Dur99, Dab98, Elt90, Eme90, EG91, Erb94, Ett92, FB93, FB98, FMO96a, FMO96b, Fer98, Fin92, Flo95, Fri97, GZ96, Gar94, GS90, Giv92, dG93, Gol99, GO92b, GO95, GR91, GS91a, GS91b, Gu93, Guc95, Hac92, HRW94, Hen92, Her96, HPO92, HZO92, Hof92, Hol97, Hon93, Hon94, HSDM91, Hud91, IS93, IK94, JR92, JP96, JB94, JKB99, Ju95, Jud98, Kal96, KG93, KLP90, Kel96, KP99, Köc94, Kon93, KV97, KZSU91, KSH91, KD92, Kyt95, Lam91, Lan99b, LRRL97, LeV90, LeV92, LMH91, Li90, LM91, Liu90, Liu93, LO90, Mar94, Mar92, MK98, Mat92a, Mat92b, MF99, Mer95, Mer97, MOS96, Mod98, MO95, Mor97a, MB93, MB95, Nak91]. methods [Nak93, Nas90a, NL96, Nei91, Nem96, RF96, OG99, PM90, PBW90, PLC94, Pem92, PL95, Pet96, Pie91, Pru93, Rak94, Ram94, RS92b, Ree91, Rei92, RKA93, Ric93, RHP91, Rig96, RT97, Roj96b, RST96, Ros90, Saa92, SM90, Sal98, SJ96, Sar92, SS97b, Sch80, Sch96, Ses98, She91, SD95, Shi95, Sho97, Shu90, Sie94, Smi93, Smi94, SBG96a, SO97, Ste93, SB96b, SS92, Szi92, Tho95, Tik95, Tor97, TC98, Tor99, Tsu91, Ull90a, Van96, Ver92, Vin98, Vin91, Vit94, VK96, Vol90a, Vol90b, VS94, VK93, Vre94, Wal96b, WO90, Whi98, Woo93, Yam96, You90, Zei95, vL97, vS99, Deĭ94]. Metod [Lae93, Lae99, Zhu92, ZAB95, Zor95, Vit94]. metodes [Kal96]. Metodi [BBCM92, KV97, VK96]. Metodos [Ald92, Vol90a]. Metody [Bur93, Deĭ94, KLP90, AIL92, BV90a, Kon93, KZSU91, Nur91, Pru93, Tik90]. Mexico [GH94]. Michigan [Var90c]. micro [Ses98]. micro-channels [Ses98]. Microcomputers [BGR90]. microstructures [FM92]. mines [VJ95]. minimal [LWW96]. Minimax [SS92]. minimisation [Nas90a]. minimization [Cal96, Lyn97, SJ96]. Misconvergence [Par90b]. Mise [Jol90]. mixed [Mat92a, Shu90, SO97, Wer96]. mixing [GC95]. mixture [FMO96a]. mnogomernykh [Lae93]. model [Arm97, CB90, Hol97, LP90, LSR90, Nog97, Pax92, PW92, Pax95, SCI^{+}96]. Modèles [Dés98]. modeli [Pru93]. Modeling [Bab95, HHRS90, Uni98a, ADK91, GC95, GB97, Hib93, Lam93, Lee97, Mod98, Mun97, Nog97, Sal99, Tho90, VJ95]. modelling [Bon93, Bul94, Bul96, DG90a, Har99, KW94, LRRL97, Wak90]. models [Bor90, Cha98, Dés98, Lam93, Lee94, Pru93, SS92]. Modern [KV97, VK96]. modification [AAK96, VJ95]. molecular [PLC94]. Monitoring [Uni98a, Hud91, Mod98]. monoxide [Rig96]. Monte [Sob91, Hei96, Kel96, Man99, Nak97, Sob91, Sob94]. Monte-Carlo [Man99]. Monte-Carlo-Methode [Sob91]. Monterey [KFC97]. Moscow [LSR90]. Motion [Ell97, Cal96, Zho96]. movie [Sha98]. moving [Fin92]. MTNS [KvSR90]. MTNS-89 [KvSR90]. multi [Cha92, Ver92]. multi-body [Cha92]. multi-phase [Ver92]. multiblock [Lac95]. Multibody [vS99]. Multidimensional [Mor97b, Lae93]. multifrontal [DGLR90]. multigrid [BGP96, Jam94a]. multilevel [DJT99, SBG96a]. multiphase [Cel90, Cro94]. multiple [Ath95]. multipole [SB96b]. multiscale [DKS96]. Multivariable [Pow97]. Multivariate [Lor92]. München [MK95, MK98]. Munich [Mar92, MK95, MK98].

N2M [JKB99]. Nachala [LG95]. NAG [DHP97]. naimenshikh [ZAB95, Zor95]. Nankai [yKaY92]. Nash [CL96]. National [Tod90]. naukowo [MM96]. naukowo-technicznych [MM96]. Navier [HMRS90, HMRS92, HRW94, Sal98, Wie96]. Nazionale [NS93]. NBS [HT91]. NBS-INA [HT91]. Nearest [Dem90]. need [SCI^{+}96]. nekorrektnykh [Tik90]. Netherlands [SO97]. Network [Cou97, Hud91]. Network-based [Cou97]. Networks [FFK90, Sie96]. Neubiberg [Mar92, MK95, MK98]. Neubiberg/München [MK95, MK98]. Neural [FFK90, Sie96]. Neutrons [JKB99]. Nevada [Cro94, VS94]. Newton [Don96]. Newton-type [Don96]. Newtonian [VS94]. Next [PWF90]. Ninth [RRV92]. NOAA [GhW90]. noisy [Pla96]. Non [Ho94, Cal96, VS94]. non-adaptive [Cal96]. Non-linear [Ho94]. non-Newtonian [VS94]. nonconforming [Li90]. Nonconnah [Tho90]. Nonlinear [CGT97, Cuy94, Rei91, STL96, SS97c, Vel98, APS96, Elt90, MAA97, Naz96, Ree91, Rhe98, RHP91, Woo92, Zei95]. nonmedical [BCF95, BCF97]. Nonnegative [BP94a]. nonnormality [CC97]. Nonoscillatory [Liu93]. nonsymmetric [Bas96]. norm [MOS96]. Note [PS96b]. Notes [ALM99, Ain95]. November [Ald92, Eme90, SM90, SJ96, Sar92]. Nozzle [Uni98c, YCC96a]. nozzle/rich [YCC96a]. nozzle/rich-burn [YCC96a]. NP [BCSS96]. nuclear [Nak97]. nucleation [Gok93]. Number [KMRT90, Ded96, Lee94, Roj96a]. numbers [GM96, MW96, Ric96, Wie96]. Numeric [HL95, The97]. numerica [QSS98]. Numerical [Act90, Ago94, Ain95, ALM99, Ald92, AH96, AI98, AG90a, Ame92, Ano92b, Ano97, AA97, Asa95, ACD^{+}93, Ath95, Atk97, Avr93, AM96, Bai98, BC93, BC94, BC95, Bak97, BGGD90, BS96, Bjo96, BP93, Bor90, Bor91, Bor97, BL94, BS92a, BJ92, BCS91, BZ93b, Bru98, BT92, Bul96, BF93, BF97, Car95, CG92, Cel90, Cer93, CGB96, Cha93a, Cha93b, Cha92, CC98, Cha98, CK94, CG97, CM99, Cro94, DT97, DPR90, Dat95, Dem97, DG90a, DDN95, Des96, DH95b, DH95a, DS90, DHP97, DDSvdV98, Dor96, Duf97, Dur99, EWK90a, EWK90b, EGL96, EMU96a, EMU96b, EG92, FS90, FB93, FB98, FMO96a, FMO96b, Fer98, Fin92, FHI90, GV96a, Gar94, Gau97a, Gau97b, Gen98, GMW91, Gir92, Giv92, Gla93, dG93, Gok93, GO96b, Gra90, GR91, GW90]. Numerical [GS91a, GS91b, GW92, GW94b, GW96, GHW98, Gu93, Hac92, HP90, HHKR93, HH91, Har99, HRW94, Hen92, HT91, HMRS90, HMRS92, Hib93, Hig96, Hig97c, HPO92, HZO92, Ho92, Hof92, Hol97, HRV90, JSW93, JB94, Ju95, Jud98, Kel93, Ket99, KP99, KC91a, KC91b, KC93, KC96, KP92, KPS94, Köc94, KUIO90, KW94, Kre98, Kul96, yKaY92, KD92, KFC97, Lam91, Lan99a, LP99, LRRL97, LeB93, Tal90, LeV90, LeV92, LMH91, Li90, Lip94, LO90, Mar94, ML90, MK98, Mat92b, MF99, Mei97, Mei96, MGW96, MS95, MOF90, Mor92, Mor97a, Mor90, MB93, MM94, MB95, Nak96, NS93, Naw94, Nei91, NL95, RF96, Oki96, Ort90a, Ort90b, PM90, PBW90, Par91, Pem92, PL95, Pie91, Poz98, PFTV92, PTV^{+}92, Pre93, Pre94, PTM96, PTV96, PS91a]. Numerical [PS91b, Pyt99, QV94, QV97, RS92a, RS92b, Rei91, Ric92, Ric93, RHP91, Rig96, RRV92, RT97, RST96, Ros90, Saa92, Sar92, SSES90, Ses98, SD95, Shu90, Sie94, SK93, SK95, Smi93, Smi94, SM91, SW96, Ste93, Stu97, Szi92, Tho95, Tho99, Tik95, Tod62, Tod78, Ton90, TC98, TB97, Tsu91, Uni98a, Uni98b, Uni98c, Uni98d, Uni99, Ueb97, Ull90b, VJ95, Van96, Ver92, Vin98, Vol90b, VS94, VK93, Vre94, VWY97, Wak90, Wan91, Wat92, WG96, WSG91, Wes90, Whi98, Wie96, YCC96a, YCC96b, Zei95, bZ93a, Zho94, Zho96, dS97, vL97, Abe98, ADK91, Aka94, APS96, AGM93, AA96, Ano92a, Ano94b, Ano96, Arm97, AS93b, Atk93, AP90, Bab95, BCF95, BCF97, BRP99, BCH93, BB92, Bis97, BL98, Blo90, Blo97, Boi97, Bon93]. numerical [Bre92, Bri95, Bro92, Bul93, Bul94, Byl94, BS92b, Cha95, CS95, CCL99, Chu91, CL90, Civ98, Cou94, CH90, CS99, Cuy94, Ded96, DSV94, DJ91, DH90, Dor97, Du 97, Dub97, DJT99, DW97, Eat97, EZ91, Dab98, Elt90, Eme90, EG91, EH97, Erb94, Ett92, Eva93, Eva95, Fau99, FM92, Flo95, Fox90a, Fox90c, FP92, FK94, Fri97, GW99a, GS90, GW94a, GW99b, GC95, GB97, Gol99, GO92b, GO95, GH94, Gro97a, GhW90, Guc95, Hag90, HR99, Her94, HM96, Her98, Ho94, Hon93, Hon94, HSDM91, Hud91, IS93, IK94, Ise96, IZ96, JM90, JR92, Jam94a, Jam94b, JKB99, Kal96, KG91, KG93, KLP90, KZSU91, KSH91, Kub90, Lac95, Lam93, Lan99b, LT94, Lee94, Lee97, Lig91, LM91, Liu90, Liu93, LG95, LL97, LSR90, Mar92, MK95]. numerical [Mat92a, MAA97, Mer95, Mer97, Mik91, MOS96, Mod98, MO95, Nak91, Nak93, Nas90a, NMW96, Nog97, Nov96, OG99, Osh94, Par90a, Pax92, PW92, Pax95, PLC94, PT96, Ply92, Poo97, PS90, Rak94, RP98, Ras94, Ree91, Rei92, RSS96, RKA93, RD93, Roc97, Roj96b, RG90, SM90, Sal98, SJ96, Sch90, Sch91b, SS97b, Sch96, SAP97, She91, Shi95, SCI^{+}96, SO97, SC96, Ste96, Ste98a, SB93, SB96a, SH96, SS92, Tao94, Tes93, Tho90, Tor97, Tor99, Tur90, Tyr97, Ull90a, Ume97, Vij95, Vin91, Vit94, VI90, Woo93, WZ94, Xu99, vS99, BV90a, BBCM92, DH91, DH93, DB94, Euv94, FT92, GR92, GR94, HH94, KU94, McV90, Mei99, MW86, Opf93, QSS98, Sch91a, Sch91c, Sch93, SB90, Sto93, Sto94, SW95, TGK91, Vol90a]. numerici [BBCM92]. numerických [Vit94]. Numerico [FT92, RRLA90]. Numericos [Ald92, Vol90a]. numerics [Ain95, GO96a]. Numerik [GO96a, GR92, GR94, Mei99, SW95]. numérique [AP90, Euv94, LT94, RP98, Sch91a]. numerische [BL98, GO95, DH91, DH93, DB94, HH94, Opf93, Sch91c, Sch93, Sch97, SB90, Sto93, Sto94, TGK91, Var90c]. numerischen [Her98]. numerycznego [KZSU91]. NUMGE98 [Civ98]. NUMIFORM [SD95].

O.D.E. [RC96]. Oahu [Oki96]. Oaxaca [GH94]. Oberwolfach [HMRS90, HMRS92]. Object [Dub97]. object-oriented [Dub97]. obliczen [MM96]. obrashcheniia [Bur93]. Observation [Nog97, Cox90]. observations [Lam93, WZ94]. obyknovennykh [AIL92, Deĭ94]. oceanic [LRRL97]. Oct [Ull90b]. Octave [Eat97]. October [AS93b, Civ98, HRW94, Ano94b]. ODE [Cor96, GS90, Her94, SR97]. ODEs [BK90, Bar96a, BS92b, Rih96]. ODEs/DAEs/PDEs [BS92b]. œuvre [Jol90]. off [CB90]. Old [vG97]. Oldenburg [AH96]. one [Bro92, FMO96b, MOS96, Sny97, Wat92]. one-dimensional [Bro92, FMO96b]. Ontario [Cel90]. operation [SC96]. Operator [DKS96, PS90, PS91a, PS91b, Sam96, SK96]. operator-valued [Sam96]. operators [Lan96]. Optimal [Bul93, Mat96, Pyt99, Cal96, Fra96, Sch96]. optimisation [BCFH97]. Optimization [CGT97, MK95, MK98, MW93, Noc97, GMW91, GH94, GKT14, Lac95, Mar92, Moh92, Naz96, SM90, Shi95, Van96]. optimizatsii [Nur91]. order [BEM96, Her96, HM96, Liu93, Met96, NL96, Rei92]. Ordinary [Ise97, SW95, AIL92, Ain95, AA97, Deĭ94, Fox90c, Hig97a, IZ96, Lam91, MS95, Wal96b, Zei95, Kal96]. oriented [Dub97]. origin [Isa90]. Origins [Dan90]. Orthogonal [Spä96, BV97]. orthotropic [NN96]. osnovy [Lae99]. Ostrava [Rak94]. output [Ano92a]. overset [Cha95]. Overview [Woz96, SB96b, Wat96]. Oxford [Boi97, Mor90].

P [BCSS96]. P.C. [WNO94]. Padé [GGM90]. Pairs [HHRS90]. palaeobiology [Har99]. Panel [SCI^{+}96]. Papers [ACD^{+}93]. parabolas [Spä96]. parabolic [Lae93]. parabolicheskikh [Lae93]. Parallel [FFK90, FP92, GHN^{+}90, GO93, HHRS90, HRP90, PTM96, PTV96, TPD90, Bas96, Byl94, DFS97, DGLR90, Erh96, GO96a, Pan96, SBG96a]. Parallele [GO96a]. Parallelism [HKP^{+}90]. Parallelization [AG90b]. parameter [Var90b]. Parameterization [GGM90]. parameters [KG93]. Parasto [Kal96]. Paris [Des96]. Park [RSS96, Ste96]. Part [JB94, Mat92a, MAA97, Tho95, YCC96a, YCC96b, Die96, KV97, Tho99, VK96]. Partial [Bet98, BD90, DG90b, GM97, HKP^{+}90, HRP90, yKaY92, MOF90, Mor97b, Ain95, Ame92, Bab95, Don96, Euv94, GV96a, GR92, GR94, HP90, KG91, KG93, KZSU91, Lan99b, LP99, MM94, QV94, QV97, Ree91, Sch91b, SBG96a, Tho95, Tho99, TGK91, TW98b]. Particles [Bun90]. partiellen [TGK91]. partieller [GR92, GR94]. partielles [Euv94]. Partitioning [TPD90]. PASCAL [KKN^{+}91b, KKN^{+}91a, KKN^{+}92, HHKR93, JR92, RS92a, RS92b]. PASCAL-XSC [KKN^{+}91b, KKN^{+}91a, KKN^{+}92, HHKR93]. past [Gol90a]. Path [STL96]. PDE [HH90, SSES90, TPD90, Ume97]. PDEs [BGGD90, BS92b]. pendulum [NL96]. People [GhW90]. perception [Gen97]. perfect [FMO96a]. Performance [DDSvdV98, FJSD96a, FJSD96b, HL95, The97, Cha98, Lyn97]. Periodic [Rei91]. periodical [Hen92]. personal [Isa90, Pea90]. Perturbation [SS90, MOS96]. perturbed [Adž96, AKM96, RST96]. PETSc [Gro97a]. phase [Gok93, Ver92]. phenomena [Ram94, SS97b]. physical [BR92]. physics [CH53, CH62, Euv94, Gar94, JJ46, JJ56, JJ99, LM97, LP97, Pan97b, Sch80, Sho97]. physique [Euv94]. pile [McV90]. pioneer [Gut90]. Plain [Mod98]. plane [Gir92]. planets [WZ94]. Planning [DDLW90]. plans [NMW96]. plasma [Tao94]. Plasticity [HR99, Hol97]. plate [BGP96]. plates [NN96]. Plovdiv [BC93, BC94, BC95]. podkhody [ZAB95]. Point [Kel93, SS97c, Sin96, DGY96, Fox90c, JP96, Shi95, SS96b]. Points [KMRT90, Pow97]. Poisson [RW96, Wer96]. Polaron [APS96]. Polish [KZSU91, MM96]. pollution [LP90]. polymeric [Whi98]. polymers [Bas96]. Polynomial [BP94b, Nem96, Sin96, Car96, EGL96, Her96, Pet96, Roj96a, Yam96]. polynomials [BV97, PTŽ96, SM91]. populations [Lam93]. posed [BCF95, BCF97, IY94, PS96a, Tik95]. possibilities [Roc97]. Posteriori [SH97, LeB93, SK96]. PostScript [GRM97]. power [Ath95]. power-stream [Ath95]. Practical [Ehr96, Eva93, Eva95, For96, CSV96, ML90, Tor97, Tor99]. practice [Asa95, CC97, Wat96]. practices [JM90]. pragmatic [SB96b]. Precise [Abe98]. Precision [CCF96]. preconditioned [Erh96]. Preconditioning [Duf97]. predicting [LP90]. Prediction [Uni99, Ano92a, Cha98, LeB93, vL97]. preferred [NMW96]. prehistory [Tod90]. presas [RRLA90]. presented [Cel90, Cro94, Eme90, Gir92, Ju95, SM90, SJ96, Sar92, Ste96, VS94]. preserving [Rei90]. presmyataniya [KV97, VK96]. pressure [Cha93a]. Preventing [Act96]. prevention [SJ96]. priblizhennogo [Deĭ94].

Priblizhennye [AIL92]. prilozheniëiĭa [ZAB95]. Primer [ST98, Sob94]. prior [Dor97]. pro [Vit94]. probabilistic [CB90]. Problem [BD90, CS99, NA90, Bro92, CB90, EZ91, GV96a, Kel96, Köc94, Lam91, LP97, Wer96, Zac96, Zac98]. Problem-Solving [BD90]. Problems [AEP90, EG91, GH93, GH95, GH97, HHKR93, Kel93, Pyt99, Rih96, SH97, Var90d, Adž96, AA96, AS93b, BCF95, BCF97, Bjo96, BCSS96, Civ98, Ded96, Fin92, Fox90c, Fri97, GZ96, Giv92, IY94, Koe96, KD92, Lae99, Tal90, LMH91, Li90, Mat92a, MOS96, Nei91, Nog97, Nov96, PS96a, RST96, Saa92, Sch90, SS97b, Sch96, Sho97, Tik95, Ume97, Var90b, VI90, VK93]. procedure [Vij95]. Procedures [Uni98d, AS93b, McV90, Mei97]. Proceedings [BC93, BC94, BC95, HMRS90, HMRS92, KvSR90, yKaY92, RRV92, Ull90b, Wes90, Ald92, AH96, AS93b, BB92, Boi97, Bru98, Civ98, DDN95, DSV94, EG91, GH94, GW90, GW92, GW94b, GHW98, Hac92, HRW94, HZO92, HRV90, KFC97, LMH91, LO90, Mar92, MK95, MK98, Mor90, NS93, Nei91, Nog97, Rak94, SD95, Smi94, SO97, VWY97]. Processes [SD95, BL94, HM96, Mik91, Olv90, Ost97, Sar92]. Processing [HRP90, KLRK98, RF96, SD95]. product [LWW96]. profile [Lyn97]. profiles [She91]. program [DFS97]. Programmed [Whe90]. Programming [Gra90, Gro90, Mor92, Roj96b, SS97c, SG98, CC98, GZ96, GMSW90, GS91a, GS91b, JR92, Lan99b, MK95, MK98, Nem96, RHP91, SC96, Zac96, Zac98]. Programs [HHKR93, RS92a, WNO94, RS92b]. progress [Ano94b]. project [Par91]. Projections [STL96]. Projects [Cra94, Gra90]. Prologue [Gol90b]. Propagation [AEP90, CB90, TC98]. properties [ZAB95]. proposed [Loz97]. Propulsion [Uni98d, Mei97]. Prototypical [BM90]. Pseudospectral [For96]. publications [Ric90]. Pullman [LSR90]. Pullman-Moscow [LSR90]. Purdue [HRV90]. Pure [LG90]. purpose [Nak97]. PVM [DHP97]. Pythagoras [Rie96].

QMR [Bas96]. quadratic

[GZ96, GMSW90]. Quadrature

[LB92, Ehr96, GN97]. quadratures [Die96]. Qualitative [IZ96]. Quality [Ano97, Boi97, VS97, Gen97, Ost97, Pan97a, Poo97]. Quasi [Kel96]. Quasi-Monte [Kel96]. quasidifferentiable [Pie91]. quasilinear [Koe96]. quench/lean [YCC96b]. Quick [YCC96b]. Quick-quench [YCC96b]. Quick-quench/lean-combustion [YCC96b].

radiative [Szi92]. Radio [Guc95]. rail [Cha93a]. Random [LWW96, Roj96a]. raschetakh [Pru93]. rate [Mer95, Mer97]. rational [BV97, Cuy94]. Rayleigh [EZ91, LM91]. Rayleigh-Benard [EZ91]. reaction [Tsu91]. reactions [FMO96a]. reactor [Bir90, Gok93]. Real [Act96, Gro97b, Mor92, Baj96, GM96, LL97, Roj96a, Roy96]. Real-time [Mor92]. realistic [Bor90]. Rechnen [GO95, GO96a, Her95]. Recipes [Cor95, PFTV92, PTV^{+}92, PTM96, PTV96, Pre93, Pre94]. recollections [Isa90]. reconstruction [Mat96]. recorded [Ste96]. Recursive [STL96, LeB93]. reduced [Rig96]. reduction [BR92]. reductions [AA96]. reference [HL96, HL98, KKN^{+}92]. refinement [Bjö90, Mat92a]. reflection [Bro92]. Reflections [Gol90b]. regional [Hud91, LP90]. regional-scale [Hud91]. regions [Cha93a]. reinforced [Avr93]. related [BCH93, FS90, PS91a, PS91b]. relationship [VY96]. relaxation [Pem92]. release [B^{+}98, Lee97]. relevant [Poo97]. Reliability [Hig97a, SJ96, Mun97]. Reliable [CH90, Bjö90, SS96a]. remarks [Rie96]. Remembrance [Gol90a].

Reminiscences [Var90c]. reordering [BHR96]. report [Ano94b, WZ94]. representation [Bir90]. Representations [BGGD90]. representing [WZ94]. Republic [GhW90, LO90, Rak94]. Research [Ano94b, Par91]. resení [Vit94]. reservoir [Pea90]. resheniia [Tik90]. resheniya [AIL92, Deĭ94, KLP90, Lae93]. residual [FM92]. residues [CM96]. Résolution [Euv94, Lam93]. resonator [Guc95]. resource [BPRB97, Bor97, Sho97, WO90]. resource-sparing [Sho97]. resources [Woo93]. responses [Naw94]. Restoration [Vel98, Blo97]. Result [AK93, Kul96]. results [Mat92a, Wan91, Nov96]. retrospection [Pea90]. return [WSG91]. revealed [LS97]. Review [Pow97, You90]. Reynolds [Lee94, Wie96]. Ricerche [NS93]. rich-burn [YCC96a]. Richardson [ACD^{+}93]. Riemann [Tor97, Tor99]. right [Gir92]. right-half [Gir92]. risk [SJ96]. River [Hib93]. różniczkowych [KZSU91]. Robert [LRRL97]. Robotic [DDLW90]. rock [PBW90, Rak94]. Rome [NS93]. roof [ADK91]. root [Vas96]. root-finding [Vas96]. rootfinding [PS96b]. roots [Car96, Her96, Pet96, Roj96a]. rotating [FS90]. rotor [Pax92, PW92, Pax95]. round [CB90]. round-off [CB90]. Rounding [Olv90]. Rousse [VWY97]. rovnic [Vit94]. równań [KZSU91]. rozwiazywania [KZSU91]. RPV [Lip94]. RPV-1 [Lip94]. Runge [NL96, Sch96]. Russian [AIL92, BV90a, Deĭ94, KLP90, Kon93, Lae93, Lae99, LG95, Pru93, Rya94, ZAB95, Zor95].

salinity [Oki96]. San [BCF95, BCF97]. satellite [GhW90]. savings [NMW96]. Scalable [Byl94]. ScaLAPACK [BCC^{+}97, DHP97]. scalar [GC95]. Scale [Noc97, Erb94, Hud91, LP90, Mu97]. Scales [BFHR97, Zei95]. scaling [VY96]. SCAN [Ull90b]. Scattered [Pow97]. schémas [Dés98]. schemes [BHR96, Dés98, Jam94a, Jam94b, Xu99]. School [Ain95, ALM99, MW86, Ste98a]. schools [Var90c]. Schrödinger [Cer93]. Schur [GMSW90]. Schur-complement [GMSW90]. schwachbesetzter [Hac91]. Schwarz [Mat92a]. science [AI98, Ano92b, Ano94b, Bet98, HSDM91, LP99, Mat92b, Poz98]. Sciences [Ald92, BP94a, BR92, Car95, Des96, Euv94, HPO92]. Scientific [Act96, AK93, Cor95, Cra94, Cra96, FJSD96a, FJSD96b, GH93, GH95, GH97, GO92b, GO93, GO95, GO96a, GO96b, GO92a, Hea97, KC93, KKN^{+}91a, LP98, MR96, MM96, Nas90b, Ort94b, Ort94a, PWF90, PFTV92, PTV^{+}92, PTM96, PTV96, Var90d, CG92, Cou97, DH95a, Dub97, Fiu95, G^{+}99, Gut90, Hat97, KC91b, KC96, KKL^{+}93, Köc94, Kul97, Mu97, Pre94, Van97, Van94, VS97, Zac96, Zac98, Her95]. Scientists [Jen77, NL95, Bor97, Ett92, Hof92, JR92, JM92]. Scilab [G^{+}99]. SDE [KPS94]. Secant [PTŽ96]. Second [BC94, Des96, yKaY92, Ort90b]. sedimentation [Tho90]. seismicheskikh [Bur93]. seismology [Bro92]. Selected [KSS95, LL97, TC98]. Selection [GL90, NA90]. Self [MDE90, Ull90b, dS97, Bun90, Ull90a]. Self-Adaptive [MDE90]. self-consistent [Bun90]. Self-Validated [dS97]. Self-Validating [Ull90b, Ull90a]. Semi [Hig90]. Semi-definite [Hig90]. semiannual [Ano94b]. Seminar [AGM93, PS90, RSS96, Hac92]. Sensitivity [SM90, Lac95]. sentiments [Fox90b]. separating [FMO96b]. Sept [HMRS90]. September [Des96, HPO92, HZO92, Rak94, RRV92, Smi94]. sequences [HM96, Met96]. SERC [Ain95]. series [Ste96]. serious [CC97]. service [Loz97, VS97]. sessions [Des96, HPO92, Nog97]. settings [BCSS96]. Seventh [Hac92, LMH91]. Several [Uni98b]. shallow [dG93, Vre94]. shallow-water [Vre94]. shape [Lac95, Rei90]. shaped [Cha93a]. shaped-rail [Cha93a]. Shaping [Blo90]. Shared [DDSv91]. Shear [Uni98b]. sheared [Tao94]. shell [RKA93]. shock [Jam94b, She91, Xu99]. short [Ree91]. Shuttle [Cha92, Cha95]. Shuzhi [WY90]. SIAM [AGM93, Blo90, RSS96]. sic [Vij95]. sided [DS90, Dob96]. sign [RR96]. signal [RF96]. silicon [Par91]. simplex [Dan90]. simplicial [Baj96]. Simulation [Eme90, KUIO90, SD95, Ton90, Uni98b, Uni98d, Bri95, CGB96, Cha92, DJT99, EZ91, Hag90, MAA97, Mei97, Mer95, Mer97, Pea90, Sar92, Tao94, Ver92, WSG91, YCC96a, YCC96b, vS99]. Simulators [BGGD90]. Simulink [CSV96]. simultaneous [Her96, Pet96, Yam96]. Sinc [LB92, Ste93]. single [Gir92]. single-link [Gir92]. singular [Gu93, MOS96]. singularities [Li90, VI90]. singularly [Adž96, RST96]. Sixteenth [Bru98]. Sixth [Ain95, GH94]. skaitliskās [Kal96]. slender [Zho96]. sliders [Cha93a]. sliding [SM91]. slope [PTŽ96]. sloped [Cou94]. slot [Ago94]. small [Wie96]. Smalltalk [PWF90]. smart [SJ96]. smog [vL97]. snap [NN96]. Society [Cel90, Eme90, Nog97, SM90, Sar92]. Sofia [DSV94]. Software [Ano97, Boi90, FFK90, GL90, HH90, HL95, HN90, MR96, The97, Mu97, Rei90, Ric92, Uni98d, Ueb97, WNO94, Bis97, Boi97, BCFH97, Byl94, BS92b, CC98, DT97, DH90, Dub97, EH97, EG92, Gen97, Hat97, Hig97b, HRV90, JM90, Köc94, Loz97, Lyn97, Mei97, MW93, Mun97, Nak91, Ost97, Poo97, Ric90, Ric93, SR97, Wan91, vS99]. Soil [Bul94, AS93b, Nog97, Wat92, WO90]. soil-structure [Nog97]. soils [Zho94]. soilware [LS97]. solids [Moh92]. Solution [Atk97, Gv97, Gro90, KC93, KP92, KPS94, MOF90, SW95, Adž96, AA96, Cox90, Deĭ94, Dab98, Euv94, Fox90c, Fra96, GR92, GR94, Hac91, Hac94, KG91, LP99, Liu93, Mei99, MM94, SS97b, Tik95, TGK91]. Solutions [DPR90, Rei91, Ded96, Elt90, Fri97, GV96a, KLP90, Koe96, KU96, Rih96, Sha96, VI90, Wal96b, Wie96]. solve [Ree91]. solved [Sch90]. Solver [SSES90]. Solvers [Gv97, TPD90, Tor97, Tor99]. Solving [Bar96a, B^{+}98, BD90, DDSv91, GH93, GH95, GH97, HKP^{+}90, Woo92, AIL92, APS96, BCF95, BCF97, Cor96, CS99, GV96a, IY94, Köc94, KZSU91, Lae93, LP97, PS96a, Pie91, Rhe98, RW96, Sch96, Sho97, Ste90, Vit94, Zac96, Zac98, Naz96]. Some [LS97, Pai90, Fox90b, Isa90, JM90, Rie96, KZSU91]. SOR [Yam96]. SOR-like [Yam96]. soundings [GhW90]. sources [Bro92, WZ94]. Space [Cha92, Cha95, Sto96]. spacecraft [FS90]. spaces [Wer96]. spacetime [LS97]. Spanish [Vol90a]. sparing [Sho97]. Sparse [BHR96, Duf97, AD90, Bas96, Erh96, GMSW90, Hac91, Hac94, Roj96a]. special [Bar96b, Des96, HPO92, Loz97]. species [FMO96a]. Specification [Uni98d, Mei97]. specimen [Ho92, Ho94]. spectral [BEM96]. spectroscopy [Lam93, PLC94]. spline [BGP96, Rei90]. splines [Baj96]. Splitflow [Uni99, Cha98]. splitting [Shu90]. sponsored [Nog97]. Sprachbeschreibung [KKN^{+}91b]. Spring [Cel90]. squared [Spä96]. squares [Bjo96, Cox90, LeB93, VY96, ZAB95, Zor95]. srodowisko [MM96]. Stability [Hig96, Rei91, Stu97, Uni99, Cha98, Naw94]. Stabilization [BFHR97]. Standards [Uni98d, Mei97, Tod90]. Stanford [LMH91]. State [AGM93, Pyt99, DW97, LP90]. static [Tes93]. Stations [WNO94]. Statistical [WO90, BL98]. statisticians [Lan99a]. Statistics [Gen98, HS93, MN99]. Statistische [BL98]. stellar [Lam93]. steps [VY96]. stiff [Pem92]. Still [vG97]. Stochastic [BRP99, Hon93, Hon94, KP92, Mar92, MK95, MK98, AA97, BL94, KD92, Mat96]. Stokes [HMRS90, HMRS92, HRW94, Sal98, Wie96]. storage [Hib93, Lee97]. Strata [Rak94]. Strategies [AG90b, Uni98a, Mod98]. stream [Ath95]. stress [FM92, RKA93, SJ96, VJ95]. strongly [Tao94]. Structural [Roc97, BS96, Ju95, Naw94]. structure [Bul94, Cou94, Jam94b, MO95, Nog97]. structures [Ano96, Avr93, MAA97, SJ96]. Student [ALM99, HL95, The97, HL97]. students [Var90a]. studies [DHP97, Lam93, Ses98]. Study [Ano96, WZ94, Chu91, Erb94, Tho90]. sub [Cha93a]. sub-ambient [Cha93a]. Subgrid [BFHR97, GC95]. Subroutine [Gee90]. Subscale [Osh94]. sum [RR96]. Summer [Ain95, ALM99, AGM93, Cro94, Ju95, RSS96, VS94, MW86, Var90c]. sunspot [Sme96]. super [Sme96]. super-long-term [Sme96]. supercomputers [dG93]. supersonic [Lee94]. supported [MAA97]. surface [Bir90, Hol97]. Survey [Hea97, Pla96, Tod62, WO90]. SUSPNDRS [MAA97]. svoistva [ZAB95]. Svremenni [KV97, VK96]. Sweden [EG91]. Switzerland [Gut90]. symbiosis [GS90]. Symbolic [Cor95, FM92, Ton90, Fiu95, RKA93]. Symmetric [AKM96]. Symmetry [AA96, Bau98, AGM93]. Symposia [Var90c]. Symposium [AH96, KvSR90]. Synthesis [KMRT90, Bor90]. System [BK90, BM90, CZ90, DG90b, GGM90, GL90, HHRS90, HKP^{+}90, HRP90, MDE90, Uni98d, FS90, GhW90, Hac91, KU96, Mei97, vS99, Boi90]. Systems [AEP90, DDSv91, DDLW90, FHI90, FFK90, Mor92, Rei91, SSES90, Stu97, AKM96, AD90, AA97, Bas96, BD96, Elt90, Hac94, Hen92, Her94, Hon93, Hon94, HRV90, KSH91, Lam91, Mei99, Mor97a, Pie91, Rhe98, Roj96a, Shu90, SH96, Wat92, Whi98, KU94].

Tahoe [Cro94, VS94]. tangential [Ago94]. tank [Lee97]. Tapia [JP96]. Tasks [HD90]. Tax [NMW96]. Tax-preferred [NMW96]. TC2 [Boi97]. TC2/WG2.5 [Boi97]. tearing [AD90, Bri95]. Technical [MK95, Mar92, MK98]. technicznych [MM96]. Technik [TGK91]. Techniques [BFHR97, ADK91, BGP96, FM92, Hac92, Lee94, Lyn97, MK95, Roc97]. technological [Des96, HPO92]. technology [Dub97, TGK91]. teorie [Vit94]. teorii [Lae99]. term [Sme96]. termination [JP96]. test [BPRB97, Loz97]. Testing [Bon93, Hig97b, Sny97, BCFH97, Mu97]. tests [Nog97]. Texas [Eme90, SM90]. TFQMR [Bas96]. Thales [Rie96]. Their [HD90, Bun90, Cer93, Hag90, Har99, Jam94a, RHP91, Sha96]. theorem [AAK96, Ded96]. theoretic [Her94]. Theoretical [BB92, MAA97, Tao94]. Theory [AEP90, HHKR93, HMRS90, HMRS92, Ise97, PT96, Sch96, SS90, AS93a, Asa95, BRP99, BS92a, BS94, Bul93, CCL99, Elt90, EG91, GM96, HR99, Hei96, Lae99, PM90, Ras94, Sal98, SD95, SS92, Vit94, Wat96, Ain95]. thermal [Ket99, LMH91]. thermally [FMO96a]. thermodydrodynamic [Vij95]. thesis [Gir92]. thin [BGP96, FS90, Par91]. things [Gol90a]. Third [BC95, Des96, EG91, Smi94, DSV94]. Thirteenth [NS93]. three [dG93]. three-dimensional [dG93]. thrust [Blo90]. Tianjin [yKaY92]. tierra [RRLA90]. Tight [Sin96]. time [KD92, Mer95, Mer97, Mor92, Nem96, Vin91, Zei95]. time-domain [Vin91]. time-scales [Zei95]. time-varying [Mer95, Mer97]. titania [Gok93]. Toeplitz [Pan96]. Toeplitz-like [Pan96]. Tome [LT94]. Tomography [Nat97]. too [Poo97]. Tool [AG90b, MAA97]. Toolbox [HHKR93, BD96, CSV96]. tools [Byl94, DT97, Kul97]. Topics [Cra96, HJ91, KSS95]. Topological [Vas96]. Toronto [Cel90]. torsion [Fri97]. torus [Tsu91]. Total [Blo97]. transfer [Eme90, Szi92, VJ95]. Transform [Ano94a]. Transforms [Wal96a]. transient [MAA97]. Transition [Rei91]. transport [Lee97, Ram94, SS97b, Szi92, Zho94]. trees [CL96]. tritium [Lee97]. turbine [Ath95, YCC96a, YCC96b]. turbomachines [Sie94]. turbulence [DJT99, GB97, Lee94]. turbulent [EZ91]. tutorial [HL96, HL98]. Twelfth [Mor90]. Two [DS90, Dob96, Kel93, Rei97, BGP96, Car96, Cha98, Fox90c, Lee94, MOS96, Rei92, Sny97]. two-equation [Lee94]. Two-Point [Kel93, Fox90c]. Two-sided [DS90, Dob96]. type [Die96, Don96, Shu90, Xu99].

U.K [BB92]. U.S. [Tod90]. U.S.A. [HRV90, LMH91]. Übungsbuch [Her98]. UCLA [HT91]. Udine [Civ98]. UK [Smi94, TC98]. unbounded [LM91]. Uncertainty [KSH91]. unconfined [Mod98]. Unconstrained [Noc97]. underflow [Hib93]. underground [Lee97]. United [Boi97, Fox90a]. Univariate [Wat97, EGL96]. Universal [MM96]. Universities [Whe90]. University [Ald92, BB92, Cel90, HRV90, yKaY92, Mar92, MK95, MK98, Mor90, Nei91, AGM93, Gee90, Ste96, Var90c]. Uniwersalne [MM96]. unsaturated [Zho94]. Uppsala [EG91]. upwind [Chu91, Jam94a]. uravnenii [AIL92, Deĭ94, KLP90, Lae93]. USA [SD95, AS93b, KFC97]. Use [FFK90, Gro90, HN90, BD96, Gro97a]. used [Nak97]. User [BCC^{+}97, HL95, The97, SSES90, HL97, Pan97a]. user-centered [Pan97a]. Using [Bar96b, CSV96, GH93, GH95, Gla93, Got95, Abe98, Bri95, Fau99, GH97, GV96a, GC95, GB97, Guc95, Köc94, Lee94, Lee97, MF99, PL95, Ses98, Sha98, Shu90, SO97, Tho90, TW98a, Van97, YCC96a, YCC96b, Zac96, Zac98]. USSR [Bab90]. Utah [Nog97, RSS96]. UV [Lam93].

V [B^{+}98, Kon93, Pru93, Rya94, Zhu92, Bar96a, MB95]. vadose [Lee97]. vagueness [KSH91]. Validated [dS97, Cor96]. Validating [Ull90b, Ull90a]. Value [Kel93, AA96, Hig97a, Koe96, Lam91]. valued [Sam96]. values [Sam96]. variation [Blo97]. Variational [Bel90, NN96]. variations [Bul93]. varying [Mer95, Mer97]. Vector [DDSv91, Bet98, Van97]. Vehicle [Cha92, Cha95, WSG91]. Verification [AK93, AS93b, Hud91, KU96, Kul96]. Verified [HHKR93]. Version [HL95, The97, HL97]. vertically [McV90]. VI [Bai98]. via [Cou97, Naz96]. vibrations [NL96]. VIII [ALM99]. virtual [Sha98]. viscoelastic [Tal90]. viscoplasticity [IS93]. Viscous [She91]. Visual [KUIO90, Ume97]. VisualDSolve [SW97]. Visualization [HL95, KUIO90, The97, Nak96]. visualizing [SW97]. Vol [BS92a, GMW91]. voln [Bur93]. Volterra [Bak97]. Volume [BP94b, McV90, Ste98b, LRRL97, Lee94, MGW96, ACD^{+}93, LT94]. Vortex [BCH93, Zho96]. VRML [Sha98]. Vvedenie [Kon93, Rya94]. vychislenii [KLP90]. vychislitelnoi [Zhu92]. vychislitelnuyu [Rya94]. vychislitelnye [Kon93]. vypukloi [Nur91].

Washington [LSR90]. Water

[Uni98a, dG93, Hud91, LSR90, Mod98, Oki96, Rak94, Vre94, Wat92, Woo93]. Wave [AEP90, Dur99, Pax92, PW92, Pax95, TC98]. Wavelets [Str96, GW99a]. Waves [Bäc99, Hag90, She91]. weather [Ano92a]. Web [BPRB97]. well [Mer95, Mer97]. WG2.5 [Boi97]. Where [Cor96]. Wilkinson [Fox90b, Gol90b, Par90a]. winners [NMW96]. Winter [Eme90, SM90, Sar92]. wissenschaftliche [GO96a]. Wissenschaftliches [GO95, Her95]. Withdrawals [Uni98a, Mod98]. WNAA [VWY97]. wood [Tes93]. Work [WNO94, Act90]. workflows [VS97]. Working [Boi97]. works [LL97]. Workshop [GH94, HRW94, TC98, VWY97, MK95, MK98, Mar92]. worry [Poo97]. Wybrane [KZSU91].

XELLPACK [BD90]. XSC [HHKR93, KKN^{+}91b, KKN^{+}91a, KKN^{+}92, KKL^{+}93, Kul97].

Years [vG97]. yinlun [WY90]. York [SD95]. York/USA/18 [SD95]. young [Lam93].

za [KV97, VK96]. zadach [Tik90]. zadachi [Lae99]. Základy [Vit94]. zero [Nov96]. zeros [Gir92, PTŽ96, Yam96]. zone [Lee97, Sha98, YCC96a]. zones [YCC96b]. zur [Her98].

References

Allgower:1996:SRN

[AA96] Eugene L. Allgower and Philip J.

Aston. Symmetry reductions for the numerical solution of boundary value problems. In Renegar et al. [RSS96], pages 29–48. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Artemiev:1997:NAS

[AA97] S. S. Artemiev and T. A. Averina. Numerical analysis of systems of ordinary and stochastic differential equations. VSP, International Science Publishers, Utrecht, The Netherlands, 1997. ISBN 90-6764-250-9. vi + 176 pp. LCCN QA371 .A96 1997.

Atanassova:1996:CIM

[AAK96] L. Atanassova, A. Andreev, and N. Kjurkchiev. Combined iterative methods based on a modification of the Koenig theorem. In Alefeld and Herzberger [AH96], pages 23–37. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Aberth:1998:PNM

[Abe98] Oliver Aberth. Precise numerical methods using C++. Academic Press, New York, NY, USA, 1998. ISBN 0-12-041750-2. xiv + 238 pp. LCCN QA76.73.C153 A32 1998.

URL http://www.loc.gov/
catdir/description/els033/
98117803.html;   http://www.
loc.gov/catdir/toc/els032/
98117803.html. With 1 CD-

ROM (Windows 95).

Ashford:1993:CPL

[ACD^{+}93] Oliver M. Ashford, H. Charnock, P. G. Drazin, J. C. R. Hunt, P. Smoker, and Ian Sutherland, editors. Collected Papers of Lewis Fry Richardson. Volume 1: Meteorology and Numerical Analysis. Cambridge University Press, ????, 1993. ISBN 0-521-38297-1. xiv + 1016 pp.

Acton:1990:NMW

[Act90] Forman S. Acton. Numerical methods that work. Mathematical Association of America, Washington, DC, USA, 1990. ISBN 0-88385-450-3. xx + 549 pp. LCCN QA297 .A33 1990.

URL http://www.loc.gov/
catdir/description/cam024/
90062538.html;   http://www.
loc.gov/catdir/toc/cam027/
90062538.html. Cover title:

Numerical methods that usually work. Updated and revised from the 1970 edition published by Harper and Row.

Acton:1996:RCM

[Act96] Forman S. Acton. Real Computing Made Real: Preventing Errors in Scientific and Engineering Calculations. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-03663-2. xv + 259 pp.

Arioli:1990:ETL

[AD90] Mario Arioli and Iain S. Duff.

Experiments in tearing large sparse systems. In Cox and Hammarling [CH90], pages 207– 226. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Ahola:1991:ANM

[ADK91] M. P. Ahola, D. A. Donato, and Nicholas P. Kripakov. Application of numerical modeling techniques to analysis of cutter roof failure. U.S. Dept. of the Interior, Bureau of Mines, Washington, DC, USA, 1991. ii + 28 pp.

Adzic:1996:LDA

[Adž96] Nevenka Adžić. A low-degree approximation for layer solution of singularly perturbed problems. In Alefeld and Herzberger [AH96], pages 9–15. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Alekseev:1990:ESI

[AEP90] A. S. Alekseev, G. N. Erokhin, and N. L. Podkolodny. Expert systems for inverse problems of wave propagation theory. In Houstis et al. [HRV90], pages 43–50. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Allgower:1990:NCM

[AG90a] E. L. (Eugene L.) Allgower and Kurt Georg. Numerical continuation methods: an introduction, volume 13 of Springer series in computational mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 3-540-12760-7, 0-387-12760-7 (New York). xiv + 388 pp. LCCN QA377 .A562 1990.

Anand:1990:TEC

[AG90b] T. S. Anand and R. Gupta. A tool for evaluating compilerbased parallelization strategies. In Houstis et al. [HRV90], pages 103–110. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Allgower:1993:ESA

[AGM93] E. L. Allgower, Kurt Georg, and Rick Miranda, editors. Exploiting symmetry in applied and numerical analysis: 1992 AMS-SIAM Summer Seminar in Applied Mathematics, July 26– August 1, 1992, Colorado State University, volume 29 of Lectures in applied mathematics. American Mathematical Society, Providence, RI, USA, 1993. ISBN 0-8218-1134-7. LCCN QA297 .S846 1992.

Agosta:1994:NAT

[Ago94] Roxana M. Agosta. Numerical analysis of tangential slot blowing on a generic chined forebody. National Technical Information Service, Washington, DC, USA, 1994. ???? pp. Shipping list no.: 95-0053-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1994] 1 microfiche.

Alefeld:1996:NME

[AH96] Götz Alefeld and Jürgen Herzberger, editors. Numerical methods and error bounds: proceedings of the IMACS GAMM International Symposium on Numerical Methods and Error Bounds held in Oldenburg, Germany, July 9–12, 1995, volume 89 of Mathematical Research. Akademie-Verlag, Berlin, Germany, 1996. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Allen:1998:NAA

[AI98] Myron B. Allen, III and Eli L. Isaacson. Numerical analysis for applied science. Pure and Applied Mathematics (New York). John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1998. ISBN 0-471-55266-6. xi + 492 pp. LCCN QA297 .A53 1998 Bar. URL http:

//www.loc.gov/catdir/bios/
wiley042/97016688.html;
http://www.loc.gov/catdir/
description/wiley033/97016688.
html;   http://www.loc.gov/
catdir/toc/wiley022/97016688.

html. A Wiley-Interscience publication.

Agafonov:1992:PMR

[AIL92] S. A. Agafonov, E. E. Ivanova, and A. A. Levitskiĭ. Priblizhennye metody resheniya obyknovennykh differentsialnykh uravnenii. (Russian) [Approximate methods for solving ordinary differential equations]. Moskovskiĭ Gosudarstvennyĭ Tekhnicheskiĭ Universitet im. N. È. Baumana, Moscow, Russia, 1992. ISBN 5-7038-0699-2. 95 pp.

Ainsworth:1995:TNO

[Ain95] Mark Ainsworth, editor. Theory and numerics of ordinary and partial differential equations: [lecture notes of the Sixth SERC Summer School in Numerical Analysis, held at Leicester from the 18th to the 29th of July, 1994], volume 4 of Oxford science publications; Advances in numerical analysis. Clarendon Press, New York, NY, USA, 1995. ISBN 0-19-851193-0. LCCN ????

Adams:1993:SCA

[AK93] E. Adams and U. Kulisch. Scientific Computing with Automatic Result Verification, volume 189 of Mathematics in Science and Engineering. Academic Press, New York, NY, USA, 1993. ISBN 0-12-044210-8. x + 612 pp. LCCN QA76 .S368 1993.

Akai:1994:ANM

[Aka94] Terrence J. Akai. Applied numerical methods for engineers. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1994. ISBN 0-471-57523-2. xix + 410 pp. LCCN TA335 .A38 1994.

Akin:1994:FEA

[Aki94] J. E. Akin. Finite elements for analysis and design. Computational Mathematics and Applications. Academic Press, New York, NY, USA, 1994. ISBN 0-12-047654-1. xii + 548 pp. With a contribution by Bala Ramaswamy, With 1 IBM-PC floppy disk (3.5 inch; HD), UNIX version of MODEL available separately.

Alefeld:1996:SLS

[AKM96] G. Alefeld, V. Kreinovich, and G. Mayer. Symmetric linear systems with perturbed input data. In Alefeld and Herzberger [AH96], pages 16–22. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Alder:1992:NME

[Ald92] H. Alder. Numerical Methods in Engineering and Applied Sciences = Metodos Numericos en Ingenieria y Cincias Aplicadas: proceedings of the International Congress on Numerical Methods in Engineering and Applied Sciences, held at University of Concepcion, Chile, 16–20 November, 1992. Centro Internacional de Metodos Numericos en Ingenieria, Barcelona, Spain, 1992. ISBN 84-87867-16-2. 613 (vol. 1), 633–1340 (vol. 2) pp. LCCN 555 I575 1992.

Ainsworth:1999:GSG

[ALM99] Mark Ainsworth, Jeremy Levesley, and Marco Marletta, editors. The Graduate Student’s Guide to Numerical Analysis ’98: Lecture Notes from the VIII EPSRC Summer School in Numerical Analysis, volume 26 of Springer series in computational mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-65752-5 (hardcover). ISSN 0179-3632. x + 250 pp. LCCN QA 297 G67 1999. URL

http://www.loc.gov/catdir/
enhancements/fy0815/99032314-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0815/
99032314-t.html.  Lecture

notes from the 8th EPSRC Numerical Analysis Summer School held at the University of Leicester, Leicester, July 5–17, 1998.

Ayyub:1996:NME

[AM96] Bilal M. Ayyub and Richard H.

McCuen. Numerical methods for engineers. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-337361-4. xix + 362 pp. LCCN TA335 .A99 1996.

Ames:1992:NMP

[Ame92] William F. Ames. Numerical methods for partial differential equations. Computer Science and Scientific Computing. Academic Press, New York, NY, USA, third edition, 1992. ISBN 0-12-056761-X. ix + 451 pp. LCCN QA374 .A46 1992.

URL http://www.loc.gov/
catdir/description/els032/
91032811.html;   http://www.
loc.gov/catdir/toc/els032/
91032811.html.

Anonymous:1992:MIN

[Ano92a] Anonymous. Methods of interpreting numerical weather prediction output for aeronautical meteorology. Number 770 in WMO. Secretariat of the World Meteorological Organization, Geneva, Switzerland, 1992. ISBN 92-63-10770-X. v + 89 pp. LCCN QC851.W6 no.195. World Meteorology Organization. Working Group on Advanced Techniques Applied to Aeronautical Meteorology.

Anonymous:1992:NMA

[Ano92b] Anonymous, editor. Numerical methods in applied science and industry, Rendiconti del Seminario matematico. Universita e politecnico, Torino, Italy, 1992. ISSN 0373-1243. LCCN TA331 .N86 1991. Papers presented at the International Conference on Numerical Methods in Applied Science and Industry held at the Politecnico di Torino, June 18–21, 1990. Fascicolo speciale, 1991.

Anonymous:1994:FFF

[Ano94a]  Anonymous.  Fundamentals

of the Fast Fourier Transform (FFT). Technical Education Videos, Redmond, WA, USA, 1994. ISBN 0-9642912-0-7. 1 videocassette (NTSC; 1/2 inch; VHS) (144 min.).

Anonymous:1994:RPA

[Ano94b]  Anonymous.  Research in

progress in applied mathematics, numerical analysis, fluid mechanics, and computer science: semiannual report, October 1, 1993 through March 31, 1994. National Technical Information Service, Washington, DC, USA, 1994. ???? pp. Institute for Computer Applications in Science and Engineering. Shipping list no.: 95-0051-M.

Anonymous:1996:SNA

[Ano96] Anonymous. Study on numerical analysis of structures. Technical Report 94k3-1402-06-01-3.

1996.2, Advanced Structure Research Station, Hanyang University, South Korea, 1996. xi + 111 pp. In Korean, abstract in English. Title also in Korean. Hanyang Taehakkyo. Advanced Structure Research Station.

Anonymous:1997:QNS

[Ano97] Anonymous. The quality of numerical software: Assessment and enhancement. In Boisvert [Boi97], pages 375–380. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Atteia:1990:EAN

[AP90] M. (Marc) Atteia and M. (Michel)

Pradel. Éléments d’analyse numérique. (French) [Elements of numerical analysis]. Cépaduès Éditions, Toulouse, France, 1990. ISBN 2-85428-229-9, 2-85428-299-9. viii + 154 pp. LCCN QA297 .A88 1990. 90F.

Akishin:1996:NSN

[APS96] Pavel G. Akishin, Igor V.

Puzynin, and Yuri S. Smirnov. On numerical solving of nonlinear Polaron equations. In Renegar et al. [RSS96], pages 19–28. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Armero:1997:FNA

[Arm97]  Francisco Armero.  Formu-

lation and numerical analysis of an anisotropic damage model with a localized dissipative mechanism. Technical Report UCB/SEMM-97/11, Dept. of Civil Engineering, University of California, Berkeley, Berkeley, CA, USA, August 1997. 44 pp.

Argyros:1993:TAI

[AS93a] Ioannis K. Argyros and Ferenc Szidarovszky. The theory and applications of iteration methods. Systems Engineering Series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1993. ISBN 0-8493-8014-6. xii + 355 pp. LCCN QA297.8 .A74 1993. URL

http://www.loc.gov/catdir/
enhancements/fy0744/93007173-
d.html.

Arulanandan:1993:VNP

[AS93b] Kandiah Arulanandan and Ronald F. Scott. Verification of numerical procedures for the analysis of soil liquefaction problems: proceedings of the International Conference on the Verification of Numerical Procedures for the Analysis of Soil Liquifaction Problems, Davis, California, USA, 17–20 October 1993. A. A. Balkema, Brookfield, VT, USA, 1993. ISBN 90-5410-360-4 (set), 90-5410-361-2 (vol. 1), 90-5410-362-0 (vol. 2). xiv + 1801 pp. LCCN TA710 .V47 1993 v.1-2 (1993-1994). The VELACS (VErification of Liquefaction Analysis by Centrifuge Studies) Project is sponsored by the National Science Foundation and involves the cooperative effort of seven universities. Contents: vol. 1. Experimental results and numerical predictions (Class A and B) — vol. 2. [without special title].

Asaithambi:1995:NAT

[Asa95] N. S. Asaithambi. Numerical analysis: theory and practice. Saunders College Pub., Fort Worth, TX, USA, 1995. ISBN 0-03-030983-2. xiv + 855 + 17 + 11 pp. LCCN QA297 .A747 1995.

Athavale:1995:NAI

[Ath95]  M. M. Athavale.  Numeri-

cal analysis of intra-cavity and power-stream flow interaction in multiple gas-turbine diskcavities. National Technical Information Service, Washington, DC, USA, 1995. ???? pp. Shipping list no.: 95-0744-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1995] 1 microfiche.

Atkinson:1993:ENA

[Atk93] Kendall E. Atkinson. Elementary numerical analysis. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, second edition, 1993. ISBN 0-471-50999-X. xiii + 425 pp. LCCN QA297 .A831 1993.

Atkinson:1997:NSB

[Atk97] Kendall E. Atkinson. The numerical solution of boundary integral equations. In Duff and Watson [DW97], pages 223–260. ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

//global.oup.com/academic/
product/the-state-of-the-
art-in-numerical-analysis-
9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Avram:1993:NAR

[Avr93] Constantin Avram. Numerical analysis of reinforced concrete structures, volume 41 of Developments in civil engineering. Elsevier, Amsterdam, The Netherlands, 1993. ISBN 0-444-98842-4. xi + 508 pp. LCCN TA683.2 .S7713 1993. Rev., extended, and updated translation of: Structuri din beton armat.

Barrow:1998:SDE

[B^{+}98] David Barrow et al. Solving differential equations with Maple V, release 4. Brooks/Cole symbolic computation series. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1998. ISBN 0-534-34555-7. viii + 253 pp. LCCN QA371.5.D37S65 1998.

Babuska:1990:DCM

[Bab90] I. Babuska. The development of computational mathematics in Czechoslovakia and the USSR. In Nash [Nas90b], pages 314– 328. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Babuska:1995:MMG

[Bab95] Ivo Babuska. Modeling, mesh generation, and adaptive numerical methods for partial differential equations, volume 75 of The IMA volumes in mathematics and its applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94542-3 (New York). li + 450 pp. LCCN QA377 .M578 1995. Errata slip inserted. Based on the proceedings of the 1993 IMA summer program.

Backstrom:1999:WFE

[Bäc99] Gunnar Bäckström. Waves by finite element analysis. Studentlitteratur, Lund, Sweden, 1999. ISBN 91-44-01007-9. viii + 189 pp.

Baines:1998:NMF

[Bai98] M. J. Baines, editor. Numerical methods for fluid dynamics VI. ICFD, Oxford, 1998. ISBN 0-9524929-1-1 (paperback). LCCN TA357.N8729 1998. Proceedings of a conference held at St Catherine’s College and the St Cross Building, March/April 1998.

Bajaj:1996:CRA

[Baj96] Chandrajit L. Bajaj. The combinatorics of real algebraic splines over a simplicial complex. In Renegar et al. [RSS96], pages 49–58. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Baker:1997:NAV

[Bak97] Christopher T. H. Baker. Numerical analysis of Volterra functional and integral equations. In Duff and Watson [DW97], pages 193–222.

ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

                                        //global.oup.com/academic/
                                        product/the-state-of-the-
                                        art-in-numerical-analysis-
                                        9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Barrow:1996:SOM

[Bar96a] David Barrow. Solving ODEs with Maple V. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, preliminary edition, 1996. ISBN 0-534-34402-X. viii + 154 pp. LCCN ????

Barth:1996:USA

[Bar96b] W. Barth. Using special arithmetic in geometric algorithms. In Alefeld and Herzberger [AH96], pages 38–47. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Basermann:1996:QTM

[Bas96] Achim Basermann. QMR and TFQMR methods for sparse nonsymmetric polymers on massively parallel systems. In Renegar et al. [RSS96], pages 59–76. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Baumann:1998:SAD

[Bau98] Gerd Baumann. Symmetry analysis of differential equations with Mathematica. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1998. ISBN 0-387-98552-2 (hardcover). xii + 521 pp. LCCN QA371.B36 1998.

Beevers:1992:TCN

[BB92] C. J. Beevers and A. F. Blom.

Theoretical concepts and numerical analysis of fatigue: proceedings of a conference held 25–27th May 1992 at The University of Birmingham, U.K. EMAS, Warley. UK, 1992. ISBN 0-947817-57-3. vi + 589 pp. LCCN TA418.38.T44 1992.

Biran:1999:ME

[BB99] Adrian Biran and Moshe Breiner.

MATLAB 5 for Engineers. Addison-Wesley, Reading, MA, USA, second edition, 1999. ISBN 0-201-36043-8. xvi + 688 pp. LCCN QA297.B52 1999.

URL http://www.awl-he.com/
titles/11606.html. Rev. ed.

of: MATLAB for engineers (1995).

Bevilacqua:1992:MN

[BBCM92] Roberto Bevilacqua, D. Bini, M. Capovani, and O. Menchi. Metodi numerici. (Italian) [Numerical methods]. Zanichelli, Bologna, Italy, 1992. ISBN 88-08-14510-7. x + 862 pp. LCCN ????

Bainov:1993:PFI

[BC93] D. Bainov and Valery Covachev, editors. Proceedings of the First International Colloquium on Numerical Analysis, Plovdiv, Bulgaria, 13–17 August 1992. VSP, International Science Publishers, Utrecht, The Netherlands, 1993. ISBN 90-6764-152-9. LCCN QA297.I4518 1992.

Bainov:1994:PSI

[BC94] D. Bainov and Valery Covachev, editors. Proceedings of the Second International Colloquium on Numerical Analysis: Plovdiv, Bulgaria, 13–17 August 1993. VSP, International Science Publishers, Utrecht, The Netherlands, 1994. ISBN 90-6764-165-0. LCCN QA297.I4518 1993.

Bainov:1995:PTI

[BC95] D. Bainov and Valery Covachev, editors. Proceedings of the Third International Colloquium on Numerical Analysis: Plovdiv, Bulgaria, 13–17 August 1994. VSP, International Science Publishers, Utrecht, The Netherlands, 1995. ISBN 90-6764-193-6. LCCN QA297.I45 1994.

Blackford:1997:SUG

[BCC^{+}97] L. S. Blackford, J. Choi, A. Cleary, E. D’Azevedo, J. Demmel, I. Dhillon, J. Dongarra, S. Hammarling, G. Henry, A. Petitet, K. Stanley, D. Walker, and R. C. Whaley. ScaLA-PACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1997. ISBN 0-89871-400-1 (paperback), 0-89871-401-X (CD-ROM), 0-89871-397-8 (set). xxvi + 325 pp. LCCN QA185.D37 S33 1997.

Barbour:1995:ENM

[BCF95] Randall L. Barbour, Mark Joseph Carvlin, and M. A. Fiddy, editors. Experimental and numerical methods for solving ill-posed inverse problems: medical and nonmedical applications, 10–11 July 1995, San Diego, California, volume 2570 of Proceedings / SPIE–the International Society for Optical Engineering. SPIE Optical Engineering Press, Bellingham, WA, USA, 1995. ISBN 0-8194-1929-X. LCCN QC350 .E96 1995.

Barbour:1997:CEN

[BCF97] Randall L. Barbour, Mark Joseph Carvlin, and M. A. Fiddy. Computational, experimental, and numerical methods for solving ill-posed inverse imaging problems: medical and nonmedical applications: 30–31 July 1997, San Diego, California, volume 3171 of SPIE proceedings series. SPIE Optical Engineering Press, Bellingham, WA, USA, 1997. ISBN 0-8194-2593-1. ix + 240 pp. LCCN RC78.7.D53 C66 1997.

Butler:1997:MTC

[BCFH97] B. Butler, M. Cox, A. Forbes, and S. Hannaby. A methodology for testing classes of approximation and optimisation software. In Boisvert [Boi97], pages 138– 151. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Beale:1993:VFR

[BCH93] J. T. Beale, G.-H. Cottet, and S. Huberson, editors. Vortex flows and related numerical methods, volume 395 of NATO ASI series. Series C, Mathematical and physical sciences. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1993. ISBN 0-7923-2250-9. LCCN QA925 .V66 1993. Published in cooperation with NATO Scientific Affairs Division. Proceedings of the NATO Advanced Research Workshop on Vortex Flows and Related Numerical Methods, Grenoble, France, June 15–19, 1992.

Brandal:1991:NLA

[BCS91] Willy Brandal, Charles O. Christenson, and Bryan A. Smith, editors. Numerical linear algebra. BCS Associates, Moscow, ID, USA, 1991. ISBN 0-914351-05-2. vi + 202 pp. LCCN QA184 .B72 1991.

Blum:1996:ASP

[BCSS96] Lenore Blum, Felipe Cucker, Mike Shub, and Steve Smale. Algebraic settings for the problems “P not equal to NP?”. In Renegar et al. [RSS96], pages 125– 144. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Bonomo:1990:XIP

[BD90] J. P. Bonomo and W. R. Dyksen. XELLPACK: An interactive problem-solving environment for elliptic partial differential equations. In Houstis et al. [HRV90], pages 331–342. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Beale:1996:FST

[BD96] Mark H. Beale and Howard B.

Demuth. Fuzzy systems toolbox for use with MATLAB. PWS productivity tools. PWS-Kent Publishing Company, Division of Wadsworth, Inc., 20 Park Plaza, Boston, MA 02116, USA, student, Mac version edition, 1996. ISBN 0-534-95258-5 (disk), 0-534-94579-1 (manual). 1 computer disk pp. LCCN QA297.

Beltzer:1990:VFE

[Bel90] A. I. Beltzer. Variational and finite element methods. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 3-540-51598-4. xii + 254 pp. A symbolic computation approach.

Bellman:1997:IMA

[Bel97] Richard Bellman. Introduction to matrix analysis, volume 19 of Classics in Applied Mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1997. ISBN 0-89871-399-4. xxviii + 403 pp. LCCN QA188 .B445 1997. Reprint of the second (1970) edition, with a foreword by Gene Golub.

Brown:1996:NAC

[BEM96] B. M. Brown, M. S. P. Eastham, and D. K. R. McCormack. A new algorithm for computing the spectral matrix for higherorder differential equations and the location of discrete eigenvalues. In Renegar et al. [RSS96], pages 145–156. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Betounes:1998:PDE

[Bet98] David Betounes. Partial differential equations for computational science: with Maple and vector analysis. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1998. ISBN 0-387-98300-7 (hardcover). xix + 517 pp. LCCN QA377.B53 1998.

Burden:1993:NA

[BF93] Richard L. Burden and J. Douglas Faires. Numerical analysis. Prindle, Weber and Schmidt series in mathematics. PWS-Kent Pub. Co., Boston, MA, USA, fifth edition, 1993. ISBN 0-534-93219-3. xiv + 768 pp. LCCN QA297 .B84 1993.

Burden:1997:NA

[BF97] Richard L. Burden and J. Douglas Faires. Numerical analysis. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, sixth edition, 1997. ISBN 0-534-95532-0. xiii + 811 pp. LCCN QA 297 B84 1997.

Brezzi:1997:STS

[BFHR97] F. Brezzi, L. P. Franca, T. J. R. Hughes, and A. Russo. Stabilization techniques and subgrid scales capturing. In Duff and Watson [DW97], pages 391– 406. ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

//global.oup.com/academic/
product/the-state-of-the-
art-in-numerical-analysis-
9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Balaban:1990:KRA

[BGGD90] D. Balaban, J. Garbarini, W. Greiman, and M. Durst. Knowledge representations for the automatic generation of numerical simulators for PDEs. In Houstis et al. [HRV90], pages 269–280. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Beatson:1996:MTT

[BGP96] R. K. Beatson, G. Goodsell, and M. J. D. Powell. On multigrid techniques for thin plate spline interpolation in two dimensions. In Renegar et al. [RSS96], pages 77–98. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Bruce:1990:MM

[BGR90] J. W. (James William) Bruce, P. J. Giblin, and P. J. Rippon. Microcomputers and mathematics. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-37515-0, 0-521-31238-8 (paperback). xvi + 425 pp. LCCN QA76.95 .B78 1989.

                                        URL http://www.loc.gov/
                                        catdir/description/cam024/
                                        90031051.html;   http://www.
                                        loc.gov/catdir/toc/cam027/
                                        90031051.html.

Bhatia:1997:MA

[Bha97] Rajendra Bhatia. Matrix Analysis. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-94846-5. xi + 347 pp.

Berry:1996:SMR

[BHR96] Michael W. Berry, Bruce Hendrickson, and Padma Raghavan. Sparse matrix reordering schemes for browsing hypertext. In Renegar et al. [RSS96], pages 99–124. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Birkhoff:1990:FDR

[Bir90] Garrett Birkhoff. Fluid dynamics, reactor computations, and surface representation. In Nash [Nas90b], pages 63–87. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Bischof:1997:ADN

[Bis97] C. H. Bischof. Automatic differentiation and numerical software design. In Boisvert [Boi97], pages 287–299. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Brand:1992:NCM

[BJ92] Ludwig Brand and Michael L. Johnson. Numerical computer methods, volume 210 of Methods in enzymology. Academic Press, New York, NY, USA, 1992. ISBN 0-12-182111-0. xxix + 718 pp. LCCN QU135 .C58 v. 210 and 240 1992–1994.

Breitschuh:1993:FEM

[BJ93] Ulrich Breitschuh and Ronald Jurisch. Die Finite-Element-Methode. (German) [The Finite-Element Method]. Akademie-Verlag, Berlin, Germany, 1993. ISBN 3-05-501571-1. 270 pp. Theorie, Praxis und Programmierung auf Personalcomputern. [Theory, practice and programming on personal computers].

Bjorck:1990:IRR

[Bjö90] Åke Björck. Iterative refinement and reliable computing. In Cox and Hammarling [CH90], pages 249–266. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Bjorck:1996:NML

[Bjo96] Ake Bjorck. Numerical methods for least squares problems. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1996. ISBN 0-89871-360-9 (paperback). xvii + 408 pp. LCCN QA 214 B56 1996.

Barnett:1990:EES

[BK90] D. Barnett and D. Kahaner. Experiences with an expert system for ODEs. In Houstis et al. [HRV90], pages 5–14. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Bouleau:1994:NMS

[BL94] Nicolas Bouleau and Dominique Lepingle. Numerical methods for stochastic processes. Wiley series in probability and mathematical statistics. Applied probability and statistics. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1994. ISBN 0-471-54641-0. xvii + 359 pp. LCCN QA274 .B67 1994. A Wiley-Interscience publication.

Blobel:1998:SNM

[BL98] Volker Blobel and Erich Lohrmann. Statistische und numerische Methoden der Datenanalyse (German) [Statistical and numerical methods of data analysis]. Teubner Studienbücher Physik. [Teubner Textbooks in Physics]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1998. ISBN 3-519-03243-0. 358 pp.

Block:1990:SEN

[Blo90] I. Edward Block. Shaping the evolution of numerical analysis in the computer age: the SIAM thrust. In Nash [Nas90b], pages 199–205. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Blomgren:1997:TVI

[Blo97] Peter Blomgren. Total variation image restoration: numerical methods and extensions. Dept. of Mathematics, University of California, Los Angeles, Los Angeles, CA, USA, October 1997. [4] pp. To appear in the proceedings of ICIP 1997.

Berkbigler:1990:GAP

[BM90] K. P. Berkbigler and P. A. Max.

The graphics advisor: a prototypical advisory expert system. In Houstis et al. [HRV90], pages 187–196. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Boisvert:1990:GAM

[Boi90] R. F. Boisvert. The Guide to Available Mathematical Software Advisory System. In Houstis et al. [HRV90], pages 167–178. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Boisvert:1997:QNS

[Boi97] Ronald F. Boisvert, editor.

Quality of numerical software: assessment and enhancement / proceedings of the IFIP TC2/WG2.5 Working Conference on the Quality of Numerical Software, Assessment and Enhancement, Oxford, United Kingdom, 8–12 July 1996. Chapman and Hall, Ltd., London, UK, 1997. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Bonnier:1993:TMN

[Bon93] Paul Godfried Bonnier. Testing, modelling and numerical analysis of the mechanical behaviour of bituminous concrete. Thesis (ph.d.), Technische Universiteit Delft, Delft, The Netherlands, 1993. 157 pp. Includes Dutch abstract.

Borges:1990:NMI

[Bor90] Carlos Freitas Borges. Numerical methods for illumination models in realistic image synthesis. Thesis (ph.d.) in applied mathematics, University of California, Davis, Davis, CA, USA, 1990. 107 pp.

Borse:1991:FNM

[Bor91] Garold J. (Garold J.) Borse. Fortran 77 and Numerical Methods for Engineers. PWS-Kent series in engineering and computer science. PWS-Kent Pub. Co., Boston, MA, USA, second edition, May 1991. ISBN 0-534-92562-6. xiv + 722 pp. LCCN TA 345 B67 1991. US$57.95. URL http:/

/www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0-534-92562-6.

Borse:1997:NMM

[Bor97] G. J. (Garold J.) Borse. Numerical methods with MATLAB: a resource for scientists and engineers. PWS-Kent Pub. Co., Boston, MA, USA, 1997. ISBN 0-534-93822-1. xvi + 640 pp. LCCN QA297 .B647 1997.

Boehm:1993:NM

[BP93] Wolfgang Boehm and Hartmut Prautzsch. Numerical methods. Friedrich Vieweg und Sohn and A. K. Peters, Ltd., Braunschweig, Germany and Wellesley, MA, USA, 1993. ISBN 3-528-06350-5, 1-56881-020-2. x + 186 pp. LCCN QA297 .B58 1993.

Berman:1994:NMM

[BP94a] Abraham Berman and Robert J.

Plemmons. Nonnegative Matrices in the Mathematical Sciences, volume 9 of Classics in Applied Mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-321-8. xx + 340 pp. LCCN QA188 .B47 1994. URL

http://www.loc.gov/catdir/
enhancements/fy0708/94037449-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0708/
94037449-t.html.  Revised

reprint of [?].

Bini:1994:PMC

[BP94b] Dario Bini and Victor Y. Pan.

Polynomial and Matrix Computations. Volume 1: Fundamental Algorithms. Birkhäuser Verlag, Basel, Switzerland, 1994. ISBN 0-8176-3786-9. xvi + 415 pp.

Boehm:1994:GCG

[BP94c] Wolfgang Boehm and Hartmut Prautzsch. Geometric concepts for geometric design. A. K. Peters, Ltd., Wellesley, MA, USA, 1994. ISBN 1-56881-004-0. xvii + 393 pp. LCCN QA445 .B63 1994b. US$49.95.

Boisvert:1997:MMW

[BPRB97] R. F. Boisvert, R. Pozo, K. Remington, and R. F. Barrett. Matrix Market: a Web resource for test matrix collections. In Boisvert [Boi97], pages 125–137. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Bevington:1992:DRE

[BR92] Philip R. Bevington and D. Keith Robinson. Data reduction and error analysis for the physical sciences. McGraw-Hill, New York, NY, USA, second edition, 1992. ISBN 0-07-911243-9. xvii + 328 pp. LCCN QA278 .B48 1992. US$14.95 (set only). With 1 IBM-PC floppy disk (5.25 inch; HD).

Brezinski:1992:BAN

[Bre92] Claude Brezinski. Biorthogonality and its applications to numerical analysis. Marcel Dekker, Inc., New York, NY, USA, 1992. ISBN 0-8247-8616-5. v + 166 pp. LCCN QA404.5 .B72 1992.

Brittnacher:1995:LAT

[Bri95] Mitchell J. Brittnacher. A linear analysis of tearing instability in magnetotail equilibria using simulation and numerical methods. Thesis (ph.d.), University of California, San Diego, San Diego, CA, USA, 1995. xvi + 135 pp.

Brookes:1992:CNM

[Bro92]  Robert W. Brookes.  Con-

vergence of numerical methods for the one-dimensional inverse problem of reflection seismology with general input sources. Thesis (ph.d.), University of California, Los Angeles, Los Angeles, CA, USA, 1992. ix + 167 pp.

Bardi:1999:SDG

[BRP99] M. Bardi, T. E. S. Raghavan, and T. Parthasarathy. Stochastic and differential games: theory and numerical methods, volume 4 of Annals of the International Society of Dynamic Games. Birkhäuser Verlag, Basel, Switzerland, 1999. ISBN 0-8176-4029-0 (hardcover). xvi + 380 pp. LCCN QA272 .S76 1999.

Bruneau:1998:SIC

[Bru98] Charles-Henri Bruneau. Sixteenth International Conference on Numerical Methods in Fluid Dynamics: proceedings of the conference held in Arcachon, France, 6–10 July 1998, volume 515 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-65153-5. ISSN 0075-8450 (print), 1616-6361 (electronic). xv + 568 pp. LCCN QA911 .I58 1998.

Braess:1992:NMA

[BS92a] Dietrich Braess and Larry L.

Schumaker. Numerical methods in approximation theory.

Vol. 9, volume 105 of International series of numerical mathematics. Birkhäuser Verlag, Basel, Switzerland, 1992. ISBN 3-7643-2746-4 (Basel), 0-8176-2746-4 (Boston). xiv + 357 pp. LCCN QA221 .N84 1992. Official proceedings of a conference on ‘Numerical methods in approximation theory’ which was held at the Mathematisches Forschungsinstitut in Oberwolfach during the week of November 24–30, 1991.

Byrne:1992:RDN

[BS92b] George D. Byrne and W. E. Schiesser, editors. Recent developments in numerical methods and software for ODEs/DAEs/ PDEs. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1992. ISBN 981-02-0557-0. LCCN QA370 .R43 1992. Papers presented at the American Institute of Chemical Enginees Annual Meeting in Chicago, Nov. 11–16, 1990.

Brenner:1994:MTF

[BS94] Susanne C. Brenner and L. Ridgway Scott. The mathematical theory of finite element methods, volume 15 of Texts in Applied Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-94193-2 (New York), 3-540-94193-2 (Berlin). xii + 294 pp. LCCN QA379 .B74

Bittnar:1996:NMS

[BS96] Zdenek Bittnar and Jiri Sejnoha.

Numerical methods in structural mechanics. Thomas Telford, London, 1996. ISBN 0-7844-0170-5. 422 pp. LCCN TA645 .B59 1996.

Buchanan:1992:NMA

[BT92] James L. Buchanan and P. R.

Turner. Numerical methods and analysis. International series in pure and applied mathematics. McGraw-Hill, New York, NY, USA, 1992. ISBN 0-07-008717-2. xv + 751 pp. LCCN QA297.B79 1992.

Bulirsch:1993:OCC

[Bul93] Roland Bulirsch. Optimal control: calculus of variations, optimal control theory, and numerical methods, volume 111 of International series of numerical mathematics. Birkhäuser Verlag, Basel, Switzerland, 1993. ISBN 3-7643-2887-8 (Basel), 0-8176-2887-8 (Boston). xiii + 349 pp. LCCN QA315 .O68 1993. Based on a conference optimal control-variationsrechnung und optimalsteuerungen which was held at the Mathematisches Forschungsinstitut of the University of Freiburg during the period May 23 to June 1, 1991.

Bull:1994:SSI

[Bul94] John W. Bull. Soil structure interaction: numerical analysis and modelling. E and FN Spon, New York, 1994. ISBN 0-419-19070-8, 1-85861-014-1. xiv + 727 pp. LCCN TA711.5 .N86 1994. Title on spine: Soilstructure interaction: numerical analysis and modelling.

Bull:1996:NAM

[Bul96] John W. Bull. Numerical analysis and modelling of composite materials. Blackie Academic and Professional, New York, 1996. ISBN 0-7514-0239-7. xvii + 441 pp. LCCN TA418.9.C6 N85 1996.

Buneman:1990:PTS

[Bun90] Oscar Buneman. Particles in their self-consistent field: from Hartree’s differential analyzer to Cray machines. In Nash [Nas90b], pages 57–62. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Burmin:1993:MCO

[Bur93] V. Iu. (Valerii Iur’evich) Burmin. Metody chislennogo obrashcheniia godografov seismicheskikh voln. Nauka, Moscow, Russia, 1993. ISBN 5-02-002289-6. 108 pp. Nauchnoe izdanie. At head of title: Rossiiskaia Akademiia nauk. Ordena Lenina Institut fiziki Zemli im. O.IU. Shmidta. Title on verso of t.p.: Methods of numerical inversion travel-time curves of seismic waves. Summary and table of contents also in English.

Badriev:1990:CM

[BV90a] I. B. Badriev and S. N. Voloshanovskaya. Chislennye metody. (Russian) [Numerical methods]. Kazanskiĭ Gosudarstvennyĭ Universitet, Kazan′, Tatarstan, 1990. 172 pp. Priblizhenie funktsii i chislennoe integrirovanie. [Approximation of functions, and numerical integration].

Beelen:1990:CAJ

[BV90b] Theo Beelen and Paul Van Dooren. Computational aspects of the Jordan canonical form. In Cox and Hammarling [CH90], pages 57–72. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Bultheel:1997:LAR

[BV97] Adhemar Bultheel and Marc Van Barel. Linear algebra, rational approximation, and orthogonal polynomials, volume 6 of Studies in Computational Mathematics. Elsevier, Amsterdam, The Netherlands, 1997. ISBN 0-444-82872-9. xvii + 446 pp. LCCN QA242 .B88 1997. URL

http://www.loc.gov/catdir/
enhancements/fy0601/97040610-
d.html.

Bylaska:1994:SPN

[Byl94] Eric J. Bylaska. Scalable parallel numerical methods and software tools for material design. Technical report CS94-388, Dept. of Computer Science and Engineering, University of California, San Diego, San Diego, CA, USA, September 1994. 6 pp. To appear, Sixth SIAM Conference on Parallel Processing for Scientific Computing, February, 1995, San Francisco, CA.

Zhou:1993:NAE

[bZ93a] Pei bai Zhou. Numerical analysis of electromagnetic fields. Electric energy systems and engineering series. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-54722-3 (Berlin), 0-387-54722-3 (New York). xxi + 406 pp. LCCN QC665.E4 Z46 1993.

Breuer:1993:NML

[BZ93b] Shlomo Breuer and Gideon Zwas. Numerical mathematics — a laboratory approach. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-44040-8. xiv + 267 pp.

Coombes:1996:DEM

[C^{+}96] Kevin Robert Coombes et al. Differential equations with Maple. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1996. ISBN 0-471-10875-8. v + 232 pp. LCCN QA371.5.D37D53 1996.

Calvin:1996:AON

[Cal96] James M. Calvin. An asymptotically optimal non-adaptive algorithm for minimization of Brownian motion. In Renegar et al.

[RSS96], pages 157–164. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Carr:1995:NAG

[Car95] James R. Carr. Numerical analysis for the geological sciences. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-02-319511-8. xxv + 592 pp. LCCN QE33.2.M3 C37 1995.

Cardinal:1996:TIM

[Car96] Jeal-Paul Cardinal. On two iterative methods for approximating the roots of a polynomial. In Renegar et al. [RSS96], pages 165–188. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Chatelin:1990:PRE

[CB90] Françoise Chatelin and Marie-Christine Brunet. A probabilistic round-off error propagation model. application to the eigenvalue problem. In Cox and Hammarling [CH90], pages 139– 160. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Chaitin-Chatelin:1997:NSC

[CC97] F. Chaitin-Chatelin. Is nonnormality a serious computational difficulty in practice? In Boisvert [Boi97], pages 300–314.

ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Chapra:1998:NME

[CC98] Steven C. Chapra and Raymond P. Canale. Numerical methods for engineers: with programming and software applications. WCB/McGraw-Hill, Boston, MA, USA, third edition, 1998. ISBN 0-07-010938-9. xix + 924 pp. LCCN TA345 .C47 1998.

Chaitin-Chatelin:1993:EM

[CCA93] Francoise Chaitin-Chatelin and Mario Ahués. Eigenvalues of matrices. Pure and applied mathematics. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1993. ISBN 0-471-93538-7. xviii + 382 pp. LCCN QA188 .C44 1993. US$60.80. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0707/93003430-
                                        d.html;   http://www.loc.
                                        gov/catdir/toc/onix04/93003430.
                                        html.

Chaitin-Chatelin:1996:LFP

[CCF96] Françoise Chaitin-Chatelin and Valérie Frayssé. Lectures on Finite Precision Computations. Software, Environments, and Tools. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1996. ISBN 0-89871-358-7. xvi + 235 pp. LCCN QA297 .C417 1996. With a foreword by Iain S. Duff.

Chen:1999:BTN

[CCL99] Changxin Chen, Shui-Nee Chow, and Kaitai Li, editors. Bifurcation theory and its numerical analysis. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 981-4021-58-X. LCCN QA380 .I58 1998. Eighteen selected papers from the 2nd International Conference, Xi’an, China, June 29–July 3, 1998.

Celik:1990:NMM

[Cel90] I. Celik, editor. Numerical methods for multiphase flows: presented at the 1990 Spring Meeting of the Fluids Engineering Division, held in conjunction with the 1990 Forum of the Canadian Society of Mechanical Engineers, University of Toronto, Toronto, Ontario, Canada, June 4–7, 1990, volume 91 of FED. American Society of Mechanical Engineers, United Engineering Center, 345 E. 47th St., New York, NY 10017, USA, 1990. ISBN 0-7918-0488-7. LCCN QA922 .N86 1990. Papers collected in this volume represent the Second Symposium on Numerical Methods for Multiphase Flows organized by the Multiphase Flow Committee of the ASME Fluids Engineering Division.

Cerjan:1993:NGM

[Cer93] C. Cerjan, editor. Numerical grid methods and their application to Schrödinger’s equation, volume 412 of NATO ASI series. Series C, Mathematical and physical sciences. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1993. ISBN 0-7923-2423-4. LCCN QC174.26.W28 N38 1993. Published in cooperation with NATO Scientific Affairs Division. Proceedings of the NATO Advances Research Workshop on Grid Methods in Atomic and Molecular Quantum Calculations Corte, Corsica, France, September 27– October 3, 1992.

Celia:1992:NMD

[CG92] Michael Anthony Celia and William G. (William Guerin) Gray. Numerical methods for differential equations: fundamental concepts for scientific and engineering applications. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-626961-3. xii + 436 pp. LCCN QA372 .C39 1992. Fundamental concepts for scientific and engineering applications.

Conca:1997:NMM

[CG97] Carlos Conca and Gabriel N. Gatica, editors. Numerical methods in mechanics, volume 371 of Pitman research notes in mathematics series. Longman Scientific and Technical, Harlow, Essex, UK, 1997. ISBN 0-582-31320-1 (paperback). ISSN 0269-3674. LCCN QA801 .N86 1997. Selected papers from the 4th French-Latinamerican Congress on Applied Mathematics, held at the University of Concepcion, Concepcion, Chile, December 11–15, 1995.

Cerrolaza:1996:NME

[CGB96] M. Cerrolaza, C. Gajardo, and C. A. Brebbia. Numerical methods in engineering simulation. Computational Mechanics Publications, Boston, 1996. ISBN 1-85312-475-3. 412 pp. LCCN TA342.I566 1996. Third International Congress on Numerical Methods in Engineering and Applied Sciences (CIMENICS ’96), Merida, Venezuela, 25–29 March 1996.

Conn:1997:MNC

[CGT97] Andrew R. Conn, Nicholas I. M. Gould, and Philippe L. Toint. Methods for nonlinear constraints in optimization calculations. In Duff and Watson [DW97], pages 363–390. ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

//global.oup.com/academic/
product/the-state-of-the-
art-in-numerical-analysis-
9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Courant:1953:MMP

[CH53] Richard Courant and David Hilbert. Methods of mathematical physics, volume 1 of Wiley classics library. Interscience Publishers, New York, NY, USA, 1953. ISBN 0-471-50447-5, 0-585-29428-3 (ebook). xv + 560 pp. LCCN QA401 .C72413 1989.

Courant:1962:MMP

[CH62] Richard Courant and David Hilbert. Methods of mathematical physics, volume 2 of Wiley classics library. Interscience Publishers, New York, NY, USA, 1962. ISBN 0-471-50439-4, 0-585-32148-5 (ebook). ix + 830 pp. LCCN QA401 .C72413 1989. URL

http://www.loc.gov/catdir/
enhancements/fy0607/89112829-
b.html;   http://www.loc.
gov/catdir/enhancements/fy0607/
89112829-d.html;   http://
www.loc.gov/catdir/toc/onix05/
89112829.html.

Cox:1990:RNC

[CH90] M. G. Cox and S. Hammarling, editors. Reliable numerical computation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Chan:1992:NMS

[Cha92] William M. Chan. Numerical methods for the simulation of complex multi-body flows with applications for the Integrated Space Shuttle Vehicle. National Technical Information Service, Washington, DC, USA, 1992. ???? pp. Shipping list no.: 93-0040-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration], 1992. 1 microfiche.

Cha:1993:NAH

[Cha93a] Ellis Taeho Cha. Numerical analysis of head-disk assembly dynamics for shaped-rail sliders with sub-ambient pressure regions. Thesis (ph.d.) in mechanical engineering, University of California, Berkeley, Berkeley, CA, USA, May 1993. vii + 148 pp.

Champion:1993:NME

[Cha93b] Edward R. Champion, Jr. Numerical methods for engineering applications, volume 84 of Mechanical Engineering. Marcel Dekker, Inc., New York, NY, USA, 1993. ISBN 0-8247-9135-5. xv + 442 pp. LCCN TA335 .C48 1993.

Chatelin:1993:EM

[Cha93c] Françoise Chatelin. Eigenvalues of Matrices. Pure and applied mathematics. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1993. ISBN 0-471-93538-7. xviii + 382 pp. LCCN QA188 .C44 1993. URL

http://www.loc.gov/catdir/
enhancements/fy0707/93003430-
d.html;   http://www.loc.

                                        gov/catdir/toc/onix04/93003430.

html. With exercises by Mario Ahués and the author, Translated from the French and with additional material by Walter Ledermann.

Chan:1995:DNM

[Cha95] William M. Chan. Development of numerical methods for overset grids with applications for the integrated space shuttle vehicle. National Technical Information Service, Washington, DC, USA, 1995. ???? pp. Shipping list no.: 95-0749-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1995] 1 microfiche.

Charlton:1998:NSC

[Cha98] Eric F. Charlton. Numerical stability and control analysis towards falling-leaf prediction capabilities of splitflow for two generic high-performance aircraft models. National Technical Information Service, Washington, DC, USA, 1998. ???? pp. Shipping list no.: 99-0618-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1999] 1 microfiche.

Chu:1991:NSU

[Chu91] Keh-Cheng Chu. A numerical study of upwind methods for the Euler equations. Thesis (m.s.), University of California, Los Angeles, Los Angeles, CA, USA, 1991. viii + 61 pp.

Cividini:1998:ANM

[Civ98] Annamaria Cividini. Application of numerical methods to geotechnical problems: proceedings of the fourth European Conference on Numerical Methods in Geotechnical Engineering, NUMGE98, Udine, Italy, October 14–16, 1998, volume 397 of CISM courses and lectures / International Centre for Mechanical Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-211-83141-X. ISSN 0254-1971. 765 pp. LCCN TA703.5 .E92 1998 Bar. Organized by: European Regional Technical Committee ERTC7 … of the International Society for Soil Mechanics and Foundation Engineering.

Cheney:1994:NMC

[CK94] Ward Cheney and David Kincaid. Numerical Mathematics and Computing. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, third edition, 1994. ISBN 0-534-20112-1. 578 pp. LCCN QA297.C426 1994.

Ciarlet:1990:HNA

[CL90] Philippe G. Ciarlet and Jacques Louis Lions. Handbook of numerical analysis. Distributors for the United States and Canada, Elsevier Science Pub. Co., New York, 1990. ISBN 0-444-70366-7 (vol. 1), 0-444-70365-9 (vol. 2), 0-444-89928-6 (vol. 3), 0-444-81794-8 (vol. 4), 0-444-82569-X (vol. 6). various pp. LCCN QA297 .H287 1990 v.1-6 (1990-1998). Contents: vol. 1. Finite difference methods (pt. 1). Solutions of equations in Rn (pt. 1)

— vol. 2. Finite element methods (pt. 1) — vol. 3. Techniques of scientific computing (pt. 1). Numerical methods for solids (pt. 1). Solution of equations in Rn (pt. 2) – vol. 4. Finite element methods (pt. 2). Numerical methods for solids (pt. 2) — vol. 6. Numerical methods for solids (pt. 3). Numerical methods for fluids (pt. 1).

Cucker:1996:NTN

[CL96] Felipe Cucker and Thomas Lickteig. Nash trees and Nash complexity. In Renegar et al. [RSS96], pages 211–224. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Cardinal:1996:AAR

[CM96] J.-P. Cardinal and B. Mourrain. Algebraic approach of residues and applications. In Renegar et al. [RSS96], pages 189–210. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Constantinides:1999:NMC

[CM99] A. Constantinides and Navid Mostoufi. Numerical methods for chemical engineers with MATLAB applications. Prentice-Hall international series in the physical and chemical engineering sciences. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-013851-7. xxv + 560 pp. LCCN QA297 .C6494 1999.

Cohen:1990:HHA

[Coh90] I. Bernard Cohen. Howard H. Aiken and the computer. In Nash [Nas90b], pages 41– 52. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Cooley:1990:HFG

[Coo90] James W. Cooley. How the FFT gained acceptance. In Nash [Nas90b], pages 133–140. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Coombes:1997:DEM

[Coo97] Kevin Robert Coombes. Differential equations with Maple. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, second (revised for Maple V, release 4) edition, 1997. ISBN 0-471-17645-1. v + 240 pp. LCCN QA371.5.D37 D53 1997 Bar.

Corless:1995:SRS

[Cor95] Robert M. Corless. Symbolic Recipes: Scientific Computing with Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94210-6. 300 pp. US$39.00.

Corliss:1996:WVO

[Cor96] George F. Corliss. Where is validated ODE solving going? In Alefeld and Herzberger [AH96], pages 48–57. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Coutermarsh:1994:ENA

[Cou94] Barry A. Coutermarsh. Experimental and numerical analysis of floating ice beam impact forces against a sloped structure. National Technical Information Service, Washington, DC, USA, 1994. ???? pp. Shipping list no.: 94-0625-M. Microfiche. [Washington, DC?]: Supt. of Docs., U.S. G.P.O., [1994] 1 microfiche.

Coughran:1997:NSC

[Cou97] W. M. Coughran. Networkbased scientific computation via Inferno. In Boisvert [Boi97], pages 267–269. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Cox:1990:LSS

[Cox90] M. (Maurice) G. Cox. The least squares solution of linear equations with block-angular observation matrix. In Cox and Hammarling [CH90], pages 227– 240. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Crandall:1994:PSC

[Cra94] Richard E. Crandall. Projects in Scientific Computation. TE-LOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1994. ISBN 0-387-97808-9. xxiv + 470 pp. LCCN Q183.9.P733 1994. US$49.95. Includes computer disk.

Crandall:1996:TAS

[Cra96] Richard E. Crandall. Topics in Advanced Scientific Computation. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1996. ISBN 0-387-94473-7. ix + 340 pp. LCCN QA9.58.C73 1996. US$39.95. Includes programs in C and Mathematica.

Crowe:1994:NMM

[Cro94] C. T. Crowe, editor. Numerical methods in multiphase flows, 1994: presented at the 1994 ASME Fluids Engineering Division Summer Meeting, Lake Tahoe, Nevada, June 19– 23, 1994, volume 185 of FED. American Society of Mechanical Engineers, United Engineering Center, 345 E. 47th St., New York, NY 10017, USA, 1994. ISBN 0-7918-1368-1. LCCN TA357.5.M84 N8 1994. Papers collected in this volume represent the Third International Symposium on Numerical Methods in Multiphase Flows organized by the Multiphase Flow Committee of the ASME Fluids Engineering Division.

Chandra:1995:ANM

[CS95] Prabir K. Chandra and R. Paul Singh. Applied numerical methods for food and agricultural engineers. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1995. ISBN 0-8493-2454-8. 500 pp. LCCN QA300.C4415 1995.

Cutlip:1999:PSC

[CS99] Michael B. Cutlip and Mordechai Shacham. Problem solving in chemical engineering with numerical methods. Prentice-Hall international series in the physical and chemical engineering sciences. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-862566-2. xxxii + 458 pp. LCCN TP168 .C88 1999.

Cavallo:1996:UMS

[CSV96] Alberto Cavallo, Roberto Setola, and Francesco Vasca. Using MATLAB, Simulink, and Control Toolbox: a practical approach. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-261058-2 (paperback). x + 405 pp. LCCN QA297.C3713 1996.

Cuyt:1994:NNM

[Cuy94] Annie Cuyt, editor. Nonlinear numerical methods and rational approximation II, volume 296 of Mathematics and its applications. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1994. ISBN 0-7923-2967-8. LCCN QA297 .N642 1994. Proceedings of an international conference held at the University of Antwerp, Belgium, Sept. 5–11, 1993.

Chiang:1990:ESM

[CZ90] M. K. Chiang and S. A. Zenios. An experts system for matrix balancing. In Houstis et al. [HRV90], pages 343–352. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

ElDabaghi:1998:ANM

[Dab98] F. El Dabaghi. Approximations and numerical methods for the solution of Maxwell’s Equations, volume 65 of The Institute of Mathematics and Its Applications conference series: new series. Clarendon Press, New York, NY, USA, 1998. ISBN 0-19-851483-2. xvi + 390 pp. LCCN QC670.75 .A66 1998. Based on the proceedings of the third international conference on approximations and numerical methods for the solution of Maxwell’s equations, organised by the Institute of Mathematics and its Applications and held at St. Johns College at the University of Oxford in April 1995.

Dantzig:1990:OSM

[Dan90] George B. Dantzig. Origins of the simplex method. In Nash [Nas90b], pages 141–151. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Datta:1995:NLA

[Dat95] Biswa Nath Datta. Numerical linear algebra and applications. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1995. ISBN 0-534-17466-3 (paperback). xxii + 680 pp. LCCN QA184 .D37 1995.

Deuflhard:1994:NMI

[DB94] Peter Deuflhard and Folkmar Bornemann. Numerische Mathematik. II. (German) [Numerical mathematics. II]. de Gruyter Lehrbuch. [de Gruyter Textbook]. Walter de Gruyter, New York, 1994. ISBN 3-11-013936-7; 3-11-013937-5. xvi + 383 pp. Integration gewöhnlicher Differentialgleichungen. [Integration of ordinary differential equations].

Doshi:1990:IAI

[DDLW90] R. S. Doshi, R. S. Desai, R. Lam, and J. E. White. Integration of artificial intelligence planning and robotic systems with Airobic. In Houstis et al. [HRV90], pages 227–240. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Deshpande:1995:FIC

[DDN95] S. M. Deshpande, S. S. Desai, and R. Narasimha. Fourteenth International Conference on Numerical Methods in Fluid Dynamics: proceedings of the conference held in Bangalore, India, 11–15 July 1994, volume 453 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-59280-6. xiii + 588 + [1] pp. LCCN QA911 .I58 1994.

Dongarra:1991:SLS

[DDSv91] Jack J. Dongarra, Iain S. Duff, Danny C. Sorensen, and Henk A. van der Vorst. Solving Linear Systems on Vector and Shared Memory Computers. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1991. ISBN 0-89871-270-X. x + 256 pp. LCCN QA184 .S65 1991.

Dongarra:1998:NLA

[DDSvdV98] Jack J. Dongarra, Iain S.

Duff, Danny C. Sorensen, and Henk A. van der Vorst. Numerical Linear Algebra for High-Performance Computers. Software, Environments, and Tools. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1998. ISBN 0-89871-428-1. xviii + 342 pp. LCCN QA76.88 .N86 1998. URL

http://www.loc.gov/catdir/
enhancements/fy0726/98044444-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0726/
98044444-t.html.

Dedieu:1996:ASN

[Ded96] Jean-Pierre Dedieu. Approximate solutions for numerical problems, condition number analysis and condition number theorem. In Renegar et al. [RSS96], pages 263–284. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Deineko:1994:MPR

[Deĭ94] V. V. Deĭneko. Metody priblizhennogo resheniya obyknovennykh differentsialnykh uravnenii. (Russian) [Methods for the approximate solution of ordinary differential equations]. “Izdatel′”, Novosibirsk, Russia, second edition, 1994. ISBN 5-88399-074-6. 239 pp.

Demmel:1990:NDM

[Dem90] James Demmel. Nearest defective matrices and the geometry of ill-conditioning. In Cox and Hammarling [CH90], pages 35–

55. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Demmel:1997:ANL

[Dem97] James W. Demmel. Applied Numerical Linear Algebra. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1997. ISBN 0-89871-389-7 (paperback). xi + 419 pp. LCCN QA184 .D455 1997. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0664/97017290-
                                        d.html;   http://www.loc.
                                        gov/catdir/enhancements/fy0664/
                                        97017290-t.html;   http://
                                        www.siam.org/books/demmel/
                                        demmel_class.

Desideri:1996:CMA

[Des96] J.-A. Desideri. Computational methods in applied sciences ’96: invited lectures and special technological sessions of the Third ECCOMAS Computational Fluid Dynamics Conference and the Second ECCO-MAS Conference on Numerical Methods in Engineering, 9–13 September 1996, Paris, France. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1996. ISBN 0-471-96226-0. xiii + 572 pp. LCCN TA357 .C5884 1996.

Desideri:1998:MDS

[Dés98] Jean-Antoine Désidéri. Modèles discrets et schémas itératifs. (French) [Discrete models and iterative schemes]. Hermès, Paris, France, 1998. ISBN 2-86601-710-2. 349 pp. Application aux algorithmes multigrilles et multidomaines. [Application to multigrid and multidomain algorithms].

Doi:1997:APP

[DFS97] S. Doi, H. Fujio, and K. Sugihara. Automatic parallel program generation for finite element analysis. In Boisvert [Boi97], pages 255–266. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Desai:1990:NMC

[DG90a] C. S. Desai and G. Gioda. Numerical methods and constitutive modelling in geomechanics, volume 311 of Courses and lectures. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 3-211-82215-1 (Wien-New York), 0-387-82215-1 (New York-Wien). 407 pp. LCCN TA350.3.I5 no.311. At head of title: International Centre for Mechanical Sciences.

Dyksen:1990:EEE

[DG90b] W. R. Dyksen and C. R. Gritter. Elliptic expert: An expert system for elliptic partial differential equations. In Houstis et al. [HRV90], pages 23–32. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Goede:1993:NMT

[dG93]  E. D. de Goede.  Numer-

ical methods for the threedimensional shallow water equations on supercomputers, volume 88 of CWI tract. CWI, Amsterdam, The Netherlands, 1993. ISBN 90-6196-417-2. 124 pp. LCCN QA297 .G63 1993.

Duff:1990:MMP

[DGLR90] Iain S. Duff, Nicholas I. M. Gould, Marc Lescrenier, and John K. Reid. The multifrontal method in a parallel environment. In Cox and Hammarling [CH90], pages 93–111. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Dedieu:1996:CDP

[DGY96] J. P. Dedieu, X. Gourdon, and J. C. Yakoubsohn. Computing the distance from a point to an algebraic hypersurface. In Renegar et al. [RSS96], pages 285– 294. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Dongarra:1990:ENS

[DH90] Jack Dongarra and Sven Hammarling. Evolution of numerical software for dense linear algebra. In Cox and Hammarling [CH90], pages 297–327. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Deuflhard:1991:NM

[DH91] Peter Deuflhard and Andreas Hohmann. Numerische Mathematik. (German) [Numerical mathematics]. de Gruyter Lehrbuch. [de Gruyter Textbook]. Walter de Gruyter, New York, 1991. ISBN 3-11-012917-5; 3-11-012918-3. xvi + 339 pp. Eine algorithmisch orientierte Einführung. [An algorithmically oriented introduction].

Deuflhard:1993:NM

[DH93] Peter Deuflhard and Andreas Hohmann. Numerische Mathematik. I. (German) [Numerical mathematics. I]. de Gruyter Lehrbuch. [de Gruyter Textbook]. Walter de Gruyter, New York, second edition, 1993. ISBN 3-11-013975-8; 3-11-013974-X. xvi + 371 pp. Eine algorithmisch orientierte Einführung. [An algorithmically oriented introduction].

Deuflhard:1995:NAF

[DH95a] P. (Peter) Deuflhard and Andreas Hohmann. Numerical analysis: a first course in scientific computation. Walter de Gruyter, New York, 1995. ISBN 3-11-014031-4 (hardcover), 3-11-013882-4 (paperback). xiii + 355 pp. LCCN QA297 .D4713 1995.

Deuflhard:1995:NA

[DH95b] Peter Deuflhard and Andreas Hohmann. Numerical analysis. de Gruyter Textbook. Walter de Gruyter, New York, 1995. ISBN 3-11-014031-4; 3-11-013882-4. xiv + 355 pp. A first course in scientific computation, Translated from the second (1993) German edition by F. A. Potra and F. Schulz.

Dongarra:1997:CSD

[DHP97] J. J. Dongarra, S. Hammarling, and A. Petitet. Case studies on the development of ScaLA-PACK and the NAG numerical PVM library. In Boisvert [Boi97], pages 236–248. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Diethelm:1996:EBC

[Die96] Kai Diethelm. Error bounds for compound quadratures for Hadamard-type finite-part integrals. In Alefeld and Herzberger [AH96], pages 58–63. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Donea:1991:ENM

[DJ91] J. Donea and P. M. Jones. Experimental and numerical methods in earthquake engineering, volume 2 of Eurocourses. Reliability and risk analysis. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1991. ISBN 0-7923-1434-4 (hardcover). x + 381 pp. LCCN TA654.6 .E98 1991. Based on the lecture notes prepared for the advanced course … held at the Joint Research Centre (JRC) of the Commission of the European Communities, Ispra Establishment, in October 1991.

Dubois:1999:DMM

[DJT99] Thierry Dubois, François Jauberteau, and Roger Temam. Dynamic multilevel methods and the numerical simulation of turbulence. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-62165-8 (hardcover). xix + 289 pp. LCCN QA913 .D88 1999.

Dahmen:1996:OEM

[DKS96] Wolfgang Dahmen, Angela Kunoth, and Reinhold Schneider. Operator equations, multiscale concepts and complexity. In Renegar et al. [RSS96], pages 225–262. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Dobronets:1996:TSM

[Dob96] Boris S. Dobronets. Two-sided methods based on defects. In Alefeld and Herzberger [AH96], pages 64–73. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Donnelly:1995:MMC

[Don95]  Karen Donnelly.  MATLAB

manual: computer laboratory exercises. Saunders College Pub., Fort Worth, TX, USA, 1995. ISBN 0-03-094896-7. v + 196 + 16 pp. LCCN ????

Dontchev:1996:LAN

[Don96] A. L. Dontchev. Local analysis of a Newton-type method based on partial linearization. In Renegar et al. [RSS96], pages 295– 306. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Dormand:1996:NMD

[Dor96] John R. Dormand. Numerical methods for differential equations: a computational approach. Library of Engineering Mathematics. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1996. ISBN 0-8493-9433-3. xiv + 368 pp. LCCN QA372 .D67 1996. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0744/95054095-

d.html. A computational approach, With 1 IBM-PC floppy disc (3.5 inch; HD).

Dorfman:1997:BET

[Dor97] Jeffrey H. Dorfman. Bayesian economics through numerical methods: a guide to econometrics and decision-making with prior information. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-98233-7 (New York). vii + 110 pp. LCCN HB139 .D674 1997.

DanielKowalski:1990:AAA

[DPR90] A. Daniel Kowalski, R. L. Peskin, and M. F. Russo. Anatomy of AGNES: An automatic generator of numerical equation solutions. In Houstis et al. [HRV90], pages 33–42. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Dobronets:1990:TSN

[DS90] B. S. Dobronets and V. V. Shaydurov. Two-sided Numerical Methods. Nauka (Siberian Department), Novosibirsk, Russia, 1990. ISBN ???? 208 pp. LCCN ????

deFigueiredo:1997:SVN

[dS97] Luiz H. de Figueiredo and Jorge Stolfi. Self-Validated Numerical Methods and Applications. Brazilian Mathematics Colloquium monograph. IMPA, Rio de Janeiro, Brazil, July 1997. ISBN ???? LCCN

???? URL http://www.dcc.
unicamp.br/~stolfi/EXPORT/
papers/affine-arith/aa-97-
07-val-num-book.ps.gz.

Dimov:1994:ANM

[DSV94] Ivan Dimov, Blagovest Sendov, and Panaiot Vasilevski, editors. Advances in numerical methods and applications: proceedings of the third international conference: Sofia, Bulgaria, 21– 26 August 1994. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1994. ISBN 981-02-1926-1. LCCN QA297 .I516 1994. Papers presented at the Third International Conference on Numerical Methods and Applications organized by the Bulgarian Academy of Sciences.

Dahlen:1997:NMS

[DT97] Morten Dahlen and Aslak Tveito. Numerical methods and software tools in industrial mathematics. Birkhäuser Verlag, Basel, Switzerland, 1997. ISBN 0-8176-3973-X (hardcover), 3-7643-3973-X. 400 pp. LCCN TA331 .N86 1997.

DuCroz:1997:DBN

[Du 97] J. Du Croz. Designing and building a new numerical library in Fortran 90. In Boisvert [Boi97], pages 197–209. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Dubois:1997:OTS

[Dub97] Paul F. Dubois. Object technology for scientific computing: object-oriented numerical software in Eiffel and C. Prentice Hall object-oriented series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-267808-X. xviii + 280 pp. LCCN QA76.9.O35 D83 1997.

Duff:1997:SNL

[Duf97] Iain S. Duff. Sparse numerical linear algebra: Direct methods and preconditioning. In Duff and Watson [DW97], pages 27– 62. ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

//global.oup.com/academic/
product/the-state-of-the-
art-in-numerical-analysis-
9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Durran:1999:NMW

[Dur99] Dale R. Durran. Numerical methods for wave equations in geophysical fluid dynamics, volume 32 of Texts in applied mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98376-7 (New York). xvii + 465 pp. LCCN QC809.F5 D87 1999.

Duff:1997:SAN

[DW97] Iain S. Duff and G. Alistair Watson, editors. The state of the art in numerical analysis, volume 63 of The Institute of Mathematics and Its Applications conference series: new series. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1997. ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

//global.oup.com/academic/
product/the-state-of-the-

                                        art-in-numerical-analysis-
                                        9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Eaton:1997:GOH

[Eat97] John W. Eaton. GNU Octave: a high-level interactive language for numerical computations. Network Theory Ltd., Bristol, UK, 1997. ISBN 0-9541617-2-6. viii + 311 pp. LCCN ????

Engquist:1991:TIC

[EG91] Björn Engquist and Bertil Gustafsson. Third International Conference on Hyperbolic Problems: theory, numerical methods and applications: proceedings, Uppsala, Sweden, June 11–15, 1990. Chartwell-Bratt, Bromley, Kent, UK, 1991. ISBN 91-44-33421-4 (Studentlitteratur), 0-86238-285-8 (Chartwell Bratt). 517 (vol. 1), 519–1024 (vol. 2) pp. LCCN QA377.I563 1990. The two earlier conferences … were held in St. Etienne, France, 1986 and Aachen, Germany, 1988.

Espelid:1992:NIR

[EG92] Terje O. Espelid and Alan Genz, editors. Numerical integration: recent developments, software, and applications, volume 357 of NATO ASI series. Series C, Mathematical and physical sciences. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1992. ISBN 0-7923-1583-9. LCCN QA299.3.N38 1991. Published in cooperation with NATO Scientific Affairs Division. Proceedings of the NATO Advanced Research Workshop on Numerical Integration: Recent Developments, Software, and Applications, Bergen, Norway, June 17–21, 1991.

Emiris:1996:NUP

[EGL96] Ioannis Z. Emiris, André Galligo, and Henri Lombardi. Numerical univariate polynomial GCD. In Renegar et al. [RSS96], pages 323–344. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Enright:1997:ENS

[EH97] W. H. Enright and H. Hayashi.

The evaluation of numerical software for delay differential equations. In Boisvert [Boi97], pages 179–196. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Ehrich:1996:PEE

[Ehr96] Sven Ehrich. Practical error estimates for the Gauss–Kronrod quadrature formula. In Alefeld and Herzberger [AH96], pages 74–79. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Elliott:1997:ACD

[Ell97] C. M. Elliott. Approximation of curvature dependent interface motion. In Duff and Watson [DW97], pages 407–440.

ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

//global.oup.com/academic/
product/the-state-of-the-
art-in-numerical-analysis-
9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Elton:1990:NTL

[Elt90] Altamont Bracy Hamilton Elton. A numerical theory of lattice gas and lattice Boltzmann methods in the computation of solutions to nonlinear advectivediffusive systems. Thesis (ph.d.) in computer science, University of California, Davis, Davis, CA, USA, 1990. 258 pp.

Emery:1990:SNM

[Eme90] A. F. Emery, editor. Simulation and numerical methods in heat transfer: presented at the Winter Annual Meeting of the American Society of Mechanical Engineers, Dallas, Texas, November 25–30, 1990, volume 157 of HTD. American Society of Mechanical Engineers, United Engineering Center, 345 E. 47th St., New York, NY 10017, USA, 1990. ISBN 0-7918-0597-2. LCCN TJ260 .A415 1990c. Engeln-Mullges:1996:NAC

[EMU96a] Gisela Engeln-Müllges and Frank Uhlig. Numerical algorithms with C. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-60530-4 (hardcover). xxii + 596 pp. LCCN QA297 .E56213 1996.

URL http://www.loc.gov/
catdir/enhancements/fy0815/
96023158-d.html. Appendix

and CD-ROM with ANSI C programs by Albert Becker, Jürgen Dietel and Uli Eggermann, Translated from the seventh (1993) German edition by Maria Schon and Uhlig.

Engeln-Mullges:1996:NAF

[EMU96b] Gisela Engeln-Mullges and Frank Uhlig. Numerical Algorithms With Fortran. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., November 1996. ISBN 3-540-60529-0. xxii + 602 pp. LCCN QA297 .E56213 1996b Permanent Reserve. US$49.95. URL http:/

/www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
3540605290.

Eaves:1996:ACC

[ER96] B. Curtis Eaves and Uriel G.

Rothblum. Approximations and complexity for computing algebraic curves. In Renegar et al. [RSS96], pages 307–322. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Erbes:1994:SFA

[Erb94] Gerhard Erbes. A study of finescale atmospheric dynamics and numerical advection methods.

Dept. of Meteorology, Stockholm University, 1994. ISBN 91-7153-205-6. various pp. LCCN QC880.4 .E8 1994. Originally presented as the author’s thesis (doctoral)–Stockholm University, 1994.

Erhel:1996:PPG

[Erh96] J. Erhel. A parallel preconditioned GMRES algorithm for sparse matrices. In Renegar et al. [RSS96], pages 345–356. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Etter:1992:FNM

[Ett92]  D. M. Etter.  Fortran 77

with numerical methods for engineers and scientists. Benjamin/ Cummings Pub. Co., Menlo Park, CA, USA, 1992. ISBN 0-8053-1770-8. xxv + 776 pp. LCCN QA76.73.F25 E847 1992. System requirements for computer disk (ASCII files): PC; FORTRAN 77 compiler.

Euvrard:1994:RNE

[Euv94] Daniel Euvrard. Résolution numérique des équations aux dérivées partielles de la physique, de la mécanique et des sciences de l’ingénieur. (French) [Numerical solution of equations of partial derivatives in physics, mechanics, and engineering]. Enseignement de la Physique: Mathématiques pour la Physique. [The Teaching of Physics: Mathematics for Physics]. Masson Editeur, Masson, France, third edition, 1994. ISBN 2-225-84509-3. xiv + 330 pp. Différences finies, éléments finis, problèmes en domaine non borné. [Finite differences, finite elements, problems in unbounded domains].

Evans:1993:PNI

[Eva93] Gwynne Evans. Practical numerical integration. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1993. ISBN 0-471-93898-X. xii + 328 pp. LCCN QA299.3 .E93 1993. US$63.75.

Evans:1995:PNA

[Eva95] Gwynne Evans. Practical numerical analysis. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1995. ISBN 0-471-95535-3. xiii + 455 pp. LCCN QA297 .E93 1995.

Elden:1990:NA

[EWK90a] Lars Eldén and Linde Wittmeyer-Koch. Numerical analysis. Academic Press, New York, NY, USA, 1990. ISBN 0-12-236430-9. x + 347 pp. An introduction, Translated from the Swedish.

Elden:1990:NAI

[EWK90b] Lars Eldén and Linde Wittmeyer-Koch. Numerical analysis: an introduction. Academic Press, New York, NY, USA, 1990. ISBN 0-12-236430-9. x + 347 pp. LCCN QA297 .E4213 1990. Rev. translation of: Numerisk analys.

Eidson:1991:FAD

[EZ91] T. M. Eidson and Thomas A. Zang. Filtering analysis of a direct numerical simulation of the turbulent Rayleigh-Benard problem. NASA, Washington, DC, USA, 1991. ???? pp. Microfiche. [Washington, DC?: National Aeronautics and Space Administration], 1991. 1 microfiche.

Fausett:1999:ANA

[Fau99] Laurene V. Fausett. Applied numerical analysis using MAT-LAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-319849-9. xii + 596 pp. LCCN QA297 .F38 1999.

Faires:1993:NM

[FB93] J. Douglas Faires and Richard L. Burden. Numerical methods. Prindle, Weber and Schmidt series in mathematics. PWS-Kent Pub. Co., Boston, MA, USA, 1993. ISBN 0-534-93136-7. xiii + 503 pp. LCCN QA297 .F35 1993. System requirements for computer disk: IBM PC or compatible; PC DOS or MS DOS; Pascal or FORTRAN compiler. s.

Faires:1998:NM

[FB98] J. Douglas Faires and Richard L. Burden. Numerical methods. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, second edition, 1998. ISBN 0-534-35187-5. xi + 594 pp. LCCN QA297 .F35 1998. With 1 IBM-PC floppy disk (3.5 inch; HD).

Ferziger:1998:NME

[Fer98] Joel H. Ferziger. Numerical methods for engineering application. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, second edition, 1998. ISBN 0-471-11621-1 (cloth). xvii + 378 pp. LCCN TA335 .F47 1998. A Wiley-Interscience publication.

Fox:1990:UNN

[FFK90] G. Fox, W. Furmanski, and J. Koller. The use of neural networks in parallel software systems. In Houstis et al. [HRV90], pages 51–62. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Ford:1990:NKS

[FHI90] B. Ford, S. J. Hague, and R. M. J. Iles. Numerical knowledge-based systems. In Houstis et al. [HRV90], pages 281–286. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Finlayson:1992:NMP

[Fin92] Bruce A. Finlayson. Numerical methods for problems with moving fronts. Ravenna Park Pub., Seattle, Wash., USA, 1992. ISBN 0-9631765-0-1. vii + 605 pp. LCCN TA347.F5 F565 1992. System requirements for computer disk: PC.

Fiume:1995:ISS

[Fiu95] Eugene L. Fiume. An introduction to scientific, symbolic, and graphical computation. A. K. Peters, Ltd., Wellesley, MA, USA, 1995. ISBN 1-56881-051-2. xvi + 306 pp. LCCN QA76.F54 1995.

Fosdick:1996:IHP

[FJSD96a] Lloyd D. Fosdick, Elizabeth R. Jessup, Carolyn J. C. Schauble, and Gitta Domik. An Introduction to High-Performance Scientific Computing. Scientific and Engineering Computation. MIT Press, Cambridge, MA, USA, 1996. ISBN 0-262-06181-

3. xxiii + 760 pp. LCCN QA76.A594 1996. US$65.00.

URL http://www.mitpress.
com/book-home.tcl?isbn=0262061813.

Fosdick:1996:IHS

[FJSD96b] Lloyd D. Fosdick, Elizabeth R. Jessup, Carolyn J. C. Schauble, and Gitta Domik. An Introduction to High-Performance Scientific Computing. MIT Press, Cambridge, MA, USA, 1996. ISBN 0-262-06181-3. xxiii + 760 pp. LCCN QA76 .A594 1996. US$70.00.

Friedman:1994:FCN

[FK94] Menahem Friedman and Abraham Kandel. Fundamentals of computer numerical analysis. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1994. ISBN 0-8493-8637-3. xi + 587 pp. LCCN QA76 .F757 1994. System requirements for computer disk: IBM-compatible PC; DOS; Fortran 77.

Flowers:1995:INM

[Flo95] Sir Brian Hilton Flowers. An introduction to numerical methods in C++. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1995. ISBN 0-19-853863-4 (paperback). xx + 491 pp. LCCN QA297.F54 1995.

Ferrari:1992:SCN

[FM92] Mauro Ferrari and Nicola Marzari. Symbolic computing and numerical techniques in the residual stress analysis of media with inhomogeneous microstructures. Technical Report UCB/SEMM-92/18, Dept. of Civil Engineering, University of California, Berkeley, Berkeley, CA, USA, June 1992. [19] pp.

Fedkiw:1996:NMM

[FMO96a] Ronald P. Fedkiw, Barry Merriman, and Stanley Osher. Numerical methods for a mixture of thermally perfect and/or calorically perfect gaseous species with chemical reactions. Dept. of Mathematics, University of California, Los Angeles, Los Angeles, CA, USA, January 1996. 51 pp.

Fedkiw:1996:NMO

[FMO96b] Ronald P. Fedkiw, Barry Merriman, and Stanley Osher. Numerical methods for a onedimensional interface separating compressible and incompressible flows. Dept. of Mathematics, University of California, Los Angeles, Los Angeles, September 1996. 44 pp.

Fornberg:1996:PGP

[For96] Bengt Fornberg. A Practical Guide to Pseudospectral Methods. Cambridge Monographs on Applied and Computational Mathematics. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-49582-2. x + 231 pp. LCCN QA320 F65 1996.

                                        URL http://www.loc.gov/
                                        catdir/description/cam027/
                                        95031986.html;   http://www.
                                        loc.gov/catdir/toc/cam024/
                                        95031986.html.

Fox:1990:ENA

[Fox90a] L. Fox. Early numerical analysis in the United Kingdom. In Nash [Nas90b], pages 280–300. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Fox:1990:EJW

[Fox90b] L. Fox. Epilogue. Jim Wilkinson: some after-dinner sentiments. In Cox and Hammarling [CH90], pages 329–335. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Fox:1990:NST

[Fox90c] L. (Leslie) Fox. The numerical solution of two-point boundary problems in ordinary differential equations. Dover, New York, NY, USA, 1990.

ISBN 0-486-66495-3. xi + 371 pp. LCCN QA379 .F69 1990. US$9.95. URL

http://www.loc.gov/catdir/
description/dover032/90048640.

html. Reprint of the 1957 original.

Freeman:1992:PNA

[FP92] T. L. (T. Len) Freeman and Chris Phillips. Parallel numerical algorithms. Prentice Hall International Series in Computer Science. Prentice-Hall International, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-651597-5 (paperback). xi + 315 pp. LCCN QA76.9.A43 F74 1992. US$35.95.

Frank:1996:OAL

[Fra96] Karin Frank. An optimal algorithm for the local solution of integral equations. In Renegar et al. [RSS96], pages 357–380. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Friedman:1997:INM

[Fri97] Zvi Friedman. Improved numerical methods for solutions of a beam’s axial, torsion and flexure problems. Thesis (ph.d.), University of California, San Diego, San Diego, CA, USA, 1997.

Froberg:1990:BCC

[Frö90] Carl-Erik Fröberg. BIT — a child of the computer. In Nash [Nas90b], pages 228–233. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Faghri:1990:NEA

[FS90] Amir Faghri and Theodore D. Swanson. Numerical and experimental analysis of a thin liquid film on a rotating disk related to development of a spacecraft absorption cooling system. NASA, Washington, DC, USA, 1990. ???? pp. Microfiche. [Washington, DC?: National Aeronautics and Space Administration], 1990. 4 microfiches.

Frontini:1992:CNN

[FT92] Marco Frontini and Aldo Tagliani. Calcolo Numerico— Esercizi Numerical Calculations— Exercises . clup CittàStudi, Milano, Italy, 1992. ISBN 88-251-7056-4. LCCN ????

Gomez:1999:ESC

[G^{+}99] Claude Gomez et al., editors. Engineering and scientific computing with Scilab. Birkhäuser Verlag, Basel, Switzerland, 1999. ISBN 0-8176-4009-6, 3-7643-4009-6. xxv + 491 pp. LCCN TA345 .E5334 1999.

Garcia:1994:NMP

[Gar94] Alejandro L. Garcia. Numerical methods for physics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-151986-7. xiii + 368 pp. LCCN QC20.G37 1994.

Gautschi:1997:NA

[Gau97a] Walter Gautschi. Numerical analysis. Birkhäuser Boston Inc., Cambridge, MA, USA, 1997. ISBN 0-8176-3895-4. xiv + 506 pp. An introduction.

Gautschi:1997:NAI

[Gau97b]  Walter Gautschi.  Numer-

ical analysis: an introduction. Birkhäuser Verlag, Basel, Switzerland, 1997. ISBN 0-8176-3895-4, 3-7643-3895-4 (Basel). xiii + 506 pp. LCCN QA297 .G35 1997.

Girimaji:1997:AMB

[GB97] Sharath S. Girimaji and S. Balachandar. Analysis and modeling of buoyancy generated turbulence using numerical data. National Technical Information Service, Washington, DC, USA, 1997. ???? pp. Shipping list no.: 99-0141-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1998] 1 microfiche. (NASA-CR ; 201736).

Girimaji:1995:AMS

[GC95] Sharath S. Girimaji and Yeh Chou. Analysis and modeling of subgrid scalar mixing using numerical data. National Technical Information Service, Washington, DC, USA, 1995. ???? pp. Shipping list no.: 96-0033-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1995] 1 microfiche.

Geers:1990:SLA

[Gee90] N. Geers. Subroutine libraries at the University of Karlsruhe Computing Center. In Houstis et al. [HRV90], pages 215–226. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Gentleman:1997:ISQ

[Gen97] W. M. Gentleman. If software quality is a perception, how do we measure it? In Boisvert [Boi97], pages 32–43. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Gentle:1998:NLA

[Gen98] James E. Gentle. Numerical Linear Algebra for Applications in Statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-95489-9. xiii + 221 pp. LCCN QA184 .G45 1998.

Gilewicz:1990:ESC

[GGM90] J. Gilewicz, M. Griffiths, and A. Mortada. An expert system for the choice and parameterization of algorithms for Padé approximation. In Houstis et al. [HRV90], pages 161–166. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Gander:1993:SPS

[GH93] Walter Gander and Jiři Hřebíček. Solving Problems in Scientific Computing Using MAPLE and MATLAB. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-57329-1, 3-540-57329-1. xiii + 268 pp. LCCN Q183.9 .G36 1993.

Gomez:1994:AON

[GH94] S. Gomez and J. P. Hennart, editors. Advances in optimization and numerical analysis: proceedings of the Sixth Workshop on Optimization and Numerical Analysis, Oaxaca, Mexico, volume 275 of Mathematics and its applications. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1994. ISBN 0-7923-2673-3. LCCN QA402.5 .W68 1992.

Gander:1995:SPS

[GH95] Walter Gander and Jiřı Hřebíček.

Solving Problems in Scientific Computing Using Maple and MATLAB. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1995. ISBN 3-540-58746-2, 0-387-58746-2. xv + 315 pp. LCCN Q183.9 .G36 1995.

Gower:1996:Bb

[GH96] John Clifford Gower and David J.

Hand. Biplots, volume 54 of Monographs on statistics and applied probability. Chapman and Hall, Ltd., London, UK, 1996. ISBN 0-412-71630-5. ISSN 0960-6696. xvi + 277 pp. LCCN ????

Gander:1997:SPS

[GH97] Walter Gander and Jiři Hřebíček.

Solving Problems in Scientific Computing using Maple and MATLAB. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1997. ISBN 3-540-61793-0 (paperback). xvii + 408 pp. LCCN Q183.9 .G36 1997. US$49.95.

Gallivan:1990:PAM

[GHN^{+}90] K. A. Gallivan, Michael T. Heath, Esmond Ng, James M. Ortega, Barry W. Peyton, R. J. Plemmons, Charles H. Romine, A. H. Sameh, and Robert G. Voigt. Parallel Algorithms for Matrix Computations. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1990. ISBN 0-89871-260-2 (paperback), 1-61197-170-5 (ebook). ix + 197 pp. LCCN QA188 .P367 1990.

Gruber:1990:INS

[GhW90] Arnold Gruber and Tsung hao Wang. The impact of NOAA satellite soundings on the numerical analysis and forecast system of the People’s Republic of China. U.S. Dept. of Commerce, National Oceanic and Atmospheric Administration, National Environmental Satellite, Data, and Information Service, May 1990. ix + 30 + [41] + 25 + [26] pp. Contents: pt. I. FGGE data — pt. II. December 1987.

Griffiths:1998:NAP

[GHW98] D. F. Griffiths, D. J. Higham, and G. A. Watson. Numerical analysis 1997: proceedings of the 17th Dundee Biennial Conference, June 24–27, 1997, volume 380 of Pitman research notes in mathematics series. Longman Scientific and Technical, Harlow, Essex, UK, 1998. ISBN 0-582-31261-2 (paperback). ISSN 0269-3674. 254 pp. LCCN QA297.D85 1997.

Girvin:1992:NAR

[Gir92] Douglas Lynn Girvin. Numerical analysis of right-half plane zeros for a single-link manipulator: a thesis presented to the academic faculty. National Technical Information Service, Washington, DC, USA, 1992. ???? pp. Shipping list no.: 93-0324-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration], 1992. 2 microfiches.

Givoli:1992:NMP

[Giv92] Dan Givoli. Numerical methods for problems in infinite domains, volume 33 of Studies in applied mechanics. Elsevier, Amsterdam, The Netherlands, 1992. ISBN 0-444-88820-9. xv + 299 pp. LCCN TA347.B69 G58 1992.

Guenin:2014:GIO

[GKT14] B. (Bertrand) Guenin, J. (Jochen)

Könemann, and Levent Tuncel. A gentle introduction to optimization. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-05344-7 (hardcover), 1-107-65879-9 (paperback), 1-139-99068-3 (e-book). xi + 269 pp. LCCN QA402.5 .G84 2014.

URL http://www.cambridge.
org/us/academic/subjects/
mathematics/optimization-

                                        or-and-risk-analysis/gentle-
                                        introduction-optimization.

Gladwell:1990:ACS

[GL90] I. Gladwell and M. Lucks. An automated consultation system for the selection of mathematical software. In Houstis et al. [HRV90], pages 179–186. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Glassey:1993:NCU

[Gla93]  Robert Glassey.  Numerical

Computation Using C. Academic Press, New York, NY, USA, 1993. ISBN 0-12-286155-8. vii + 283 pp.

Gradel:1996:DCT

[GM96] Erich Grädel and Klaus Meer. Descriptive complexity theory over the real numbers. In Renegar et al. [RSS96], pages 381– 404. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Guichard:1997:PDE

[GM97] Frédéric Guichard and Jean-Michel Morel. Partial differential equations and image iterative filtering. In Duff and Watson [DW97], pages 525–562. ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

                                        //global.oup.com/academic/
                                        product/the-state-of-the-
                                        art-in-numerical-analysis-
                                        9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Gill:1990:SCM

[GMSW90] Philip E. Gill, Walter Murray, Michael A. Saunders, and Margaret H. Wright. A Schurcomplement method for sparse quadratic programming. In Cox and Hammarling [CH90], pages 113–138. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Gill:1991:NLA

[GMW91] Philip E. Gill, Walter Murray, and Margaret H. Wright. Numerical linear algebra and optimization. Vol. 1. Addison-Wes-ley, Reading, MA, USA, 1991. ISBN 0-201-12649-4 (vol. 1). xx + 426 pp. LCCN QA184 .N86 1989.

Gladwell:1997:EGA

[GN97] I. Gladwell and M. A. Napierala. Efficiency of global adaptive quadrature. In Boisvert [Boi97], pages 326–329. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Golub:SCD92

[GO92a] Gene H. Golub and James M.

Ortega. Scientific Computing and Differential Equations. Academic Press, New York, NY, USA, 1992. ISBN 0-12-289255-0. xi + 337 pp. LCCN QA371 .G62 1992. Prepared with LaTEX.

Golub:1992:SCD

[GO92b] Gene H. (Gene Howard) Golub and James M. Ortega. Scientific computing and differential equations: an introduction to numerical methods. Academic Press, New York, NY, USA, 1992. ISBN 0-12-289255-0. xi + 337 pp. LCCN QA371 .G62 1992.

URL http://www.loc.gov/
catdir/description/els032/
91020106.html;   http://www.
loc.gov/catdir/toc/els032/
91020106.html. Rev. ed. of: An

introduction to numerical methods for differential equations / James M. Ortega. 1981.

Golub:1993:SCI

[GO93] Gene H. (Gene Howard) Golub and James M. Ortega. Scientific Computing: An Introduction with Parallel Computing. Academic Press, New York, NY, USA, 1993. ISBN 0-12-289253-4. x + 442 pp. LCCN QA76.58 .G64 1993.

URL http://www.jstor.org/
stable/2153127;   http://
www.loc.gov/catdir/description/
els032/93009408.html;   http:
//www.loc.gov/catdir/toc/
els032/93009408.html.

Golub:1995:WRD

[GO95] Gene H. Golub and James M. Ortega. Wissenschaftliches Rechnen und Differentialgleichungen: Eine Einführung in die numerische Mathematik (German) [English: Scientific computing and differential equations: An introduction to numerical methods], volume 6 of Berliner Studienreihe zur Mathematik [Berlin Study Series on Mathematics]. Heldermann Verlag, Berlin, Germany, 1995. ISBN 3-88538-106-0. xii + 330 pp. LCCN ???? Translated from the 1992 English original by Heinz Dalkowski.

Golub:1996:SC

[GO96a] Gene Golub and James M.

Ortega. Scientific computing: Eine Einführung in das wissenschaftliche Rechnen und Parallele Numerik (German) [An introduction to scientific computing and parallel numerics]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1996. ISBN 3-519-02969-3. 534 pp. URL http://

link.springer.com/openurl?
genre=book\%26isbn=978-3-

519-02969-4. Translated from the 1993 English original [GO93] and with a foreword by Rolf Dieter Grigorieff and Hartmut Schwandt.

Golub:1996:SCD

[GO96b] Gene Howard Golub and James M.

Ortega. Scientific Computing and Differential Equations: an Introduction to Numerical Methods. Academic Press, New York, NY, USA, fourth edition, 1996. ISBN 0-12-289255-0. xi + 337 pp. LCCN QA371 1996. URL http://

                                        www.gbv.de/dms/bowker/toc/
                                        9780122892554.pdf.

Gokoglu:1993:NAI

[Gok93] Suleyman A. Gokoglu. Numerical analysis of an impinging jet reactor for the CVD and gas-phase nucleation of titania. National Technical Information Service, Washington, DC, USA, 1993. ???? pp. Shipping list no.: 94-0808-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1994] 1 microfiche.

Goldstine:1990:RTP

[Gol90a] Herman H. Goldstine. Remembrance of things past. In Nash [Nas90b], pages 5–16. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Golub:1990:PRJ

[Gol90b] Gene Golub. Prologue. reflections on Jim Wilkinson. In Cox and Hammarling [CH90], pages 1–5. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Golberg:1999:BIM

[Gol99] Michael A. Golberg. Boundary integral methods: numerical and mathematical aspects. International series on computational engineering v.1. Computational Mechanics Publications, WIT Press, 1999. ISBN 1-85312-529-6. 374 pp. LCCN TA347.B69 G65 1999.

Gottling:1995:MAC

[Got95] James G. Gottling. Matrix Analysis of Circuits Using MAT-LAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-127044-3. viii + 130 pp. LCCN TK7867.G633 1995.

Greenspan:1991:NM

[GR91] Donald Greenspan and P. Rozsa, editors. Numerical methods, volume 59 of Colloquia Mathematica Societatis Janos Bolyai. North-Holland Publishing Co., Amsterdam, The Netherlands, 1991. ISBN 0-444-98694-4. ISSN 0139-3383. LCCN QA297 .N8618 1991. Papers from the fifth in a series of conferences on numerical methods.

Grossmann:1992:NPD

[GR92] Christian Großmann and Hans-Görg Roos. Numerik partieller Differentialgleichungen. (German) [Numerical solution of partial differential equations]. Teubner Studienbücher Mathematik. [Teubner Mathematical Textbooks]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1992. ISBN 3-519-02089-0. 477 pp.

Grossmann:1994:NPD

[GR94] Christian Großmann and Hans-Görg Roos. Numerik partieller Differentialgleichungen. (German) [Numerical solution of partial differential equations].

Teubner Studienbücher Mathematik. [Teubner Mathematical Textbooks]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, second edition, 1994. ISBN 3-519-12089-5. 477 pp.

Grandine:1990:NMP

[Gra90] Thomas A. Grandine. The Numerical Methods Programming Projects Book. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-853385-3 (hardcover), 0-19-853387-X (paperback). viii + 146 pp. LCCN QA76.6 .G718 1990.

Goossens:1997:LGC

[GRM97] Michel Goossens, Sebastian Rahtz, and Frank Mittelbach. The LaTEX Graphics Companion: Illustrating Documents with TEX and PostScript. Tools and Techniques for Computer Typesetting. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-85469-4. xxi + 554 pp. LCCN Z253.4.L38G663 1997. US$39.75.

Gropp:1990:CPL

[Gro90] W. Gropp. CLAM: a programming language incorporating expert use of linear solution algorithms. In Houstis et al. [HRV90], pages 319–330. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Gropp:1997:WWC

[Gro97a]  W. D. Gropp.  Why we

couldn’t use numerical libraries for PETSc. In Boisvert [Boi97], pages 249–254. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Grosse:1997:RI

[Gro97b] Eric Grosse. Real Inferno. In Boisvert [Boi97], pages 270–279. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

URL ftp://cm.bell-labs.
com/inferno/real.ps.

Goering:1993:FEM

[GRT93] Herbert Goering, Hans-Görg Roos, and Lutz Tobiska. Finite-Element-Methode. (German) [Finite-Element Method]. Akademie-Verlag, Berlin, Germany, third edition, 1993. ISBN 3-05-501564-9. 252 pp. Eine Einführung. [An introduction].

Gear:1990:DOM

[GS90] C. W. Gear and R. D. Skeel. The development of ODE methods: a symbiosis between hardware and numerical analysis. In Nash [Nas90b], pages 88–105. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Griffiths:1991:NMEa

[GS91a] D. V. Griffiths and I. M.

Smith. Numerical methods for engineers: a programming approach. Blackwell Scientific Publications, Oxford, UK, 1991. ISBN 0-632-02753-3 (paperback), 0-632-02751-7 (hard). xiii + 330 pp. LCCN TA335 .G75 1991 Bar.

Griffiths:1991:NMEb

[GS91b] D. V. Griffiths and I. M. Smith. Numerical methods for engineers: a programming approach. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1991. ISBN 0-8493-8610-1. xiii + 330 pp. LCCN TA335 .G75 1991.

Gu:1993:NMS

[Gu93] Yunyun Gu. Numerical methods for singular integration. Thesis (m.s.) in applied mathematics, University of California, Davis, Davis, CA, USA, 1993. 51 pp.

Guclu:1995:RFR

[Guc95] Ceylan Celil Guclu. Radio frequency resonator analysis using analytical, numerical and experimental methods. Thesis (ph.d.) in radiological sciences, University of California, Irvine, Irvine, CA, USA, 1995. xv + 96 pp.

Gutknecht:1990:PDS

[Gut90] Martin H. Gutknecht. The pioneer days of scientific computing in Switzerland. In Nash [Nas90b], pages 301–313. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Ganzha:1996:NSP

[GV96a] Victor G. (Victor Grigorevich) Ganzha and Evgenii V. (Evgenii Vasilevich) Vorozhtsov. Numerical solutions for partial differential equations: problem solving using Mathematica. Symbolic and Numeric Computation Series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1996. ISBN 0-8493-7379-4. xii + 347 pp. LCCN QA377.G2345 1996. Problem solving using ıt Mathematica, With 1 IBM-PC floppy disk (3.5 inch; HD).

Golub:1996:MC

[GV96b] Gene H. Golub and Charles F.

Van Loan. Matrix Computations. Johns Hopkins Studies in the Mathematical Sciences. The Johns Hopkins University Press, Baltimore, MD, USA, third edition, 1996. ISBN 0-8018-5413-X (hardcover), 0-8018-5414-8 (paperback). xxvii + 694 pp. LCCN QA188 .G65 1996. URL

http://www.loc.gov/catdir/
bios/jhu052/96014291.html;
 http://www.loc.gov/catdir/
description/jhu051/96014291.
html.

Golub:1997:CSI

[Gv97] Gene H. Golub and Henk A. van der Vorst. Closer to the solution: Iterative linear solvers. In Duff and Watson [DW97], pages 63– 92. ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

//global.oup.com/academic/
product/the-state-of-the-
art-in-numerical-analysis-
9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Griffiths:1990:NAP

[GW90] D. F. Griffiths and G. A. Watson. Numerical analysis 1989: proceedings of the 13th Dundee Conference, June 1989, volume 228 of Pitman research notes in mathematics. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1990. ISBN 0-470-21586-0. ISSN 0269-3674. 283 pp. LCCN QA297.D85 1989.

Griffiths:1992:NAP

[GW92] D. F. Griffiths and G. A. Watson. Numerical analysis, 1991: proceedings of the 14th Dundee Conference, June 1991, volume 260 of Pitman research notes in mathematics. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1992. ISBN 0-582-08908-5. ISSN 0269-3674. 292 pp. LCCN QA297.D85 1991.

Gerald:1994:ANA

[GW94a] Curtis F. Gerald and Patrick O. Wheatley. Applied numerical analysis. Addison-Wesley, Reading, MA, USA, fifth edition, 1994. ISBN 0-201-56553-6. various pp. LCCN QA 297 G47 1994.

Griffiths:1994:NAP

[GW94b] D. F. Griffiths and G. A. Watson. Numerical analysis 1993: proceedings of the 15th Dundee Conference, June-July 1993, volume 303 of Pitman research notes in mathematics series. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1994. ISBN 0-582-22568-X (paperback). ISSN 0269-3674. 275 pp. LCCN QA297.D85 1993. UK£30.00.

Griffiths:1996:NA

[GW96] D. F. Griffiths and G. A. Watson, editors. Numerical analysis. World Scientific Publishing Co. Inc., River Edge, NJ, USA, 1996. ISBN 981-02-2719-1. x + 370 pp. A. R. Mitchell 75th birthday volume.

Gustafson:1998:ACM

[GW98] Grant B. Gustafson and Calvin H.

Wilcox. Analytical and Computational Methods of Advanced Engineering Mathematics, volume 28 of Texts in applied mathematics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98265-5. xxii + 729 pp. LCCN TA330 .G87 1998. URL http:

//www.springeronline.com/
sgw/cda/frontpage/0,10735,
4-40109-22-1522238-0,00.html.

Gasquet:1999:FAA

[GW99a] Claude Gasquet and Patrick Witomski. Fourier analysis and applications: filtering, numerical computation, wavelets, volume 30 of Texts in applied mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98485-2 (hardcover). xviii + 442 pp. LCCN QA403.5 .G37 1999.

Gerald:1999:ANA

[GW99b] Curtis F. Gerald and Patrick O. Wheatley. Applied numerical analysis. Addison-Wesley, Reading, MA, USA, sixth edition, 1999. ISBN 0-201-87072-X. various pp. LCCN QA297 .G47 1999.

Galligani:1996:EAE

[GZ96] E. Galligani and L. Zanni. Error analysis of elimination methods for equality constrained quadratic programming problems. In Alefeld and Herzberger [AH96], pages 107–112. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Hackbusch:1991:ILG

[Hac91] Wolfgang Hackbusch. Iterative Lösung großer schwachbesetzter Gleichungssysteme. (German) [Iterative solution of large sparse system of equations], volume 69 of Leitfäden der Angewandten Mathematik und Mechanik [Guides to Applied Mathematics and Mechanics]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1991. ISBN 3-519-02372-5. 382 pp. Teubner Studienbücher Mathematik. [Teubner Mathematical Textbooks].

Hackbusch:1992:NTB

[Hac92] W. Hackbusch, editor. Numerical techniques for boundary element methods: proceedings of the Seventh GAMM-Seminar, Kiel, January 25– 27, 1991, volume 33 of Notes on numerical fluid mechanics. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1992. ISBN 3-528-07633-X. ISSN 0179-9614. LCCN TA347.B69 G35 1991.

Hackbusch:1994:ISL

[Hac94] Wolfgang Hackbusch. Iterative solution of large sparse systems of equations, volume 95 of Applied Mathematical Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-94064-2. xxi + 429 pp. LCCN QA1 .A647 vol. 95 QA188. URL

http://www.loc.gov/catdir/
enhancements/fy0812/93002098-

d.html. Translated and revised from the 1991 German original.

Hagstrom:1990:AAD

[Hag90] Thomas Hagstrom. Asymptotic analysis of dissipative waves with applications to their numerical simulation. National Technical Information Service, Washington, DC, USA, August 1990. 13 pp.

Hall:1994:BEM

[Hal94] W. S. Hall. The boundary element method, volume 27 of Solid Mechanics and its Applications. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1994. ISBN 0-7923-2580-X. x + 227 pp.

Harper:1999:NPC

[Har99] D. A. T. Harper. Numerical palaeobiology: computer-based modelling and analysis of fossils and their distributions. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1999. ISBN 0-471-97405-6. x + 468 pp. LCCN QE721.2.D37 N86 1999.

Hatton:1997:EES

[Hat97] L. Hatton. The T -experiments: errors in scientific software. In Boisvert [Boi97], pages 12–31. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Hiebert-Dodd:1990:AET

[HD90] K. L. Hiebert-Dodd. Assisting experts in their data analysis tasks — an AI approach. In Houstis et al. [HRV90], pages 197–206. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Heath:1997:SCI

[Hea97]  Michael T. Heath.  Scien-

tific Computing: An Introductory Survey. McGraw-Hill, New York, NY, USA, 1997. ISBN 0-07-027684-6. xx + 408 pp. LCCN Q183.9.H4 1997.

Heinrich:1996:CTM

[Hei96] S. Heinrich. Complexity theory of Monte Carlo algorithms. In Renegar et al. [RSS96], pages 405–420. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Hench:1992:NMP

[Hen92] John Josef Hench. Numerical methods for periodical linear systems. Thesis (ph.d.), University of California, Santa Barbara, Santa Barbara, CA, USA, 1992. xiv + 166 pp.

Hermann:1994:LON

[Her94] Robert Hermann. Lie-theoretic ODE numerical analysis, mechanics, and differential systems, volume 29 of Interdisciplinary mathematics. Math Sci Press, Brookline, Mass., 1994. ISBN 0-915692-45-7. 286 pp. LCCN QA372 .H485 1994.

Herzberger:1995:WRG

[Her95] Jürgen Herzberger, editor. Wissenschaftliches Rechnen (German) [Scientific Computation]. Akademie-Verlag, Berlin, Germany, 1995. ISBN 3-05-501680-7 (paperback), 3-05-501694-7. 321 pp. URL http:/

/d-nb.info/945248083/04;
http://www.gbv.de/dms/hbz/
toc/ht006736464.pdf. Eine

Einführung in das Scientific Computing. [An introduction to scientific computing].

Herzberger:1996:OCC

[Her96] J. Herzberger. On the R-order of convergence of a class of simultaneous methods of approximating the roots of a polynomial. In Alefeld and Herzberger [AH96], pages 113–119. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455

Herzberger:1998:UNM

[Her98] Jürgen Herzberger. Übungsbuch zur numerischen Mathematik. (German) [Exercise book for numerical mathematics]. Vieweg Mathematik. [Vieweg Mathematics]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1998. ISBN 3-528-06948-1. x + 132 pp. Typische Aufgaben mit ausgearbeiteten Lösungen zur Numerik und zum wissenschaftlichen Rechnen. [Typical problems with worked-out solutions on numerics and on scientific computing].

Hestenes:1990:CG

[Hes90] Magnus R. Hestenes. Conjugacy and gradients. In Nash [Nas90b], pages 167–179. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Hake:1990:PSD

[HH90] J.-F. Hake and W. Homberg. PDE software for distributed computing. In Houstis et al. [HRV90], pages 207–214. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Hammerlin:1991:NM

[HH91] Günther (Günther) Hämmerlin and Karl-Heinz (Karl-Heinz) Hoffmann. Numerical mathematics. Undergraduate texts in mathematics. Readings in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97494-6 (New York), 3-540-97494-6 (Berlin). xi + 422 pp. LCCN QA297 .H2513 1991. URL

http://www.loc.gov/catdir/
enhancements/fy0812/90022552-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0812/
90022552-t.html. Translated

from the German by Larry Schumaker, Readings in Mathematics.

Hammerlin:1994:NM

[HH94] Günther Hämmerlin and Karl-Heinz Hoffmann. Numerische Mathematik. (German) [Numerical mathematics]. Springer-Lehrbuch. [Springer Textbook]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 1994. ISBN 3-540-58033-

6. xiv + 449 pp. Grundwissen Mathematik. [Basic Knowledge in Mathematics].

Hammer:1993:NTV

[HHKR93] R. Hammer, M. Hocks, U. Kulisch, and D. Ratz. Numerical Toolbox for Verified Computing. I: Basic Numerical Problems — Theory, Algorithms, and Pascal-XSC Programs, volume 21 of Springer series in computational mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-57118-3 (Berlin), 0-387-57118-3 (New York). various pp. LCCN QA300 .N86 1993.

Houstis:1990:AMS

[HHRS90] C. E. Houstis, E. N. Houstis, J. R. Rice, and S. M. Samartzis.

The algorithm mapper: a system for modeling and evaluating parallel applications/ architecture pairs. In Houstis et al. [HRV90], pages 87–102. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Hibbs:1993:NMH

[Hib93] Barry Joseph Hibbs. Numerical modeling and hydrologic analysis of floodplain hydrogeology, back storage interactions, and underflow in the Colorado River alluvial aquifer. Thesis (ph.d.), University of Texas at Austin, Austin, TX, USA, 1993. xiii + 329 pp.

Higham:1990:ACD

[Hig90] Nicholas J. Higham. Analysis of the Cholesky decomposition of a semi-definite matrix. In Cox and Hammarling [CH90], pages 161– 185. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Higham:1996:ASN

[Hig96] Nicholas J. Higham. Accuracy and Stability of Numerical Algorithms. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1996. ISBN 0-89871-355-2 (paperback). xxviii + 688 pp. LCCN QA297.H53 1996. US$39.00. Typeset with LaTEX2e.

Higham:1997:RLE

[Hig97a] D. J. Higham. Reliability of local error control algorithms for initial value ordinary differential equations. In Boisvert [Boi97], pages 315–325. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Higham:1997:TLA

[Hig97b] N. J. Higham. Testing linear algebra software. In Boisvert [Boi97], pages 109–124. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Higham:1997:RDD

[Hig97c] Nicholas J. Higham. Recent developments in dense numerical linear algebra. In Duff and Watson [DW97], pages 1– 26. ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

//global.oup.com/academic/
product/the-state-of-the-
art-in-numerical-analysis-
9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Horn:1990:MA

[HJ90] Roger A. Horn and Charles R.

Johnson. Matrix analysis. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-30586-1, 0-521-38632-2 (paperback). xiii + 561 pp. LCCN QA188 .H66 1990.

Horn:1991:TMA

[HJ91] Roger Horn and Charles R. Johnson. Topics in Matrix Analysis. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-30587-X. viii + 607 pp. LCCN QA188 .H664 1991.

Houstis:1990:LPE

[HKP^{+}90] E. N. Houstis, M. Kantzouraki, Papatheodorou, T. S., and V. Sotiropoulou. Logic parallelism in an expert system for solving partial differential equations. In Houstis et al. [HRV90], pages 111–132. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Hanselman:1995:SEM

[HL95] Duane C. Hanselman and Bruce Littlefield. The Student Edition of Matlab: High-Performance Numeric Computation and Visualization Software: Version 4 User’s Guide. The MATLAB curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-184979-4, 0-13-184995-6 (Windows version), 0-13-184987-5 (Macintosh version), 0-13-459207-7 (Macintosh/Power Macintosh version). xii + 833 pp. LCCN QA297 .S8427 1995.

Hanselman:1996:MMC

[HL96] Duane Hanselman and Bruce Littlefield. Mastering MAT-LAB: a comprehensive tutorial and reference. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-191594-0. xii + 542 pp. LCCN QA297.H295 1996.

Hanselman:1997:SEM

[HL97] Duane C. Hanselman and Bruce Littlefield. The student edition of MATLAB: version 5, user’s guide. The MATLAB curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-272550-9 (paperback user’s guide), 0-13-272485-5 (user’s guide), 0-13-272477-4 (set), 0-13-022598-3 (CD-ROM). xxxix + 429 pp. LCCN QA297 .S8436 1997.

Hanselman:1998:MMC

[HL98] Duane C. Hanselman and Bruce Littlefield. Mastering MAT-LAB 5: a comprehensive tutorial and reference. The MAT-LAB curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-858366-8. xviii + 638 pp. LCCN QA297.H296 1998.

Herzberger:1996:OOC

[HM96] J. Herzberger and L. Metzner.

On the Q-order and R-order of convergence for coupled sequences arising in iterative numerical processes. In Alefeld and Herzberger [AH96], pages 120–131. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Heywood:1990:NSE

[HMRS90] John G. Heywood, Kyûya Masuda, Reimund Rautmann, and Vsevolod A. Solonnikov, editors. The Navier–Stokes Equations Theory and Numerical Methods: Proceedings of a Conference held at Oberwolfach, FRG, Sept. 18– 24, 1988, volume 1431 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. CO-DEN LNMAA2. ISBN 0-387-52770-2, 3-540-52770-2 (print), 3-540-47141-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L35

v.1431. URL http://link.
springer.com/book/10.1007/
BFb0086050;   http://www.
springerlink.com/content/
978-3-540-47141-7.

Heywood:1992:NSE

[HMRS92] John G. Heywood, Kyûya Masuda, Reimund Rautmann, and Vsevolod A. Solonnikov, editors. The Navier–Stokes Equations II — Theory and Numerical Methods: Proceedings of a Conference held in Oberwolfach, Germany, August 18–24, 1991, volume 1530 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 0-387-56261-3, 3-540-56261-3 (print), 3-540-47498-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L35

v.1530. URL http://link.

springer.com/book/10.1007/
BFb0090329;   http://www.
springerlink.com/content/
978-3-540-47498-2.

Hawla:1990:KAE

[HN90] D. L. Hawla and N. Neishlos.

Knowledge acquisition and efficient use of engineering software and finite element example. In Houstis et al. [HRV90], pages 287–300. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Ho:1992:NAI

[Ho92] H. Ho. Numerical analysis of the Iosipescu specimen for composite materials. National Technical Information Service, Washington, DC, USA, 1992. ???? pp. Shipping list no.: 93-0034-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration], 1992. 1 microfiche.

Ho:1994:NNA

[Ho94] Henjen Ho. Non-linear numerical analysis of the Iosipescu specimen for composite materials. National Technical Information Service, Washington, DC, USA, 1994. ???? pp. Shipping list no.: 94-0866-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1994] 1 microfiche.

Hoffman:1992:NME

[Hof92] Joe D. Hoffman. Numerical methods for engineers and scientists. McGraw-Hill, New York, NY, USA, 1992. ISBN 0-07-029213-2. xviii + 825 pp. LCCN QA297 .H588 1992.

Holland:1997:NMI

[Hol97] Thomas Joseph Holland. Numerical methods for implementing the bounding surface plasticity model for clays. Thesis (ph.d.) in civil engineering, University of California, Davis, Davis, CA, USA, 1997.

Honerkamp:1993:SDS

[Hon93] J. Honerkamp. Stochastic dynamical systems: concepts, numerical methods, data analysis. VCH, New York, NY, USA, 1993. ISBN 1-56081-563-9. xiii + 226 pp. LCCN QA274.H6613 1993.

Honerkamp:1994:SDS

[Hon94] J. Honerkamp. Stochastic dynamical systems: concepts, numerical methods, data analysis. VCH, New York, NY, USA, 1994. ISBN 1-56081-563-9, 3-527-89563-9. xvi + 535 pp. LCCN QA274 .H6613 1994.

Hall:1990:NAP

[HP90] Charles A. Hall and Thomas A. Porsching. Numerical analysis of partial differential equations. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1990. ISBN 0-13-626557-X. xv + 319 pp. LCCN QA374 .H29 1990.

Hirsch:1992:CMA

[HPO92] Ch Hirsch, Jacques Periaux, and E. Onate. Computational methods in applied sciences: invited lectures and special technological sessions of the First European Computational Fluid Dynamics Conference and the First European Conference on Numerical Methods in Engineering, 7–11 September 1992, Brussels, Belgium. Elsevier, Amsterdam, The Netherlands, 1992. ISBN 0-444-89795-X. vii + 305 pp. LCCN TA357 .E88 1992.

Han:1999:PMT

[HR99] Weimin Han and B. Dayanand Reddy. Plasticity: mathematical theory and numerical analysis, volume 9 of Interdisciplinary applied mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98704-5 (hardcover). xiii + 371 pp. LCCN QA931 .H25 1999.

Hromadka:1993:BAM

[Hro93] Theodore V. Hromadka, II.

The best approximation method in computational mechanics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-19798-2 (Berlin), 0-387-19798-2 (New York). xii + 250 pp. LCCN QA320 .H76 1993.

Houstis:1990:PIE

[HRP90] E. N. Houstis, J. R. Rice, and T. S. Papatheodorou. Parallel (II) Ellpack: An expert system for parallel processing of partial differential equations. In Houstis et al. [HRV90], pages 63–74. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Houstis:1990:IMS

[HRV90] E. N. Houstis, John Rischard Rice, and Robert Vichnevetsky, editors. Intelligent mathematical software systems: proceedings of the First IMACS/IFAC International Conference on Expert Systems for Numerical Computing, Purdue University, U.S.A., 5–7 December, 1988. North-Holland Publishing Co., Amsterdam, The Netherlands, 1990. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Hebeker:1994:NMN

[HRW94] Friedrich-Karl Hebeker, Rolf Rannacher, and Gabriel Wittum, editors. Numerical methods for the Navier–Stokes equations: proceedings of the international workshop held at Heidelberg, October 25–28, 1993, volume 47 of Notes on numerical fluid mechanics. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1994. ISBN 3-531-07647-7, 3-528-07647-X. ISSN 0179-9614. LCCN QA 929 N85 1993.

Hardle:1993:CIM

[HS93] Wolfgang Härdle and Léopold Simar, editors. Computer Intensive Methods in Statistics. Statistics and Computing. Physica-Verlag, Vienna, Austria, 1993. ISBN 3-7908-0677-3. ISSN 1431-8784. vii + 176 pp. LCCN QA273.A1-274.9; QA274-274.9.

Hostetter:1991:ANC

[HSDM91] G. H. Hostetter, Mohammed S.

Santina, and Paul D’Carpio-Montalvo. Analytical, numerical, and computational methods for science and engineering. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1991. ISBN 0-13-026055-X. xxi + 488 pp. LCCN Q183.9 .H67 1991. US$48.00.

Hestenes:1991:NIN

[HT91] Magnus Rudolph Hestenes and John Todd. NBS-INA, the Institute for Numerical Analysis, UCLA 1947–1954. National Technical Information Service, Washington, DC, USA, August 1991. various pp. Cover title: Mathematicians learning to use computers: the Institute for Numerical Analysis, UCLA, 1947– 1954. Spine title: The Institute for Numerical Analysis, UCLA, 1947–1954. Sponsored in part by the Mathematical Association of America. Shipping list no.: 92-1306-M.

Hudak:1991:AMR

[Hud91] Paul F. Hudak. Analytical methods for regional-scale ground water monitoring network design and numerical verification. Thesis (ph.d.) in geography, University of California, Santa Barbara, Santa Barbara, CA, USA, December 14, 1991. xvii + 388 pp.

Hirsch:1992:NME

[HZO92] Ch Hirsch, O. C. Zienkiewicz, and E. Onate. Numerical methods in engineering ’92: proceedings of the First European Conference on Numerical Methods in Engineering, 7–11 September 1992, Brussels, Belgium. Elsevier, Amsterdam, The Netherlands, 1992. ISBN 0-444-89794-1. xv + 879 pp. LCCN TA335 .E97 1992 Bar.

Isaacson:1994:ANM

[IK94] Eugene Isaacson and Herbert Bishop Keller. Analysis of numerical methods. Dover, New York, NY, USA, 1994. ISBN 0-486-68029-0 (paperback). xv + 541 pp. LCCN QA297 .I8 1994. URL

                                        http://www.loc.gov/catdir/
                                        description/dover031/94007740.
                                        html.

Ionescu:1993:FNM

[IS93] Ioan R. Ionescu and Mircea Sofonea. Functional and numerical methods in viscoplasticity. Oxford science publications. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1993. ISBN 0-19-853590-2. xvii + 265 pp. LCCN QA931 .I55 1993. US$67.50.

Isaacson:1990:OMC

[Isa90] Eugene Isaacson. The origin of mathematics of computation and some personal recollections. In Nash [Nas90b], pages 211– 216. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Iserles:1996:FCN

[Ise96] Arieh Iserles. A first course in the numerical analysis of differential equations. Cambridge Texts in Applied Mathematics. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-55376-8 (hardcover), 0-521-55655-4 (paperback). xvi + 378 pp. LCCN QA371 .I813 1996.

URL http://www.loc.gov/
catdir/description/cam028/
96139924.html;   http://www.
loc.gov/catdir/toc/cam027/
96139924.html.

Iserles:1997:BCT

[Ise97]  Arieh Iserles.  Beyond the

classical theory of computational ordinary differential equations. In Duff and Watson [DW97], pages 171–192. ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

//global.oup.com/academic/
product/the-state-of-the-
art-in-numerical-analysis-
9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Ingham:1994:BEM

[IY94] D. B. Ingham and Y. Yuan. The boundary element method for solving improperly posed problems, volume 19 of Topics in Engineering. Computational Mechanics Publications, Southampton, UK, 1994. ISBN 1-85312-291-2. xiv + 145 pp. With a contribution by H. Han.

Iserles:1996:QNA

[IZ96] Arieh Iserles and Antonella Zanna. Qualitative numerical analysis of ordinary differential equations. In Renegar et al. [RSS96], pages 421–442. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Jameson:1994:ADNa

[Jam94a] Antony Jameson. Analysis and design of numerical schemes for gas dynamics 1: artificial diffusion, upwind biasing, limiters and their effect on accuracy and multigrid convergence. National Technical Information Service, Washington, DC, USA, 1994. ???? pp. Shipping list no.: 95-0062-M.

Jameson:1994:ADNb

[Jam94b] Antony Jameson. Analysis and design of numerical schemes for gas dynamics 2: artificial diffusion, discrete shock structure. National Technical Information Service, Washington, DC, USA, 1994. ???? pp. Shipping list no.: 95-0062-M.

Johnson:1994:NCM

[JB94] Michael L. Johnson and Ludwig Brand. Numerical computer methods. Part B, volume 240 of Methods in enzymology. Academic Press, New York, NY, USA, 1994. ISBN 0-12-182141-2. ISSN 0076-6879. xxviii + 857 pp. LCCN QP601 .C56 v.240.

Jennings:1977:MCE

[Jen77] Alan Jennings. Matrix Computation for Engineers and Scientists. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1977. ISBN 0-471-99421-9 (hardcover), 0-471-27832-7 (paperback). xv + 330 pp. LCCN TA347 .D4J461.

Jeffreys:1946:MMP

[JJ46] Sir Harold Jeffreys and Bertha Swirles Jeffreys. Methods of mathematical physics. Cambridge University Press, Cambridge, UK, 1946. vii + 1 + 679 pp. LCCN QA401 .J4 1946.

Jeffreys:1956:MMP

[JJ56] Sir Harold Jeffreys and Bertha Swirles Jeffreys. Methods of mathematical physics. Cambridge University Press, Cambridge, UK, third edition, 1956. 714 pp. LCCN QA401 .J4 1956.

Jeffreys:1999:MMP

[JJ99] Sir Harold Jeffreys and Bertha Swirles Jeffreys. Methods of mathematical physics. Cambridge mathematical library. Cambridge University Press, Cambridge, UK, third edition, 1999. ISBN 0-521-66402-0 (paperback). viii + 718 pp. LCCN QA401 .J4 1999.

URL http://www.loc.gov/
catdir/description/cam022/
2001265472.html;   http://
www.loc.gov/catdir/toc/cam027/
2001265472.html.

Johnson:1999:NNM

[JKB99] M. R. Johnson, G. J. Kearley, and H. G. Buttner. Neutrons and numerical methods — N2M: Grenoble, France, December 1998, volume 479 of AIP conference proceedings. American Institute of Physics, Woodbury, NY, USA, 1999. ISBN 1-56396-838-X. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465. ix + 238 pp. LCCN QC793.5.N462 N484 1999. Sponsored by: Institut Laue-Langevin and Molecular Simulations, Inc.

Jacobs:1990:ESS

[JM90] D. (David) A. H. Jacobs and G. Markham. Experiences with some software engineering practices in numerical software. In Cox and Hammarling [CH90], pages 277–296. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Jennings:1992:MC

[JM92] Alan Jennings and J. J. McKeown. Matrix computation for engineers and scientists. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, second edition, 1992. ISBN 0-471-93527-1. xiv + 427 pp. LCCN TA347.D4 J46 1992. £19.95.

Joly:1990:MOM

[Jol90] Pascal Joly. Mise en œuvre de la méthode des éléments finis. (French) [Implementation of the finite-element method], volume 2 of Mathématiques & Applications (Paris) [Mathematics and Applications]. Ellipses, Paris, France, 1990. 157 pp.

Ji:1996:TIF

[JP96] Jun Ji and Florian A. Potra.

Tapia indicators and finite termination of infeasible-interior-point methods for degenerate LCP. In Renegar et al. [RSS96], pages 443–454. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

James:1992:PPN

[JR92] K. R. James and W. O. Riha.

Pascal programming and numerical methods for scientists and engineers. A Chartwell-Bratt student text. Studentlitteratur, Lund, Sweden, 1992. ISBN 0-86238-303-X, 91-44-36061-4 (Studentlitteratur). vii + 208 pp. LCCN QA76.73.P2 .J357 1992 Bar.

James:1993:ANM

[JSW93] M. L. James, G. M. Smith, and J. C. Wolford. Applied Numerical Methods for Digital Computation. HarperCollins College Publishers, New York, NY, USA, fourth edition, 1993. ISBN 0-06-500494-9. x + 709 pp. LCCN QA297.J3 1993. US$87.41.

Ju:1995:NMS

[Ju95] J. W. Ju, editor. Numerical methods in structural mechanics: presented at the 1995 Joint ASME Applied Mechanics and Materials Summer Meeting, Los Angeles, California, June 28– 30, 1995, volume 204 of AMD. American Society of Mechanical Engineers, United Engineering Center, 345 E. 47th St., New York, NY 10017, USA, 1995. ISBN 0-7918-1319-3. LCCN TA640.2 .N85 1995.

Judd:1998:NME

[Jud98] Kenneth L. Judd. Numerical methods in economics. MIT Press, Cambridge, MA, USA, 1998. ISBN 0-262-10071-1 (hardcover). xiii + 633 pp. LCCN HB137 .J83 1998.

Kalis:1996:PDS

[Kal96]  H. Kalis.  Parasto difer-

enciālvienādojumu, skaitliskās metodes. (Latvian) [Ordinary differential equations, numerical methods]. Latv. Univ., Riga, Latvia, 1996. 80 pp.

Kincaid:1991:NA

[KC91a] David Kincaid and Ward Cheney. Numerical analysis. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1991. ISBN 0-534-13014-3. x + 690 pp. Mathematics of scientific computing.

Kincaid:1991:NAM

[KC91b] David (David Ronald) Kincaid and E. W. (Elliott Ward) Cheney. Numerical analysis: mathematics of scientific computing. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1991. ISBN 0-534-13014-3. viii + 690 pp. LCCN QA297 .K563 1991.

Kincaid:1993:SMN

[KC93] David Kincaid and Ward Cheney. Solution Manual for Numerical Analysis: Mathematics of Scientific Computing. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1993. 219 pp.

Kincaid:1996:NAM

[KC96] David (David Ronald) Kincaid and E. W. (Elliott Ward) Cheney. Numerical analysis: mathematics of scientific computing. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, second edition, 1996. ISBN 0-534-33892-5. xii + 804 pp. LCCN QA297 .K563 1996.

Kushner:1992:NMS

[KD92] Harold J. (Harold Joseph) Kushner and Paul Dupuis. Numerical methods for stochastic control problems in continuous time, volume 24 of Applications of mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97834-8 (New York), 3-540-97834-8 (Berlin). ix + 439 pp. LCCN QA402.37 .K87 1992.

Keller:1993:NMT

[Kel93]  Herbert B. Keller.  Nu-

merical Methods for Two-Point Boundary-Value Problems. Dover, New York, NY, USA, 1993. ISBN 0-486-66925-4. xii + 397 pp. Corrected and expanded republication of work first published by Blaisdell, Waltham, MA, 1964.

Keller:1996:QMC

[Kel96] Alexander Keller. Quasi-Monte Carlo methods in computer graphics: The global illumination problem. In Renegar et al. [RSS96], pages 455–470. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Ketkar:1999:NTA

[Ket99] Satish P. Ketkar. Numerical thermal analysis. American Society of Mechanical Engineers, United Engineering Center, 345 E. 47th St., New York, NY 10017, USA, 1999. ISBN 0-7918-0073-3. viii + 103 pp. LCCN TJ260 .K38 1999.

Kutler:1997:FIC

[KFC97] Paul Kutler, J. Flores, and J. J. Chattot. Fifteenth International Conference on Numerical Methods in Fluid Dynamics: proceedings of the conference held in Monterey, CA, USA, 24–28 June 1996, volume 490 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-63054-6 (Berlin). ISSN 0075-8450 (print), 1616-6361 (electronic). xvi + 655 pp. LCCN QA911 .I58 1996.

Kaper:1991:AAN

[KG91] H. G. Kaper and Marc Garbey, editors. Asymptotic analysis and the numerical solution of partial differential equations, volume 130 of Lecture notes in pure and applied mathematics. Marcel Dekker, Inc., New York, NY, USA, 1991. ISBN 0-8247-8538-X. LCCN QA377 .A84 1991. Proceedings of a workshop … held at Argonne National Laboratory, February 26–28, 1990.

Kaper:1993:ANM

[KG93] H. G. Kaper and Marc Garbey, editors. Asymptotic and numerical methods for partial differential equations with critical parameters, volume 384 of NATO ASI series. Series C, Mathematical and physical sciences. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1993. ISBN 0-7923-2061-1. LCCN QA377.N37 1992. Published in cooperation with NATO Scientific Affairs Division. Proceedings of the NATO Advanced Workshop on Asymptotic-Induced Numerical Methods for Partial Differential Equations, Critical Parameters, and Domain Decomposition, Beaune, France, May 25– 28, 1992.

Klatte:1993:CXC

[KKL^{+}93] Rudi Klatte, U. Kulisch, C. Lawo, M. Rauch, and A. Wiethoff. C-XSC: a C++ class library for extended scientific computing. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-56328-8 (Berlin), 0-387-56328-8 (New York). xii + 269 pp. LCCN QA76.73.C153 C9 1993. DM74.00.

Klatte:1991:PXP

[KKN^{+}91a] R. Klatte, U. Kulisch, M. Neaga, D. Ratz, and Ch. Ullrich. PASCAL-XSC: A PASCAL Extension for Scientific Computation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-55137-9 (New York), 3-540-55137-9 (Berlin). x + 344 pp. LCCN QA76.73.P2 P4213 1992. DM64.00. English translation of [KKN^{+}91b].

Klatte:1991:PSB

[KKN^{+}91b] R. Klatte, U. Kulisch, M. Neaga, D. Ratz, and Ch. Ullrich. PASCAL-XSC: Sprachbeschreibung mit Beispielen. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-53714-7. x + 345 pp. LCCN QA76.73.P2 P42 1991. English translation is available as [KKN^{+}91a].

Klatte:1992:PXP

[KKN^{+}92] Rudi Klatte, Ulrich Kulisch, Michael Neaga, Dietmar Ratz, and Christian Ullrich. PASCAL-XSC: language reference with examples. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-540-55137-9 (Berlin), 0-387-55137-9 (New York). x + 344 pp. LCCN QA76.73.P2 P4213 1992. DM 64.00. Translated by G. F. Corliss and others.

Karchevskii:1990:MVC

[KLP90] M. M. Karchevskiĭ, A. D.

Lyashko, and M. F. Pavlova. Metody vychislenii: chislennye metody resheniya differentsialnykh uravnenii. (Russian) [Computational methods: numerical methods for solutions of differential equations]. Kazanskiĭ Gosudarstvennyĭ Universitet, Kazan′, Tatarstan, 1990. ISBN 5-7464-0319-9. 127 pp.

Kreinovich:1998:CCF

[KLRK98] Vladik Kreinovich, Anatoly V.

Lakeyev, Jirí Rohn, and P. Kahl. Computational complexity and feasibility of data processing and interval computations, volume 10 of Applied optimization. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1998. ISBN 0-7923-4865-6 (hardcover). xii + 459 pp. LCCN QA267.7 .C68 1998. URL

http://www.loc.gov/catdir/
enhancements/fy0814/97043085-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0814/
97043085-t.html.

King:1990:KFC

[KMRT90] D. King, T. Mullen, B. Rice, and P. Topping. Knowledge-based FFT and convolution synthesis for any number of points. In Houstis et al. [HRV90], pages 143–154. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Kockler:1994:NMS

[Köc94] Norbert Köckler. Numerical methods and scientific computing: using software libraries for problem solving. Oxford Science Publications. Clarendon Press, New York, NY, USA, 1994. ISBN 0-19-859698-7. xviii + 328 pp. LCCN TA345 .K653 1994. US$48.50. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0637/94008064-
                                        d.html;   http://www.loc.
                                        gov/catdir/enhancements/fy0637/
                                        94008064-t.html. Using soft-

ware libraries for problem solving.

Koeber:1996:ISI

[Koe96] M. Koeber. Inclusion of solutions of initial value problems for quasilinear hyperbolic equations. In Alefeld and Herzberger [AH96], pages 132–137. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Konovalov:1993:VVV

[Kon93] A. N. Konovalov. Vvedenie v vychislitelnye metody lineinoi algebry. (Russian) [Introduction to computational methods of linear algebra]. VO “Nauka”, Novosibirsk, Russia, 1993. ISBN 5-02-030307-0. 159 pp.

Kloeden:1992:NSS

[KP92] Peter E. Kloeden and Eckhard Platen. Numerical Solution of Stochastic Differential Equations, volume 23 of Applications of Mathematics, Stochastic Modelling and Applied Probability. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-642-08107-X. ISSN 0172-4568. xxxvi + 636 pp. LCCN QA273.A1-274.9; QA274-274.9.

Kim:1999:NMD

[KP99] A. V. Kim and V. G. Pimenov. Numerical methods for delay differential equations, volume 44 of Lecture Notes Series. Seoul National University, Research Institute of Mathematics, Global Analysis Research Center, Seoul, Korea, 1999. x + 96 pp. Application of i-smooth calculus.

Kloeden:1994:NSS

[KPS94] Peter E. Kloeden, Eckhard Platen, and Henri Schurz. Numerical Solution of SDE Through Computer Experiments. Universitext. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-57074-8. ISSN 0172-5939 (print), 2191-6675 (electronic). LCCN QA273.A1-274.9; QA274-274.9.

Kress:1998:NA

[Kre98] Rainer Kress. Numerical analysis, volume 181 of Graduate Texts in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98408-9 (hardcover). xii + 326 pp. LCCN QA297 .K725 1998. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0815/97043748-
                                        d.html;   http://www.loc.
                                        gov/catdir/enhancements/fy0815/
                                        97043748-t.html.

Knupp:1993:FGG

[KS93] Patrick M. Knupp and Stanly Steinberg. The fundamentals of grid generation. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1993. ISBN 0-8493-8987-9. xvii + 286 pp. LCCN QA377 .K58 1993. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0646/93011639-
                                        d.html.

Knupp:1994:FGG

[KS94] Patrick Knupp and Stanly Steinberg. Fundamentals of grid generation. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1994. ISBN 0-8493-8987-9. xviii + 286 pp. With 1 IBM-PC floppy disk (3.5 inch; HD).

Kruse:1991:UVK

[KSH91] Rudolf Kruse, E. Schwecke, and J. Heinsohn. Uncertainty and vagueness in knowledge based systems: numerical methods. Artificial intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 3-540-54165-9 (Berlin), 0-387-54165-9 (New York). xi + 491 pp. LCCN Q335 .K75 1991.

Kowalski:1995:STA

[KSS95] Marek A. Kowalski, Krzysztof A.

Sikorski, and Frank Stenger. Selected Topics in Approximation and Computation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1995. ISBN 0-19-508059-9. xiv + 349 pp. URL http:/

/site.ebrary.com/lib/utah/
Doc?id=10087215.

Krommer:1994:NIA

[KU94] Arnold R. Krommer and Christoph W.

Ueberhuber. Numerical integration on advanced computer systems, volume 848 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNCSD9. ISBN 3-540-58410-2 (Berlin), 0-387-58410-2 (New York). ISSN 0302-9743. xiii + 341 pp. LCCN QA299.3 .K76 1994. DM72.00. URL http://

link.springer-ny.com/link/
service/series/0558/tocs/
t0848.htm;   http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=848.

Korn:1996:FVL

[KU96] C. Falcó Korn and C. P. Ullrich. Fast verification of linear system solutions. In Alefeld and Herzberger [AH96], pages 80–106. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Kublanovskaya:1990:CLM

[Kub90] V. N. Kublanovskaya. The contribution of Leningrad mathematicians to the development of numerical linear algebra, 1950– 1986. In Nash [Nas90b], pages 329–338. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Konno:1990:IVD

[KUIO90] C. Konno, Y. Umetani, M. Igai, and T. Ohta. Interactive/visual DEQSOL: Interactive creation, debugging, diagnosis and visualization of numerical simulation. In Houstis et al. [HRV90], pages 301–318. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Kulisch:1996:NAA

[Kul96] Ulrich Kulisch. Numerical algorithms with automatic result verification. In Renegar et al. [RSS96], pages 471–502. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Kulisch:1997:XTE

[Kul97] U. Kulisch. The XSC tools for extended scientific computing. In Boisvert [Boi97], pages 280– 286. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Konstantinov:1997:SMM

[KV97] Mikhail Mikhaĭlov Konstantinov and Nikolaĭ Lyubomirov Vlchanov. Svremenni matematicheski metodi za kompyutrni presmyataniya. Chast 2. (Bulgarian [Modern mathematical methods for computer calculations. Part 2]), volume 2 of Studii na BIAP. Matematicheski Nauki [Studies of the Bulgarian Institute of Analysis and Research. Mathematical Sciences]. Fondatsiya “Blgarski Institut za Analizi i Prouchvaniya”, Sofia, Bulgaria, 1997. ISBN 954-8949-02-4. 288 pp. Chislena lineina algebra. [Numerical linear algebra].

Kaashoek:1990:PIS

[KvSR90] M. A. Kaashoek, J. H. van Schuppen, and A. C. M. Ran, editors. Proceedings of the International Symposium MTNS-89, volume 5 of Progress in systems and control theory. Birkhäuser Verlag, Basel, Switzerland, 1990. ISBN 0-8176-3470-3 (vol. 2), 0-8176-3471-1 (vol. 3), 0-8176-3468-1 (set). LCCN QA402 .P766 1990 v. 1–3 (c1990). The 1989 International Symposium on the Mathematical Theory of Networks and Systems (MTNS) … held in Amsterdam, The Netherlands, June 19–23, 1989. Contents: vol. 1. Realization and modelling in system theory — vol. 2. Robust control of linear systems and nonlinear control — vol. 3. Signal processing, scattering and operator theory, and numerical methods.

Kowalski:1994:NAM

[KW94] Jan Krzysztof Kowalski and Andrzej Wakulicz. Numerical analysis and mathematical modelling, volume 29 of Banach Center publications. Institute of Mathematics, Polish Academy of Sciences, Warszawa, Poland, 1994. ISSN 0137-6934. 304 pp. LCCN QA297 .N842 1994.

Kythe:1995:IBE

[Kyt95] Prem K. Kythe. An introduction to boundary element methods. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1995. ISBN 0-8493-7377-8. xvi + 368 pp. LCCN TA347.B69 K98 1995. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0744/95007136-

d.html. With 1 IBM-PC floppy disk (3.5 inch; HD).

Korbicz:1991:WMN

[KZSU91] Józef Korbicz, Zbigniew Zajda, Włodzimierz Solnik, and Dariusz Uciński. Wybrane metody numerycznego rozwia̧zywania równań różniczkowych cza̧stkowych. (Polish) [Some numerical methods of solving partial differential equations]. Wydawnictwo Wyższej Szkoły Inżynierskiej, Zielona Góra, Poland, 1991. 208 pp.

Lacasse:1995:IMS

[Lac95] James M. Lacasse. Implementation of a multiblock sensitivity analysis method in numerical aerodynamic shape optimization: abstract. National Technical Information Service, Washington, DC, USA, 1995. ???? pp. Shipping list no.: 96-0197-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1995] 1 microfiche.

Laevskii:1993:MKE

[Lae93]  Yu. M. Laevskiĭ.  Metod

konechnykh elementov resheniya mnogomernykh parabolicheskikh uravnenii. (Russian) [Finiteelement method for solving multidimensional parabolic equations]. Novosibirsk. Gos. Univ., Novosibirsk, Russia, 1993. ISBN 5-230-13543-3. 104 pp.

Laevskii:1999:MKE

[Lae99] Yu. M. Laevskiĭ. Metod konechnykh elementov (osnovy teorii, zadachi). (Russian) [Finiteelement method (basic theory, problems)]. Novosibirsk. Gos. Univ., Novosibirsk, Russia, 1999. 167 pp.

Lambert:1991:NMO

[Lam91] J. D. (John Denholm) Lambert.

Numerical methods for ordinary differential systems: the initial value problem. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1991. ISBN 0-471-92990-5. x + 293 pp. LCCN QA328 .L36 1991. US$45.90.

Lamb:1993:IHR

[Lam93] Susan A. Lamb. IRAS high resolution studies and modeling of closely interacting galaxies: galaxy collisions:infrared observations and analysis of numerical models and UV spectroscopy of massive young stellar populations in interacting galaxies. National Technical Information Service, Washington, DC, USA, 1993. ???? pp. Shipping list no.: 94-0111-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1994] 1 microfiche.

Lanczos:1996:LDO

[Lan96] Cornelius Lanczos. Linear differential operators, volume 18 of Classics in Applied Mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1996. ISBN 0-89871-370-6 (paperback), 1-61197-118-7 (e-book). xviii + 564 pp. LCCN QA432 .L3 1996. Corrected reprint of the 1961 original.

Lange:1999:NAS

[Lan99a] Kenneth Lange. Numerical analysis for statisticians. Statistics and Computing. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-94979-8 (hardcover). xv + 356 pp. LCCN QA297 .L34 1999. URL

http://www.loc.gov/catdir/
enhancements/fy0815/98016688-
d.html;   http://www.loc.

gov/catdir/enhancements/fy0815/
98016688-t.html.

Langtangen:1999:CPD

[Lan99b] Hans Petter Langtangen. Computational partial differential equations: numerical methods and Diffpack programming, volume 2 of Lecture notes in computational science and engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-65274-4. xxiii + 682 pp. LCCN QA377 .L355 1999.

Lund:1992:SMQ

[LB92] J. Lund and K. L. Bowers.

Sinc Methods for Quadrature and Differential Equations. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1992. ISBN 0-89871-298-X. x + 304 pp. LCCN QA372 .L86 1992.

LeBorne:1993:NAP

[LeB93] Richard Charles LeBorne. Numerical analysis of the a posteriori recursive least squares lattice prediction algorithm. Thesis (ph.d.), Department of Mathematics, University of California, San Diego, San Diego, CA, USA, 1993. xii + 90 pp.

Lee:1994:ASF

[Lee94] J. Lee. An analysis of supersonic flows with low-Reynolds number compressible two-equation turbulence models using LU finite volume implicit numerical techniques. NASA, Washington, DC, USA, 1994. ???? pp. Shipping list no.: 94-0863-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1994] 1 microfiche.

Lee:1997:AVZ

[Lee97] Kenrick Hugh Lee. Analysis of vadose zone tritium transport from an underground storage tank release using numerical modeling and geostatistics. Thesis (ph.d.) in engineeringmaterials science and mineral engineering, University of California, Berkeley, Berkeley, CA, USA, December 1997.

LeVeque:1990:NMC

[LeV90] Randall J. LeVeque. Numerical methods for conservation laws. Lectures in mathematics ETH Zurich. Birkhäuser Verlag, Basel, Switzerland, 1990. ISBN 3-7643-2464-3, 0-8176-2464-3 (U.S.). 214 pp. LCCN QA377 .L4157 1990.

LeVeque:1992:NMC

[LeV92] Randall J. LeVeque. Numerical methods for conservation laws. Lectures in mathematics ETH Zurich. Birkhäuser Verlag, Basel, Switzerland, second edition, 1992. ISBN 3-7643-2723-5 (Basel), 0-8176-2723-5 (Boston). 214 pp. LCCN QA377 .L4157 1992.

Lawson:1990:LAP

[LG90] Charles L. Lawson and Kajal K. Gupta. The Lanczos algorithm for a pure imaginary Hermitian matrix. In Cox and Hammarling [CH90], pages 25–34. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Lokutsievskii:1995:NCA

[LG95] O. V. Lokutsievskiĭ and M. B.

Gavrikov. Nachala chislennogo analiza. (Russian) [Beginning numerical analysis]. TOO “Yanus”, Moscow, Russia, 1995. ISBN 5-88929-004-5. 581 pp. With a preface by V. M. Tikhomirov.

Li:1990:NME

[Li90] Zi-Cai Li. Numerical methods for elliptic problems with singularities: boundary methods and nonconforming combinations. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1990. ISBN 981-02-0292-X. xv + 258 pp. LCCN QA377 .L49 1990.

Light:1991:ANA

[Lig91] W. A. Light, editor. Advances in numerical analysis, Oxford science publications. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1991. ISBN 0-19-853438-8 (vol. 1), 0-19-853439-6 (vol. 2). LCCN QA297 .A38 1991 v.1-2 (1991-1992). US$45.00. Proceedings of the fourth Summer School in Numerical Analysis, which was held at Lancaster University from 15th July to 3rd August 1990 … sponsored by the Science and Engineering Research Council of Great Britain. Contents: vol. 1. Nonlinear partial differential equations and dynamical systems — vol. 2. Wavelets, subdivision algorithms, and radial basis functions.

Lippsmeyer:1994:NCA

[Lip94] Jeff Lippsmeyer. Numerical control and analysis of the RPV-1. Thesis (m.s.) in mechanical engineering, University of California, Davis, Davis, CA, USA, 1994.

Liu:1990:HNM

[Liu90] Jian-Guo Liu. Homogenization and numerical methods for hyperbolic equations. Thesis (ph.d.), University of California, Los Angeles, Los Angeles, CA, USA, 1990. ix + 97 pp.

Liu:1993:NHO

[Liu93] Xu-Dong Liu. Nonoscillatory high order methods for the numerical solution of conservation laws. Thesis (ph.d.), University of California, Los Angeles, Los Angeles, CA, USA, 1993. viii + 107 pp.

Lorentz:1997:MLA

[LL97] G. G. Lorentz and Rudolph A. Lorentz. Mathematics from Leningrad to Austin: George G. Lorentz’ selected works in real, functional, and numerical analysis. Contemporary mathematicians. Birkhäuser Verlag, Basel, Switzerland, 1997. ISBN 0-8176-3710-9 (vol. 1: hardcover), 3-7643-3710-9 (vol. 1: Basel: hardcover), 0-8176-3922-5 (vol. 2: hardcover), 3-7643-3922-5 (vol. 2: Basel: hardcover), 0-8176-3923-3 (set: hardcover), 3-7643-3923-3 (set: Basel: hardcover). xxxvi + 548 pp. LCCN QA300.5 .L67 1997 v.1-2 (1997). English and German.

Liou:1991:CNM

[LM91] W. W. Liou and P. J. Morris. A comparison of numerical methods for the Rayleigh equation in unbounded domains. Institute for Computational Mechanics in Propulsion and Center for Modeling of Turbulence and Transition, Lewis Research Center, 1991. ???? pp. Shipping list no.: 92-0454-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1991] 1 microfiche.

Landau:1997:CP

[LM97] Rubin H. Landau and Manuel José Páez Mejía. Computational physics. A Wiley-Interscience Publication. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1997. ISBN 0-471-11590-8. xxx + 520 pp. Problem solving with computers, With contributions by Hans Kowallik and Henri Jansen, With 1 IBM-PC floppy disk (3.5 inch; DD).

Lewis:1991:NMT

[LMH91] R. W. Lewis, K. Morgan, and G. M. Homsy, editors. Numerical methods in thermal problems: proceedings of the seventh international conference held in Stanford, U.S.A. on July 8–12th, 1991, volume VII of Proceedings of the Seventh International Conference on Numerical Methods in Thermal Problems. Pineridge Press, Swansea, UK, 1991. ISBN 0-906674-76-X. LCCN QC319.8 .N84 1991 Bar. Two volumes.

Luxmoore:1990:NMF

[LO90] A. R. Luxmoore and D. R. J. Owen, editors. Numerical methods in fracture mechanics: proceedings of the Fifth International Conference held in Freiburg, Federal Republic of Germany, 23rd–27th April, 1990. Pineridge Press, Swansea, UK, 1990. ISBN 0-906674-73-5. LCCN TA409.I448 1990. UK£94.00. Spine title: Proceedings of the 5th International Conference on Numeral Methods in Fracture Mechanics.

Lorentz:1992:MBI

[Lor92] Rudolph A. Lorentz. Multivariate Birkhoff Interpolation, volume 1516 of Lecture Notes in Mathematics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNMAA2. ISBN 3-540-55870-5 (print), 3-540-47300-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 192 pp. LCCN QA3 .L28

no. 1516. URL http://link.
springer.com/book/10.1007/
BFb0088788;   http://www.
springerlink.com/content/
978-3-540-47300-8.

Lozier:1997:PST

[Loz97] D. W. Lozier. A proposed software test service for special functions. In Boisvert [Boi97], pages 167–178. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Laigna:1990:MMP

[LP90] K. J. Laigna and E. L. Potashnik. Mathematical model for predicting the atmospheric pollution state in the areas of local and regional scale. Estonian Scientific Research Institute of Scientific-Technical Information and Technical-Economic Research, Tallinn, U.S.S.R., 1990. 342 pp. Contents: Ch. 1. The present state of research concerning the problem of how to protect the surrounding environment from pollution for atmospheric and water naturalhousehold objects on local and regional scale. — Ch. 2. Mathematical formulation of a spatial numerical model about the regional pollution transference. — Ch. 3. Methods for the model’s numerical realization.– Ch. 4. Prognosis model as to the atmospheric pollution state in Estonia and its north-eastern region — Ch. 5. Appendices.

Landau:1997:CPP

[LP97] Rubin H. Landau and Manuel José Páez Mejía. Computational physics: problem solving with computers. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1997. ISBN 0-471-11590-8 (hardcover). xxviii + 520 pp. LCCN QC20.82 .L36 1997. URL http:/

                                        /www.loc.gov/catdir/bios/
                                        wiley044/96051776.html;
                                        http://www.loc.gov/catdir/
                                        description/wiley037/96051776.
                                        html;   http://www.loc.gov/
                                        catdir/toc/wiley022/96051776.
                                        html.

Lucquin:1998:ISC

[LP98] Brigitte Lucquin and Olivier Pironneau. Introduction to Scientific Computing. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, January 1998. ISBN 0-471-97266-5 (paper), 0-471-97934-1 (PPC). xviii + 361 pp. LCCN QC20.7.D5 L8313 1998. US$64.95. URL

                                        http://www.loc.gov/catdir/
                                        description/wiley032/97022466.
                                        html;   http://www.loc.gov/
                                        catdir/toc/wiley022/97022466.

html. Translated from the French by Michel Kern.

Lapidus:1999:NSP

[LP99] Leon Lapidus and George F. Pinder. Numerical solution of partial differential equations in science and engineering. A Wiley-Interscience Publication.

John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1999. ISBN 0-471-35944-0. xiv + 677 pp.

Laprise:1997:NMA

[LRRL97] René Laprise, André J. Robert, Harold Ritchie, and Charles Augustin Lin, editors. Numerical methods in atmospheric and oceanic modelling: the André J. Robert memorial volume. Canadian Meteorological and Oceanographic Society = Société canadienne de météorologie et d’océanographie, Ottawa, ON, Canada, 1997. ISBN 0-9698414-4-2. LCCN QC996.N78 1997. A collection of papers based on presentations at the André J. Robert Memorial Symposium on Numerical Methods in Atmospheric and Oceanic Sciences, held at Université du Québec à Montréal Oct. 5–7, 1994. Includes foreword, summaries and supplement in French. “Primary publications of André J. Robert”.

Lee:1997:SFL

[LS97] S. L. Lee and W. E. Schiesser.

Some fundamental limitations of mathematical soilware revealed by the calculation of spacetime curvature. In Boisvert [Boi97], pages 345–348. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Lum:1990:GNM

[LSR90] W. E. Lum, James L. Smoot, and Dale R. Ralston. Geohydrology and numerical model analysis of ground-water flow in the Pullman-Moscow area, Washington and Idaho. Books and Open-File Reports Section [distributor], Denver, Colo., 1990. vi + 73 pp. Shipping list no.: 90-189-P.

Lascaux:1994:ANM

[LT94] P. Lascaux and R. Théodor. Analyse numérique matricielle appliquée à l’art de l’ingénieur. Tome 2. (French) [Matrix numerical analysis applied to the art of the engineer. Volume 2]. Masson Editeur, Masson, France, second edition, 1994. ISBN 2-225-84546-8. i–xxvi and 327–636 pp. Méthodes itératives. [Iterative methods].

Lozier:1996:EBL

[LT96] D. W. Lozier and P. R. Turner. Error-bounding in level-index computer arithmetic. In Alefeld and Herzberger [AH96], pages 138–145. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Li:1996:RPH

[LWW96] T. Y. Li, Tianjun Wang, and Xioshen Wang. Random product homotopy with minimal BKK bound. In Renegar et al. [RSS96], pages 503–512. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Lyness:1997:EMS

[Lyn97] J. N. Lyness. Evaluation of minimization software based on performance profile techniques.

In Boisvert [Boi97], pages 152– 154. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

McCallen:1997:SNS

[MAA97] David McCallen and Abolhassan Astaneh-Asl. SUSPN-DRS: a numerical simulation tool for the nonlinear transient analysis of cable supported bridge structures. part I, theoretical development. Technical Report UCRL-ID-127626, Lawrence Livermore National Laboratory, Livermore, CA, USA, June 1997. 138 pp.

Manno:1999:IMC

[Man99] István Manno. Introduction to the Monte-Carlo method. Akadémiai Kiadó, Budapest, Hungary, 1999. ISBN 963-05-7615-5. 162 pp. LCCN MLCM 99/02850 (Q).

Marti:1992:SON

[Mar92] Kurt Marti, editor. Stochastic optimization: numerical methods and technical applications: proceedings of a GAMM/IFIPworkshop held at the Federal Armed Forces University Munich, Neubiberg, FRG, May 29– 31, 1990, volume 379 of Lecture notes in economics and mathematical systems. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-540-55225-1 (Berlin), 0-387-55225-1 (New York). LCCN QA402.5 .S73 1992; HD30 .L432 no.379. Proceedings of a GAMM/IFIPworkshop held at the Federal Armed Forces University Munich, Neubiberg, FRG, May 29– 31, 1990.

Marchuk:1994:NMA

[Mar94] G. I. Marchuk. Numerical methods and applications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1994. ISBN 0-8493-8947-X. x + 272 pp. LCCN QA297 .N8619 1994.

Mathew:1992:SAI

[Mat92a] Tarek P. Mathew. Schwarz alternating and iterative refinement methods for mixed formulations of elliptic problems. Part I, Algorithms and numerical results. Dept. of Mathematics, University of California, Los Angeles, Los Angeles, CA, USA, January 1992. 22 pp.

Mathews:1992:NMM

[Mat92b] John H. Mathews. Numerical methods for mathematics, science, and engineering. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1992. ISBN 0-13-624990-6. x + 646 pp. LCCN QA297 .M39 1992. Rev. ed. of: Numerical methods for computer science, engineering, and mathematics. 1987.

Mathe:1996:ORS

[Mat96] Peter Mathé. Optimal reconstruction of stochastic evolutions. In Renegar et al. [RSS96], pages 523–532. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Morton:1993:NMF

[MB93] K. W. Morton and M. J. Baines, editors. Numerical methods for fluid dynamics 4, Oxford science publications. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1993. ISBN 0-19-853696-8. LCCN TA357 .N87261 1993. Proceedings of a conference held at Reading University in April 1992. Spine title: Numerical methods for fluid dynamics.

Morton:1995:NMF

[MB95] K. W. Morton and M. J.

Baines, editors. Numerical methods for fluid dynamics V. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1995. ISBN 0-19-851480-8 (hardcover). LCCN TA357 .N8728 1995. Proceedings of a conference held at Oxford University, April 1995.

McVay:1990:DDP

[McV90] M. C. McVay. Development of design procedures for estimating capacity and deformation of pile groups Volume 2, Numerical analysis of vertically loaded pile groups. Dept. of Civil Engineering, University of Florida, Gainesville, FL, USA, 1990. vi + 154 pp. Final report. Microfiche. Springfield, Va.: National Technical Information Service, 1990. 2 microfiches: negative ; 11 x 15 cm. Prepared for the Florida Dept. of Transportation, Bureau of Materials and Research in cooperation with the Federal Highway Administration under grant no. HPR-86-29.

Macedo:1990:KSD

[MDE90] C. G. Macedo, J. C. Diaz, and R. E. Ewing. A knowledgebased system for the determination of activity indicators for self-adaptive grid methods. In Houstis et al. [HRV90], pages 133–142. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Meinguet:1996:NAN

[Mei96] Jean Meinguet, editor. Numerical analysis: a numerical analysis conference in honour of Jean Meinguet. Belgian Mathematical Society, Brussels, Belgium, 1996. LCCN QA297.N826 1996. This volume is published as a supplement to the Bulletin of the Belgian Mathematical Society, Simon Stevin … December 1996. Contents: Matrix computation and the theory of moments / Gene H. Golub (pp. 1–9), — Convergence of rational interpolants / Herbert Stahl — Variations on Richardson’s method and acceleration / Claude Brezinski — Andre Louis Cholesky / Claude Brezinski — On the convergence of multivariate Padé approximants / Annie Cuyt, Doron Lubinsky — Leastsquares fitting of circles and ellipses / Walter Gander, Gene H. Golub, Rolf Strebel (pp. 63– 84) — Entropy of Hermite polynomials with application to the harmonic oscillator / Walter van Assche. 6 papers in English, 1 in French.

Meinert:1997:NPS

[Mei97] Karen M. Meinert. Numerical propulsion system simulation, software standards and procedures, specification. National Technical Information Service, Washington, DC, USA, 1997. ???? pp. Shipping list no.: 98-0733-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1997] 1 microfiche.

Meister:1999:NLG

[Mei99] Andreas Meister. Numerik linearer Gleichungssysteme. (German) [Numerical solution of linear equation systems]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1999. ISBN 3-528-03135-2. x + 222 pp. Eine Einführung in moderne Verfahren. [An introduction to modern procedures].

Merritt:1995:CTF

[Mer95] Michael L. Merritt. Computation of the time-varying flow rate from an artesian well in central Dade County, Florida, by analytical and numerical simulation methods. Earth Science Information Center, Open-File Reports Section [distributor], Reston, VA, USA, 1995. ???? pp. Shipping list no.: 97-0273-M. Microfiche. [Denver, Colo.: U.S. Geological Survey, 1997] 1 microfiche.

Merritt:1997:CTF

[Mer97] Michael L. Merritt. Computation of the time-varying flow rate from an artesian well in central Dade County, Florida, by analytical and numerical simulation methods. U.S. Geological Survey, Branch of Information Services, Denver, CO, USA, 1997. iv + 44 pp. Prepared in cooperation with the Metropolitan-Dade Department of Environmental Resources Management. U.S. Department of the Interior, U.S. Geological Survey. Shipping list no.: 98-0001-P.

Metropolis:1990:AE

[Met90] N. Metropolis. The Los Alamos experience, 1943–1954. In Nash [Nas90b], pages 237–250. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Metzner:1996:COO

[Met96] Lars Metzner. On the calculation of the Q-order and R-order for coupled sequences. In Alefeld and Herzberger [AH96], pages 146–151. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Mathews:1999:NMU

[MF99] John H. Mathews and Kurtis D. Fink. Numerical methods using MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 1999. ISBN 0-13-270042-5. viii + 662 pp. LCCN QA297 .M39 1999. Rev. ed. of:

Numerical methods for mathematics, science, and engineering. 2nd ed. 1992.

Mitchell:1996:NAA

[MGW96] A. R. Mitchell, D. F. Griffiths, and G. A. Watson. Numerical analysis: A.R. Mitchell 75th birthday volume. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1996. ISBN 981-02-2719-1. ix + 370 pp. LCCN QA297 .N827 1996.

Mikhlin:1991:EAN

[Mik91] S. G. (Solomon Grigor’evich)

Mikhlin. Error analysis in numerical processes. Pure and applied mathematics. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1991. ISBN 0-471-92133-5. 283 pp. LCCN QA275 .M6413 1991. US$70.00. Rev. translation of: Fehler in numerischen Prozessen. A Wiley-Interscience publication.

Marti:1995:SPN

[MK95] Kurt Marti and Peter Kall, editors. Stochastic programming: numerical techniques and engineering applications: proceedings of the 2nd GAMM/IFIP-Workshop on “Stochastic Optimization: Numerical Methods and Technical Applications”, held at the Federal Armed Forces University Munich, Neubiberg/ München, Germany, June 15– 17, 1993, volume 423 of Lecture notes in economics and mathematical systems. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-58996-1 (New York), 3-540-58996-1 (Berlin). LCCN T57.79.G36 1993.

Marti:1998:SPM

[MK98] Kurt Marti and Peter Kall, editors. Stochastic programming methods and technical applications: proceedings of the 3rd GAMM/IFIP-Workshop on “Stochastic Optimization: Numerical Methods and Technical Applications”, held at the Federal Armed Forces University Munich, Neubiberg/München, Germany, June 17–20, 1996, volume 458 of Lecture notes in economics and mathematical systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-63924-1. ISSN 0075-8442. LCCN T57.79 .G36 1998.

Maron:1990:NAP

[ML90] Melvin J. Maron and Robert J. Lopez. Numerical analysis: a practical approach. Wadsworth Publishing Company, Pacific Grove, CA, USA, third edition, 1990. ISBN 0-534-12372-4. xxi + 743 pp. LCCN QA297 .M344 1991.

Morton:1994:NSP

[MM94] K. W. Morton and D. F. (David Francis) Mayers. Numerical solution of partial differential equations: an introduction. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-41855-0 (hardcover), 0-521-42922-6 (paperback). x + 227 pp. LCCN QA377 .M69 1994.

URL http://www.loc.gov/
catdir/description/cam026/
94006670.html. An introduc-

tion.

Mrozek:1996:MUS

[MM96] Bogumila Mrozek and Zbigniew Mrozek. MATLAB, Uniwersalne srodowisko do obliczen naukowo-technicznych (Polish) [MATLAB, Universal Environment for Scientific and Engineering Computing]. Wydawnictwo PLJ, Warszawa, Poland, 1996. ISBN 83-7101-325-6. 261 pp. LCCN ????

Magnus:1999:MDC

[MN99] Jan R. Magnus and Heinz Neudecker. Matrix Differential Calculus with Applications in Statistics and Econometrics. John Wiley and Sons, Inc., Chichester, UK, revised edition, 1999. ISBN 0-471-98633-X. xviii + 395 pp.

Morand:1995:FSI

[MO95] H. J.-P Morand and R. Ohayon.

Fluid structure interaction: applied numerical methods. Masson Editeur, Masson, France, 1995. ISBN 0-471-94459-9 (Wiley), 2-225-84682-0 (Masson). viii + 212 pp. LCCN TA357.5.F58 M67 1995. Translation by Claude Andrew James of: Interactions fluidesstructures. Paris: Masson, 1992.

Modica:1998:AMN

[Mod98] Edward Modica. Analytical methods, numerical modeling, and monitoring strategies for evaluating the effects of groundwater withdrawals on unconfined aquifers in the New Jersey Coastal Plain. Branch of Information Services [distributor], Denver, CO, USA, 1998. vii + 65 pp. Prepared in cooperation with the New Jersey Department of Environmental Protection.

Moore:1990:TAN

[MOF90] P. K. Moore, C. Ozturan, and J. E. Flaherty. Towards the automatic numerical solution of partial differential equations. In Houstis et al. [HRV90], pages 15–22. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Mohr:1992:FES

[Moh92] G. A. Mohr. Finite elements for solids, fluids, and optimization. Oxford Science Publications. Clarendon Press, New York, NY, USA, 1992. ISBN 0-19-856369-8 (hardcover), 0-19-856368-X (paperback). xvi + 604 pp. LCCN TA347.F5 M64 1992. ^{1}60.00 (US$90.00 U.S.)^{1}27.50 (US$41.50 U.S.).

                                        URL http://www.loc.gov/
                                        catdir/enhancements/fy0638/
                                        92009625-d.html;   http://
                                        www.loc.gov/catdir/enhancements/
                                        fy0638/92009625-t.html.

Morton:1990:TIC

[Mor90] K. W. Morton. Twelfth International Conference on Numerical Methods in Fluid Dynamics: proceedings of the conference held at the University of Oxford, England on 9–13 July 1990, volume 371 of Lecture notes in physics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 0-387-53619-1 (U.S.). xiv + 562 pp. LCCN QA911 .I58 1990.

Morgan:1992:NM

[Mor92] Don Morgan. Numerical Methods: Real-time and Embedded Systems Programming. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1992. ISBN 1-55851-232-2 (book only), 1-55851-233-0 (disk only), 1-55851-234-9 (book and disk set). 496 pp. LCCN QA76.6 .M669 1992. US$34.95.

Morgan:1997:NMD

[Mor97a] Don Morgan. Numerical methods for DSP systems in C. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1997. ISBN 0-471-13232-2 (paperback). xx + 332 pp. LCCN TK5102.9 .M65 1997 Bar. Wiley Computer Publishing.

Morton:1997:AMH

[Mor97b] K. W. Morton. Approximation of multidimensional hyperbolic partial differential equations. In Duff and Watson [DW97], pages 473–502.

ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

//global.oup.com/academic/
product/the-state-of-the-
art-in-numerical-analysis-
9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Miller:1996:FNM

[MOS96] John J. H. Miller, E. O’Riordan, and G. I. Shishkin. Fitted numerical methods for singular perturbation problems: error estimates in the maximum norm for linear problems in one and two dimensions. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1996. ISBN 981-02-2462-1. xiv + 166 pp. LCCN QA371 .M55 1996.

Mackens:1996:SEI

[MR96] W. Mackens and S. M. Rump. Software Engineering im Scientific Computing. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1996. ISBN 3-528-06904-X. x + 315 pp. LCCN QA76.758 .S6464527 1996.

Mitsui:1995:NAO

[MS95] T. Mitsui and Y. Shinohara, editors. Numerical analysis of ordinary differential equations and its applications. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1995. ISBN 981-02-2229-7. LCCN QA372 .N85 1995. Based on papers from the Kyoto Workshop on Numerical Analysis of ODEs held Nov. 1994 at Kyoto University.

Mu:1997:STE

[Mu97] M. Mu. Software testing and evaluation in large-scale scientific applications. In Boisvert [Boi97], pages 330–332. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Munson:1997:FAS

[Mun97] J. C. Munson. A functional approach to software reliability modeling. In Boisvert [Boi97], pages 61–76. ISBN 0-412-80530-

8. LCCN QA297.I35 1996. US$146.50.

Mohamed:1986:NAB

[MW86] J. L. Mohamed and Joan E.

Walsh, editors. Numerical algorithms: [based on lectures given at a joint Liverpool–Manchester summer school in 1984], Manchester/Liverpool summer schools in numerical analysis. Clarendon Press, New York, NY, USA, 1986. ISBN 0-19-853364-

0. LCCN QA297 .N79 1986.

URL http://catdir.loc.gov/
catdir/enhancements/fy0637/
87001546-d.html;   http://
catdir.loc.gov/catdir/enhancements/
fy0637/87001546-t.html;
http://www.gbv.de/dms/hbz/
toc/ht002933977.pdf.

More:1993:OSG

[MW93] Jorge J. More and Stephen J. Wright, editors. Optimization software guide, volume 14 of Frontiers in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1993. ISBN 0-89871-322-6. LCCN QA402.5 .M67 1993.

Maller:1996:CCA

[MW96] Michael Maller and Jennifer Whitehead. Computational complexity over the 2-adic numbers. In Renegar et al. [RSS96], pages 513–522. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Noda:1990:ASP

[NA90] M.-T. Noda and S. Asagawa. Algorithm selection problem in the hybrid computation. In Houstis et al. [HRV90], pages 155–160. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Nakamura:1991:ANM

[Nak91] Shoichiro Nakamura. Applied numerical methods with software. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1991. ISBN 0-13-041047-0. xiii + 570 pp. LCCN QA297 .N283 1991. US$27.75.

Nakamura:1993:ANM

[Nak93] Shoichiro Nakamura. Applied numerical methods in C. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1993. ISBN 0-13-042052-2. xiv + 604 pp. LCCN QA76.73.C15 N35 1993.

Nakamura:1996:NAG

[Nak96] Shoichiro Nakamura. Numerical analysis and graphic visualization with MATLAB. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-051518-3. xi + 477 pp. LCCN T385 .N34 1996.

Nakagawa:1997:DEG

[Nak97] M. Nakagawa. Development of efficient general purpose Monte Carlo codes used in nuclear engineering. In Boisvert [Boi97], pages 349–360. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Nash:1990:CNM

[Nas90a] John C. Nash. Compact numerical methods for computers: linear algebra and function minimisation. Adam Hilger Ltd., Bristol, UK, second edition, 1990. ISBN 0-85274-318-1 (case), 0-85274-319-X (paperback). xii + 278 pp. LCCN QA184 .N37 1990.

Nash:1990:HSC

[Nas90b] Stephen G. Nash, editor. A History of Scientific Computing. ACM Press history series. Addison-Wesley and ACM Press, Addison-Wesley and New York, NY 10036, USA, 1990. ISBN 0-201-50814-1. xix + 359 pp. LCCN QA76.17 .H59 1990.

Natterer:1997:AT

[Nat97] Frank Natterer. Algorithms in tomography. In Duff and Watson [DW97], pages 503–524. ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

//global.oup.com/academic/
product/the-state-of-the-
art-in-numerical-analysis-
9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Nawrotzki:1994:NSA

[Naw94] Peter Nawrotzki. Numerical stability analysis of arbitrary structural responses, volume 24 of SFB 151 Berichte Nr. SFB 151 Tragwerksdynamik, Ruhr-Universität Bochum, Bochum, Germany, February 1994. 124 pp. LCCN 500 S663 24. Wissenschaftliche Mitteilungen. Projekt B3: Dynamisches Stabilitatsverhalten allgemeiner, insbesondere flachenhafter Strukturen.

Nazareth:1996:LGS

[Naz96] J. L. Nazareth. Lagrangian globalization: Solving nonlinear equations via constrained optimization. In Renegar et al. [RSS96], pages 533–542. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Neittaanmaki:1991:NMF

[Nei91] P. Neittaanmaki, editor. Numerical methods for free boundary problems: proceedings of a conference held at the Department of Mathematics, University of Jyvaskyla, Finland, July 23–27, 1990, volume 99 of International series of numerical mathematics = Internationale Schriftenreihe zur numerischen Mathematik. Birkhäuser Verlag, Basel, Switzerland, 1991. ISBN 3-7643-2641-7 (Basel), 0-8176-2641-7 (Boston). LCCN QA379 .N87 1991. 124.00F (approx.). Papers from the Conference on Numerical Methods for Free Boundary Problems.

Nemirovski:1996:PTM

[Nem96] Arkadi Nemirovski. Polynomial time methods in convex programming. In Renegar et al. [RSS96], pages 543–590. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Nyhoff:1995:FNM

[NL95] Larry R. Nyhoff and Sanford Leestma. Fortran 77 and Numerical Methods for Engineers and Scientists. Prentice-Hall, Upper Saddle River, NJ 07458, USA, April 1995. ISBN 0-02-388741-9 (paperback). xviii + 764 pp. LCCN QA 76.73 F25 N89 1995. US$65.00. URL http:/

/www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=

0023887419. Includes an introduction to FORTRAN 90.

Nath:1996:CHO

[NL96] Y. Nath and B. Lenka. Comparison of higher order Runge Kutta methods for chaotic vibrations of double pendulum. In Alefeld and Herzberger [AH96], pages 152–157. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Nichols:1996:TMS

[NMW96] Len M. Nichols, Marilyn Moon, and Susan Wall. Tax-preferred medical savings accounts and catastrophic health insurance plans: a numerical analysis of winners and losers. Technical Report 06571-002, Urban Institute, Washington, DC, April 1996. 41 pp. Supported by the Robert Wood Johnson Foundation.

Nath:1996:VAS

[NN96] Y. Nath and V. Nayar. Variational approach to snap through buckling analysis of orthotropic curved plates. In Alefeld and Herzberger [AH96], pages 158– 163. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Nocedal:1997:LSU

[Noc97] Jorge Nocedal. Large scale unconstrained optimization. In Duff and Watson [DW97], pages 311–338. ISBN 0-19-850014-9. LCCN QA297.S775 1997.

                                        URL https://global.oup.
                                        com/academic/product/the-
                                        state-of-the-art-in-numerical-
                                        analysis-9780198500148. Based

on the proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Nogami:1997:OMN

[Nog97] Toyoaki Nogami. Observation and modeling in numerical analysis and model tests in dynamic soil-structure interaction problems: proceedings of sessions sponsored by the Geo-Institute of the American Society of Civil Engineers held in conjunction with Geo-Logan ’97, Logan, Utah, July 16–17, 1997, volume 64 of Geotechnical special publication. American Society of Civil Engineers, Dallas, Texas, 1997. ISBN 0-7844-0252-3. vii + 141 pp. LCCN TA711.5 .O27 1997.

Novak:1996:BAN

[Nov96] Erich Novak. The Baysian approach to numerical problems: Results for zero finding. In Alefeld and Herzberger [AH96], pages 164–171. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Napolitano:1993:TIC

[NS93] M. Napolitano and F. Sabetta.

Thirteenth International Conference on Numerical Methods in Fluid Dynamics: proceedings of the conference held at the Consiglio Nazionale delle Ricerche, Rome, Italy, 6–10 July 1992, volume 414 of Lecture notes in physics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-56394-6 (Berlin), 0-387-56394-6 (U.S.). xiv + 541 pp. LCCN QA911 .I58 1992.

Nurminskii:1991:CMV

[Nur91] Evgenii Alekseevich Nurminskii. Chislennye metody vypukloi optimizatsii. Nauka, Moscow, Russia, 1991. ISBN 5-02-000127-9. 165 pp. 2.00rub. Nauchnoe izdanie. Title on t.p. verso: Numerical methods for convex optimization. At head of title: Akademiia nauk SSSR. Dal’nevostochnoe otdelenie. Institut prikladnoi matematiki. Summary and table of contents in English.

Oden:1990:HCF

[Ode90] J. Tinsley Oden. Historical comments on finite elements. In Nash [Nas90b], pages 152–166. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Ortega:1999:ICN

[OG99] James M. Ortega and Andrew Swift Grimshaw. An introduction to C++ and numerical methods. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-511767-0 (paperback). x + 273 pp. LCCN QA297 .O778 1998.

Oki:1996:NAG

[Oki96] Delwyn S. Oki. Numerical analysis of ground-water flow and salinity in the Ewa Area, Oahu, Hawaii. U.S. Geological Survey, Branch of Information Services [distributor], Denver, CO, USA, 1996. ???? pp. Shipping list no.: 97-0816-M.

Olver:1990:REA

[Olv90] F. (Frank) W. J. Olver. Rounding errors in algebraic processes — in level-index arithmetic. In Cox and Hammarling [CH90], pages 197–205. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Opfer:1993:NMA

[Opf93] Gerhard Opfer. Numerische Mathematik für Anfänger. (German) [Numerical Methods for Beginners], volume 65 of Vieweg Studium: Grundkurs Mathematik [Vieweg Studies: Basic Mathematics Course]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1993. ISBN 3-528-07265-2. xiv + 288 pp.

Ortega:1990:NA

[Ort90a] James M. Ortega. Numerical analysis, volume 3 of Classics in Applied Mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, second edition, 1990. ISBN 0-89871-250-5. xiv + 201 pp. A second course.

Ortega:1990:NAS

[Ort90b] James M. Ortega. Numerical analysis: a second course, volume 3 of Classics in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1990. ISBN 0-89871-250-

5. xiii + 201 pp. LCCN QA297 .O78 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0708/89048562-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0708/
89048562-t.html.

Ortega:1994:IFSb

[Ort94a] James M. Ortega. An Introduction to Fortran 90 for Scientific Computing. Saunders College Pub., Fort Worth, TX, USA, August 1994. ISBN 0-03-010198-

0. vi + 234 pp. LCCN QA 76.73 F25 O75 1994b. US$30.00; US$20.00. URL http://

www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0030101980.

Ortega:1994:IFSa

[Ort94b] James M. Ortega. An Introduction to Fortran for Scientific Computing. Saunders College Pub., Fort Worth, TX, USA, January 1994. ISBN 0-03-003128-1. viii + 204 pp. LCCN QA76.73.F25 O75 1994. US$30.00. URL http:/

/www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0030031281.

Osher:1994:SCN

[Osh94] Stanley Osher. Subscale capturing in numerical analysis. Dept. of Mathematics, University of California, Los Angeles, Los Angeles, CA, USA, July 1994. 15 pp.

Osterweil:1997:IQS

[Ost97] L. J. Osterweil. Improving the quality of software quality determination processes. In Boisvert [Boi97], pages 90–108. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Paige:1990:SAG

[Pai90] Chris Paige. Some aspects of generalized QR factorizations. In Cox and Hammarling [CH90], pages 73–91. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Palm:1998:IME

[Pal98] William J. (William John)

Palm. Introduction to MAT-LAB for engineers. McGraw-Hill’s BEST–basic engineering series and tools. WCB/McGraw-Hill, Boston, MA, USA, 1998. ISBN 0-07-047328-5. xx + 316 pp. LCCN QA297.P32 1998.

Pan:1996:EPC

[Pan96] Victor Y. Pan. Effective parallel computations with Toeplitz and Toeplitz-like matrices filled with integers. In Renegar et al. [RSS96], pages 591–642. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Pancake:1997:IQT

[Pan97a] C. M. Pancake. Improving quality through user-centered design. In Boisvert [Boi97], pages 44–60. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Pang:1997:ICP

[Pan97b] Tao Pang. An introduction to computational physics. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-48143-0 (hardcover), 0-521-48592-4 (paperback). xvii + 374 pp. LCCN QC20.7.E4 P36 1997.

                                        URL http://www.loc.gov/
                                        catdir/description/cam028/
                                        96045574.html;   http://www.
                                        loc.gov/catdir/toc/cam023/
                                        96045574.html.

Parlett:1990:CJH

[Par90a] B. N. Parlett. The contribution of J. H. Wilkinson to numerical analysis. In Nash [Nas90b], pages 17–30. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Parlett:1990:MLA

[Par90b] Beresford N. Parlett. Misconvergence in the Lanczos algorithm. In Cox and Hammarling [CH90], pages 7–24. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Park:1991:NAL

[Par91] Hee Kuwon Park. Numerical analysis of laser melting of thin silicon films: research project. Master of science, plan ii, Dept. of Mechanical Engineering, University of California, Berkeley, Berkeley, CA, USA, 1991. various pp.

Paxson:1992:GNM

[Pax92] Daniel E. Paxson. A general numerical model for wave rotor analysis. National Technical Information Service, Washington, DC, USA, 1992. ???? pp. Shipping list no.: 93-0047-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration], 1992. 1 microfiche.

Paxson:1995:NMD

[Pax95] Daniel E. Paxson. A numerical model for dynamic wave rotor analysis. National Technical Information Service, Washington, DC, USA, 1995. ???? pp. Shipping list no.: 96-0021-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1995] 1 microfiche.

Pandey:1990:NMR

[PBW90] G. N. Pandey, G. Beer, and J. R. Williams. Numerical methods in rock mechanics. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1990. ISBN 0-471-92021-5. ix + 327 + [4] pp. LCCN TA706 .P34 1990. US$64.40.

Paris:1997:BEM

[PC97] Federico París and José Cañas. Boundary element method: fundamentals and applications. Oxford Science Publications. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1997. ISBN 0-19-856537-2, 0-19-856543-7 (paperback). xv + 392 pp. LCCN TA347.B69 P34 1997. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0637/96034311-
                                        d.html;   http://www.loc.
                                        gov/catdir/enhancements/fy0637/
                                        96034311-t.html. Fundamen-

tals and applications, With 1 Macintosh floppy disk (3.5 inch; HD).

Peaceman:1990:PRR

[Pea90] Donald W. Peaceman. A personal retrospection of reservoir simulation. In Nash [Nas90b], pages 106–129. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Pember:1992:NMH

[Pem92] Richard Bissell Pember. Numerical methods for hyperbolic conservation laws with stiff relaxation. Thesis (ph.d.) in applied mathematics, University of California, Berkeley, Berkeley, CA, USA, May 1992. viii + 109 pp.

Part-Enander:1999:MH

[PES99] Eva Pärt-Enander and Anders Sjöberg. The MATLAB 5 handbook. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-39845-1. xv + 559 pp. LCCN QA297.P47 1999. Rev. ed. of: The MATLAB handbook (1996).

Petkovic:1996:AMS

[Pet96] Miodrag S. Petković. Asynchronous methods for simultaneous inclusion of polynomial roots. In Alefeld and Herzberger [AH96], pages 178–187. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Press:1992:NRC

[PFTV92] W. H. Press, B. P. Flannery, S. A. Teukolsky, and W. T. Vetterling. Numerical Recipes in C: The Art of Scientific Computing. Cambridge University Press, Cambridge, UK, second edition, 1992. ISBN 0-521-43108-5, 0-521-43720-2. xxvi + 994 pp. LCCN QA297 .N864 1992.

Pepper:1992:FEM

[PH92] Darrell W. (Darrell W.) Pepper and Juan C. Heinrich. The finite element method: basic concepts and applications. Series in Computational and Physical Processes in Mechanics and Thermal Sciences. Hemisphere Publishing Corp., Washington, DC, USA, 1992. ISBN 1-56032-104-0. xii + 240 pp. LCCN TA347.F5 P46 1992. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0653/91037924-

d.html. Basic concepts and applications, With 1 IBM-PC floppy disk (5.25 inch; HD).

Pielczyk:1991:NMS

[Pie91]  Andreas Pielczyk.  Numeri-

cal methods for solving systems of quasidifferentiable equations, volume 124 of Mathematical systems in economics. A. Hain, Frankfurt am Main, Germany, 1991. ISBN 3-445-09827-1. 70 pp. LCCN QA372 .P645 1991 Bar.

Penny:1995:NMU

[PL95] J. E. T. (John E. T.) Penny and G. R. Lindfield. Numerical methods using MATLAB. Ellis Horwood, New York, NY, USA, 1995. ISBN 0-13-030966-4 (paperback). xii + 328 pp. LCCN QA297 .P45 1995. US$40.00. Errata slip inserted.

Plaskota:1996:SCC

[Pla96] Leszek Plaskota. Survey of computational complexity with noisy information. In Renegar et al. [RSS96], pages 651–664. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Pelikan:1994:ANM

[PLC94] Peter Pelikan, Marek Liska, and Michal Ceppan. Applications of numerical methods in molecular spectroscopy. Chemometrics series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1994. ISBN 0-8493-7322-0. 341 pp. LCCN QD96.M65 P45 1994.

Plybon:1992:IAN

[Ply92] Benjamin F. Plybon. An introduction to applied numerical analysis. Prindle, Weber and Schmidt series in mathematics. PWS-Kent Pub. Co., Boston, MA, USA, 1992. ISBN 0-534-92284-8, 0-534-97205-5 (international student ed.). xi + 644 pp. LCCN QA297.P58 1992.

Pandey:1990:NME

[PM90] G. N. Pandey and J. Middleton, editors. Numerical methods in engineering: theory and appreciations. Elsevier, Amsterdam, The Netherlands, 1990. ISBN 1-85166-462-9 (set), 1-85166-463-7 (vol. 1), 1-85166-464-5 (vol. 2). LCCN TA329 .I54 1990 v. 1–2 (c1990). Two volumes. NUMETA 90. Proceedings of the Third International Conference on Advances in Numerical Methods in Engineering: Theory and Applications, (NU-META 90), 7–11 January, 1990, University College of Swansea, Wales, UK. Contents: vol. 1. Numerical techniques of engineering analysis and design — vol. 2. Transient/dynamic analysis and constitutive laws for engineering materials.

Pool:1997:NSR

[Poo97] J. C. T. Pool. Is numerical software relevant? Is it too late to worry about quality? In Boisvert [Boi97], pages 3–

11. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Powell:1997:RMM

[Pow97] M. J. D. Powell. A review of methods for multivariable interpolation at scattered data points. In Duff and Watson [DW97], pages 283–310. ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

                                        //global.oup.com/academic/
                                        product/the-state-of-the-
                                        art-in-numerical-analysis-
                                        9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Pozrikidis:1998:NCS

[Poz98] C. (Constantine) Pozrikidis. Numerical computation in science and engineering. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-511253-9 (hardcover). xi + 627 pp. LCCN QA297 .P69 1998. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0604/97038185-
                                        d.html;   http://www.loc.
                                        gov/catdir/enhancements/fy0604/
                                        97038185-t.html.

Press:1993:NRFb

[Pre93] William H. Press. Numerical recipes in FORTRAN. Cambridge University Press, Cambridge, UK, 2nd, v2.0 edition, 1993. ISBN 0-521-43716-4. 1 computer disk pp. LCCN QA297.N866 1992.

Press:1994:NRF

[Pre94] William H. Press. Numerical recipes in FORTRAN: the art of scientific computing. Cambridge University Press, Cambridge, UK, 2d, corrected to software version 2.04 edition, 1994. ISBN 0-521-43064-X (book), 0-521-43719-9 (FORTRAN diskette IBM 3.5, 720K). xxvi + 963 pp. LCCN QA76.73.C15 N865 1992.

Prusakov:1993:MMM

[Pru93] G. M. (Grigorii Moiseevich)

Prusakov. Matematicheskie modeli i metody v raschetakh na çEVM. (Russian) [Mathematical models and methods in computer calculations]. Fizmatlit “Nauka”, Moscow, Russia, 1993. ISBN 5-02-015138-6. 141 pp. LCCN QA76.9.M35 P78 1993.

Prossdorf:1990:SAO

[PS90] Siegfried Prossdorf and Bernd Silbermann. Seminar analysis: operator equations and numerical analysis, 1989/90. Karl-Weierstrass-Institut für Mathematik, Berlin, 1990. 102 pp. LCCN QA329 .S47 1990. English and German.

Prossdorf:1991:NAIa

[PS91a] Siegfried Prossdorf and Bernd Silbermann. Numerical analysis for integral and related operator equations, volume 84 of Mathematische Lehrbucher und Monographien. II. Abteilung, Mathematische Monographien. Akademie-Verlag, Berlin, Germany, 1991. ISBN 3-05-500696-8. 542 pp. LCCN QA3 .M328 v.84.

Prossdorf:1991:NAIb

[PS91b] Siegfried Prossdorf and Bernd Silbermann. Numerical analysis for integral and related operator equations, volume 52 of Operator theory, advances and applications. Birkhäuser Verlag, Basel, Switzerland, 1991. ISBN 0-8176-2620-4 (Boston), 3-7643-2620-4 (Basel). 542 pp. LCCN QA431 .P77 1991. SFr. 198.00 (approx.).

Pereverzev:1996:EDS

[PS96a] S. V. Pereverzev and S. G. Solodsky. An efficient discretization for solving ill-posed problems. In Renegar et al. [RSS96], pages 643–650. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Potra:1996:NBR

[PS96b] F. A. Potra and Y. Shi. A note on Brent’s rootfinding method. In Alefeld and Herzberger [AH96], pages 188–197. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455

Phillips:1996:TAN

[PT96] G. M. (George McArtney)

Phillips and Peter John Taylor. Theory and applications of numerical analysis. Academic Press, New York, NY, USA, second edition, 1996. ISBN 0-12-553560-0. xii + 447 pp. LCCN QA297.P52 1996.

Press:1996:NRFa

[PTM96] William H. Press, Saul A.

Teukolsky, and Michael Metcalf. Numerical Recipes in Fortran 90: The Art of Parallel Scientific Computing, volume 2 of Fortran Numerical Recipes. Cambridge University Press, Cambridge, UK, second edition, September 1, 1996. ISBN 0-521-57439-0 (hardcover). 200 pp. LCCN QA76.73.F25 N85 1996. US$44.95; US$40.45.

URL http://www.cbooks.com/
sqlnut/SP/search/gtsumt?source=
&isbn=0521574390;   http:/
/www.nr.com/nronline_switcher.
html.

Press:1992:NRF

[PTV^{+}92] William H. Press, Saul A.

Teukolsky, William T. Vetterling, Brian P. Flannery, Michael Metcalf, et al. Numerical Recipes in FORTRAN: The Art of Scientific Computing. Cambridge University Press, Cambridge, UK, second edition, October 1, 1992. ISBN 0-521-43064-X (book), 0-521-43721-0 (example book) 0-521-43717-2 (diskette), 0-521-43719-9 (diskette), 0-521-43716-4 (diskette). xxvi + 963 pp. LCCN QA76.73.C15 N865 1992. US$39.95; US$54.95.

                                        URL http://www.cbooks.com/
                                        sqlnut/SP/search/gtsumt?source=
                                        &isbn=052143064X;   http:/
                                        /www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0521437164;   http://www.
                                        cbooks.com/sqlnut/SP/search/
                                        gtsumt?source=&isbn=0521437172;
                                         http://www.cbooks.com/sqlnut/
                                        SP/search/gtsumt?source=&
                                        isbn=0521437199.  Includes

MacIntosh disk.

Press:1996:NRFb

[PTV96] William H. Press, Saul A. Teukolsky, and William T. Vetterling. Numerical Recipes in Fortran 77 and Fortran 90: The Art of Scientific and Parallel Computing. Cambridge University Press, Cambridge, UK, September 1996. ISBN 0-521-57440-4. xxxi + 933, xx + 935– 1486 pp. US$39.95. URL http:/

                                        /www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0521574404;   http://www.
                                        nr.com/; http://www.nr.com/
                                        nronline_switcher.html.

Petkovic:1996:SSM

[PTŽ96] Lj. Petković, S. Tričković, and D. Živković. Secant slope method for inclusion of complex zeros of polynomials. In Alefeld and Herzberger [AH96], pages 172–177. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Paxson:1992:INM

[PW92] Daniel E. Paxson and Jack Wilson. An improved numerical model for wave rotor design and analysis. National Technical Information Service, Washington, DC, USA, 1992. ???? pp. Shipping list no.: 93-0304-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration], 1992. 1 microfiche.

Peskin:1990:SNG

[PWF90] R. L. Peskin, S. S. Walther, and A. M. Froncioni. Smalltalk — the next generation scientific computing interface? In Houstis et al. [HRV90], pages 257–268. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Pytlak:1999:NMO

[Pyt99]  Radosław Pytlak.  Numeri-

cal Methods for Optimal Control Problems with State Constraints, volume 1707 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CO-DEN LNMAA2. ISBN 3-540-66214-6 (print), 3-540-48662-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xv + 215 pp. LCCN QA3 .L35

v.1707. URL http://link.
springer.com/book/10.1007/
BFb0097244;   http://www.
springerlink.com/content/
978-3-540-48662-6.

Quarteroni:1998:MN

[QSS98] A. Quarteroni, R. Sacco, and F. Saleri. Matematica numerica. (Italian) [Numerical mathematics]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 88-470-0010-6. xiv + 488 pp.

Quarteroni:1994:NAP

[QV94] Alfio Quarteroni and Alberto (Alberto) Valli. Numerical approximation of partial differential equations, volume 23 of Springer Series in Computational Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-57111-6 (Berlin), 0-387-57111-6 (New York). xvi + 543 pp. LCCN QA377 .Q37 1994.

Quarteroni:1997:NAP

[QV97] Alfio Quarteroni and A. (Alberto) Valli. Numerical approximation of partial differential equations, volume 23 of Springer series in computational mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2nd, corr. print. edition, 1997. ISBN 3-540-57111-6. ISSN 0179-3632. xvi + 543 pp. LCCN QA377 .Q37 1997. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0812/97160884-
                                        d.html.

Rakowski:1994:GSM

[Rak94] Zikmund Rakowski. Geomechanics 93: Strata mechanics, numerical methods, water jet cutting, mechanical rock disintegration: proceedings of the international conference Geomechanics 93, Hradec, Ostrava, Czech Republic, 28–30 September 1993. A. A. Balkema, Brookfield, VT, USA, 1994. ISBN 90-5410-354-X. xiv + 460 pp. LCCN TA703.5 G44 1994. Sponsored by the International Bureau of Strata Mechanics of World Mining Congress and the International Society of Water Jet Technology.

Ramachandran:1994:BEM

[Ram94] P. A. Ramachandran. Boundary element methods in transport phenomena. International Series on Computational Engineering. Computational Mechanics Publications and Elsevier Applied Science, Southampton, UK and London, UK, 1994. ISBN 1-85312-260-2 (Southampton), 1-56252-184-5 (Boston), 1-85861-026-5 (Elsevier Applied Science). xviii + 406 pp. LCCN TA347.B69 R36 1994. With 1 IBM-PC floppy disk (5.25 inch).

Rassias:1994:FAA

[Ras94] John Michael Rassias. Functional analysis, approximation theory, and numerical analysis. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1994. ISBN 981-02-0737-9. xi + 325 pp. LCCN QA321 .F854 1994. US$68.00. Dedicated to Stefan Banach on his 100th birthday, Alexander Markowic Ostrowski on his 99th birthday, Stanislaw Marcin Ulam on his 83rd birthday.

Redfern:1996:MDL

[RC96] Darren Redfern and Ed Chandler. Maple O.D.E. lab book. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94733-7 (softcover). ix + 160 pp. LCCN QA371.5.D37R43 1996.

Redfern:1998:MH

[RC98] Darren Redfern and Colin Campbell. The MATLAB 5 handbook. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-94200-9 (softcover). xi + 488 pp. LCCN QA297 .R395 1998.

Richardson:1993:MNA

[RD93] Lewis Fry Richardson and P. G. Drazin. Meteorology and numerical analysis, volume 1 of Collected papers of Lewis Fry Richardson. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-38297-1. xiv + 1016 pp. LCCN QC861.2 .R4725 1993 Floor 3.

Reed:1991:SDN

[Ree91] Brenda Lee Reed. A short discussion of numerical methods to solve linear and nonlinear elliptic partial differential equations. Thesis (m.a.), University of California, Santa Cruz, Santa Cruz, CA, USA, 1991. v + 131 pp.

Reinsch:1990:SSP

[Rei90] Christian H. Reinsch. Software for shape preserving spline interpolation. In Cox and Hammarling [CH90], pages 267–276. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Reithmeier:1991:PSN

[Rei91] Eduard Reithmeier. Periodic Solutions of Nonlinear Dynamical Systems: Numerical Computation, Stability, Bifurcation and Transition to Chaos, volume 1483 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CODEN LNMAA2. ISBN 3-540-54512-3 (print), 3-540-38427-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 171 pp. LCCN QA3 .L28

no. 1483. URL http://link.
springer.com/book/10.1007/
BFb0094521;   http://www.
springerlink.com/content/
978-3-540-38427-4.

Reider:1992:DHO

[Rei92] Marc B. Reider. Development of higher order numerical methods for two dimensional incompressible flows with applications to flows around circular cylinders and airfoils. Thesis (ph.d.), University of California, Los Angeles, Los Angeles, CA, USA, 1992. x + 176 pp.

Reid:1997:TAE

[Rei97] J. K. Reid. Two approaches to exception handling in Fortran. In Boisvert [Boi97], pages 210– 223. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

ORuanaidh:1996:NBM

[RF96] Joseph J. K. O. Ruanaidh and William J. Fitzgerald. Numerical Bayesian methods applied to signal processing. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94629-2 (New York). xiv + 244 pp. LCCN TK5102.9 .O78 1996.

Rutishauser:1990:LNM

[RG90] Heinz Rutishauser and Martin Gutknecht, editors. Lectures on numerical mathematics. Birkhäuser Boston Inc., Cambridge, MA, USA, 1990. ISBN 0-8176-3491-6. xv + 546 pp. LCCN QA297 .R8713 1990. Edited by Martin Gutknecht with the assistance of Peter Henrici, Peter Läuchli and Hans-Rudolf Schwarz, With a foreword by Gutknecht and a preface by Henrici, Läuchli and Schwarz, Translated from the German and with a preface by Walter Gautschi.

Rheinboldt:1998:MSS

[Rhe98] Werner C. Rheinboldt. Methods for solving systems of nonlinear equations, volume 70 of CBMS-NSF Regional Conference Series in Applied Mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, second edition, 1998. ISBN 0-89871-415-X (paperback). ix + 148 pp. LCCN QA214 .R44 1998. URL

http://www.loc.gov/catdir/
enhancements/fy0726/98027321-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0726/
98027321-t.html.

Richter:1991:NMN

[RHP91] Claus Richter, Horst Hollatz, and Diethard Pallaschke, editors. Numerical methods of nonlinear programming and their implementations, volume 60 of Mathematical research. Akademie-Verlag, Berlin, Germany, 1991. ISBN 3-05-500883-9. ISSN 0138-3019. LCCN QA402.5 .N89 1991. The papers published in this volume are based on lectures given at the conference Opti-Soft ’89, Quedlinburg, July 7–14, 1989.

Rice:1990:MSA

[Ric90] John R. Rice. Mathematical software and ACM publications. In Nash [Nas90b], pages 217–

227. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Rice:1992:NMS

[Ric92] John R. Rice. Numerical Methods, Software, and Analysis. Academic Press, New York, NY, USA, second edition, 1992. ISBN 0-12-587755-2. xiv + 720 pp. LCCN QA297 .R49 1993.

                                        URL http://www.loc.gov/
                                        catdir/description/els032/
                                        93114990.html;   http://www.
                                        loc.gov/catdir/toc/els032/
                                        93114990.html.

Rice:1993:NMS

[Ric93] John Rischard Rice. Numerical methods, software, and analysis. Academic Press, New York, NY, USA, second edition, 1993. ISBN 0-12-587755-2. xiv + 720 pp. LCCN QA297 .R49 1993.

Richardson:1996:LAE

[Ric96] Daniel Richardson. Lazy analysis and elementary numbers. In Renegar et al. [RSS96], pages 665–676. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Rieder:1996:PTC

[Rie96] G. Rieder. Pythagoras, Thales and a classical finite element — some remarks on hypercircle bounding. In Alefeld and Herzberger [AH96], pages 205–296. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Rightley:1996:NAC

[Rig96] Maria Lynn Jason Rightley. Numerical analysis of carbon monoxide flames by reduced kinetic and asymptotic methods. Thesis (ph.d.), University of California, San Diego, San Diego, CA, USA, 1996.

Rihm:1996:PES

[Rih96] R. Rihm. Problems in enclosing solutions of ODEs. In Alefeld and Herzberger [AH96], pages 297–210. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Rengarajan:1993:CSN

[RKA93] Govind Rengarajan, Norman F.

Knight, and Mohammad A. Aminpour. Comparison of symbolic and numerical integration methods for an assumed-stress hybrid shell element. National Technical Information Service, Washington, DC, USA, 1993. ???? pp. Shipping list no.: 95-0048-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1994] 1 microfiche.

Roca:1997:SAH

[Roc97] P. Roca, editor. Structural analysis of historical constructions: possibilities of numerical and experimental techniques, Theory and engineering applications of computational methods. CIMNE, Barcelona, Spain, 1997. ISBN 84-87867-66-9. LCCN TA646 S769 1997. Papers presented at the International Seminar on Structural Analysis of Historical Constructions, held November 8–10, 1995 in Barcelona.

Rojas:1996:ANR

[Roj96a] J. Maurice Rojas. On the average number of real roots of certain random sparse polynomial systems. In Renegar et al. [RSS96], pages 689–700. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Rojiani:1996:PCN

[Roj96b] Kamal B. Rojiani. Programming in C with numerical methods for engineers. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-726498-4. xviii + 1130 pp. LCCN QA76.73.C15 R65 1996.

Rosenberg:1990:NMF

[Ros90] Joseph Robert Rosenberg. Numerical methods for fluids with hyperbolic constitutive equations. Thesis (ph.d.) in chemical engineering, Dept. of Chemical Engineering, University of California, Berkeley, Berkeley, CA, USA, May 1990. vi + 180 pp.

Roy:1996:CRA

[Roy96] Marie-Françoise Roy. Computations in real algebraic geometry. In Renegar et al. [RSS96], pages 701–714. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Rappaz:1998:IAN

[RP98] Jacques Rappaz and Marco Picasso. Introduction à l’analyse numérique. (French) [Introduction to numerical analysis]. Mathématiques. [Mathematics].

Presses Polytechniques et Universitaires Romandes, Lausanne, Switzerland, 1998. ISBN 2-88074-363-X. x + 256 pp.

Ratschek:1996:DES

[RR96] H. Ratschek and J. Rokne. Determination of the exact sign of sum. In Alefeld and Herzberger [AH96], pages 198–204. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995. Rodriguez-Roa:1990:ANP

[RRLA90] Fernando Rodriguez-Roa, Guillermo Noguera Larrain, and Juan L. Osorio Aguayo. Analisis numerico de presas de tierra. Technical Report 90-7, Departamento de Ingenieria Estructural, Pontificia Universidad Catolica de Chile, ??, Chile, July 1990. 211 pp. Proyectos de investigacion DIUC no. 7/86 FONDECYT no. 88/0267.

Rizzi:1992:PNG

[RRV92] Arthur Rizzi, Inge L. Ryhming, and Jan B. Vos, editors. Proceedings of the Ninth GAMM Conference on Numerical Methods in Fluid Mechanics, Lausanne, September 25– 27, 1991, volume 35 of Notes on numerical fluid mechanics. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1992. ISBN 3-528-07635-6. ISSN 0179-9614. LCCN TA 357 G35 1991.

Rao:1992:NMPa

[RS92a] S. Balachandra Rao and C. K.

Shantha. Numerical Methods With Programs in Basic, Fortran and Pascal. Stosius Inc/Advent Books Division, London, UK, December 1992. ISBN 0-86311-370-2 (paperback). xiii + 446 pp. US$20.00. URL http:/

/www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0863113702.

Rao:1992:NMPb

[RS92b] S. Balachandra Rao and C. K. Shantha. Numerical methods: with programs in BASIC, FOR-TRAN and Pascal. Universities Press (India), Hyderabad, India, 1992. ISBN 0-86311-370-2. xiii + 446 pp.

Renegar:1996:MNA

[RSS96] James Renegar, Michael Shub, and Steve Smale, editors. The mathematics of numerical analysis: 1995 AMS-SIAM Summer Seminar in Applied Mathematics, July 17–August 11, 1995, Park City, Utah, volume 32 of Lectures in Applied Mathematics — American Mathematical Society. American Mathematical Society, Providence, RI, USA, 1996. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Roos:1996:NMS

[RST96] Hans-Gorg Roos, M. Stynes, and L. Tobiska. Numerical methods for singularly perturbed differential equations: convectiondiffusion and flow problems, volume 24 of Springer series in computational mathematics.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-60718-8. xvi + 348 pp. LCCN QA377 .R66 1996 Bar.

Rogers:1997:NMF

[RT97] L. C. G. Rogers and D. Talay.

Numerical methods in finance. Publications of the Newton Institute. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-57354-8 (hardback). x + 326 pp. LCCN HG174 .N86 1997.

Ritter:1996:ACC

[RW96] Klaus Ritter and Grzegorz W.

Wasilkowski. On the average case complexity of solving Poisson equations. In Renegar et al. [RSS96], pages 677–688. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Ryabenkii:1994:VVV

[Rya94] V. S. Ryaben′kiĭ. Vvedenie v vychislitelnuyu matematiku. (Russian) [Introduction to computational mathematics]. Fizmatlit “Nauka”, Moscow, Russia, 1994. ISBN 5-02-014544-0. 335 pp.

Saad:1992:NML

[Saa92] Youcef Saad. Numerical methods for large eigenvalue problems. Algorithms and Architectures for Advanced Scientific Computing. Halsted Press, New York, USA, 1992. ISBN 0-7190-3386-1 (UK), 0-470-21820-7 (US). xii + 346 pp. LCCN QA188 .S18 1992.

Salvi:1998:NET

[Sal98] Rodolfo Salvi. Navier–Stokes equations: theory and numerical methods, volume 388 of Pitman research notes in mathematics series. Longman Scientific and Technical, Harlow, Essex, UK, 1998. ISBN 0-582-35643-1 (paperback). ISSN 0269-3674. 347 pp. LCCN QA911 .N38 1998. Conference proceedings of the “International Conference on Navier– Stokes Equations: Theory and Numerical Methods” held in Villa Monastero, 2–6 June 1997, at Varenna, Lecco.

Salomon:1999:CGG

[Sal99] David (David) Salomon. Computer graphics and geometric modeling. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98682-0. xviii + 851 pp. LCCN T385 .S243 1999. URL

http://www.loc.gov/catdir/
enhancements/fy0816/98033424-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0816/
98033424-t.html.

Samoilenko:1996:ACV

[Sam96] M. N. Samoilenko. On approximation of characteristic values of meromorphic operatorvalued functions. In Alefeld and Herzberger [AH96], pages 211–216. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Shampine:1997:FNC

[SAP97] Lawrence F. Shampine, Richard C.

Allen, Jr., and Steven Pruess. Fundamentals of numerical computing. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1997. ISBN 0-471-16363-5 (hardcover). x + 268 pp. LCCN QA297 .A52 1997. URL

http://www.loc.gov/catdir/
description/wiley033/96022074.
html;   http://www.loc.gov/
catdir/toc/onix04/96022074.
html.

Saran:1992:NMS

[Sar92] M. J. Saran, editor. Numerical methods for simulation of industrial metal forming processes: presented at the Winter Annual Meeting of the American Society of Mechanical Engineers, Anaheim, California, November 8– 13, 1992, volume 156 of AMD. American Society of Mechanical Engineers, United Engineering Center, 345 E. 47th St., New York, NY 10017, USA, 1992. ISBN 0-7918-1133-6 (paperback). LCCN TS205 .N85 1992.

Stoer:1990:NM

[SB90] Josef Stoer and Roland Bulirsch. Numerische Mathematik.

2. (German) [Numerical mathematics. 2]. Springer-Lehrbuch. [Springer Textbook]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1990. ISBN 3-540-51482-1. xiv + 341 pp. Eine Einführung—unter Berücksichtigung von Vorlesungen von F. L. Bauer. [An introduction, with reference to lectures by F. L. Bauer].

Stoer:1993:INA

[SB93] Josef Stoer and Roland Bulirsch. Introduction to numerical analysis, volume 12 of Texts in Applied Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1993. ISBN 0-387-97878-X (New York), 3-540-97878-X (Berlin). xiii + 660 pp. LCCN QA297 .S8213 1993. Translated from the German by R. Bartels, W. Gautschi and C. Witzgall.

Stoer:1996:INA

[SB96a] Josef Stoer and Roland Bulirsch. Introduction to numerical analysis, volume 12 of Texts in applied mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1996. ISBN 0-387-97878-X (New York), 3-540-97878-X (Berlin). xiii + 660 pp. LCCN QA297 .S8213 1996.

Strickland:1996:POF

[SB96b] James H. Strickland and Roy S. Baty. A pragmatic overview of fast multipole methods. In Renegar et al. [RSS96], pages 807–830. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Smith:1996:DDP

[SBG96a] Barry F. Smith, Petter E.

Björstad, and William Gropp. Domain decomposition: parallel multilevel methods for elliptic partial differential equations. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-49589-X (hardcover). xii + 224 pp. LCCN QA377 .S586 1996.

URL http://www.loc.gov/
catdir/description/cam027/
95042362.html;   http://www.
loc.gov/catdir/toc/cam023/
95042362.html.

Smith:1996:DD

[SBG96b] Barry F. Smith, Petter E.

Bjørstad, and William D. Gropp. Domain decomposition. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-49589-X. xii + 224 pp. Parallel multilevel methods for elliptic partial differential equations.

Stenerson:1996:CNC

[SC96] Jon Stenerson and Kelly Curran. Computer numerical control: operation and programming. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-348962-0 (hardcover). xii + 388 pp. LCCN TJ1189 .S72 1997.

Schutz:1980:GMM

[Sch80] Bernard F. Schutz. Geometrical methods of mathematical physics. Cambridge University Press, Cambridge, UK, 1980. ISBN 0-521-23271-6, 0-521-29887-3 (paperback). xii + 250 pp. LCCN QC20.7.D52 S34.

URL http://www.loc.gov/
catdir/description/cam022/
80040211.html;   http://www.
loc.gov/catdir/enhancements/
fy0729/80040211-b.html;
http://www.loc.gov/catdir/
toc/cam029/80040211.html.

Scheid:1990:SPN

[Sch90] Francis J. Scheid. 2000 solved problems in numerical analysis. Schaum’s solved problems series. McGraw-Hill, New York, NY, USA, 1990. ISBN 0-07-055233-9. iv + 698 pp. LCCN QA297.3 .S34 1990.

Schatzman:1991:AN

[Sch91a] Michelle Schatzman. Analyse numérique. (French) [Numerical analysis]. InterEditions, Paris, France, 1991. ISBN 2-7296-0373-5. 358 pp. Cours et exercices pour la licence. [Course and exercises for the bachelor’s degree].

Schiesser:1991:NML

[Sch91b] W. E. Schiesser. The numerical method of lines: integration of partial differential equations. Academic Press, New York, NY, USA, 1991. ISBN 0-12-624130-9. xiii + 326 pp. LCCN QA377 .S354 1991.

URL http://www.loc.gov/
catdir/description/els032/
90025389.html;   http://www.
loc.gov/catdir/toc/els031/

90025389.html. Integration of partial differential equations.

Schneider:1991:NM

[Sch91c]  M. Schneider.  Numerische

Mathematik. (German) [Numerical mathematics], volume 10 of Mathematik für Lehrer [Mathematics for Teachers]. Deutscher Verlag der Wissenschaften, Berlin, Germany, 1991. ISBN 3-326-00591-1. 165 pp. Studienbücherei. [Textbook Library].

Schwarz:1991:MFE

[Sch91d] Hans-Rudolf Schwarz. Methode der finiten Elemente. (German) [Finite-element method], volume 47 of Leitfäden der Angewandten Mathematik und Mechanik [Guides to Applied Mathematics and Mechanics]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, third edition, 1991. ISBN 3-519-22349-X. 435 pp. Eine Einführung unter besonderer Berücksichtigung der Rechenpraxis. [An introduction with special reference to computational practice], Teubner Studienbücher Mathematik. [Teubner Mathematical Textbooks].

Schwarz:1993:NM

[Sch93]  Hans-Rudolf Schwarz.  Nu-

merische Mathematik. (German) [Numerical mathematics]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, third edition, 1993. ISBN 3-519-22960-9. 575 pp. With a contribution by Jörg Waldvogel.

Schwartz:1996:TIN

[Sch96] Adam Lowell Schwartz. Theory and implementation of numerical methods based on Runge– Kutta integration for solving optimal control problems. Thesis (ph.d.) in electrical engineering and computer sciences, University of California, Berkeley, Berkeley, CA, USA, May 1996. viii + 276 pp. Electronics Research Laboratory Memorandum no. UCB/ERL M96/18.

Schwarz:1997:NM

[Sch97] Hans Rudolf Schwarz. Numerische Mathematik. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, fourth edition, 1997. ISBN 3-519-32960-3. 653 pp. With a contribution by Jörg Waldvogel.

Shub:1996:PDD

[SCI^{+}96] M. Shub, Alexandre Chorin, Arieh Iserles, Beresford Parlett, Steve Smale, and Shmuel Winograd. Panel discussion: Does numerical analysis need a model of computation? In Renegar et al. [RSS96], pages 1–18. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Shen:1995:SMP

[SD95] Shan-Fu Shen and Paul R. Dawson. Simulation of materials processing: theory, methods and applications: proceedings of the Fifth International Conference on Numerical Methods in Industrial Forming Processes — NUMIFORM ’95, Ithaca/New York/USA/18–21 June, 1995. A. A. Balkema, Brookfield, VT, USA, 1995. ISBN 90-5410-553-4. xviii + 1221 pp. LCCN TS205 .I57 1995.

Sedgewick:1992:AC

[Sed92] Robert Sedgewick. Algorithms in C++. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-36118-3, 0-201-51059-6. xiv + 656 pp. LCCN QA76.73.C153 S38 1992.

Seshasayee:1998:NSF

[Ses98] Srikumar Seshasayee. Numerical studies of flows in microchannels, using finite difference and finite element methods. Thesis (m.s.), University of California, Santa Barbara, Santa Barbara, CA, USA, 1998.

Smith:1998:PFE

[SG98] I. M. (Ian Moffat) Smith and D. V. Griffiths. Programming the finite element method. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, third edition, 1998. ISBN 0-471-96542-1 (hardcover), 0-471-96543-X (paperback). xii + 534 pp. LCCN TA347.F5 S64 1997. URL

http://www.loc.gov/catdir/
description/wiley035/96054905.
html;   http://www.loc.gov/
catdir/enhancements/fy0609/
96054905-b.html;   http://
www.loc.gov/catdir/toc/wiley022/
96054905.html.

Stuart:1996:DSN

[SH96] A. M. Stuart and A. R. Humphries. Dynamical systems and numerical analysis, volume 2 of Cambridge monographs on applied and computational mathematics. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-49672-1. xxii + 685 pp. LCCN QA614.8 .S78 1996.

Suli:1997:FEM

[SH97] Endre Süli and Paul Houston. Finite element methods for hyperbolic problems: a posteriori error analysis and adaptivity. In Duff and Watson [DW97], pages 441–472. ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

                                        //global.oup.com/academic/
                                        product/the-state-of-the-
                                        art-in-numerical-analysis-
                                        9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Shary:1996:ASI

[Sha96] Sergey P. Shary. Algebraic solutions to interval linear equations and their applications. In Alefeld and Herzberger [AH96], pages 224–233. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Shahi:1998:CVM

[Sha98] Avantika Shahi. A cybertheatre a virtual movie zone designed using VRML and JavaScript. ????, ????, 1998. 44 pp.

Shearer:1991:VPN

[She91] Michael Shearer, editor. Viscous profiles and numerical methods for shock waves. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1991. ISBN 0-89871-283-1. LCCN QA927 .V57 1991. Proceedings of a workshop at North Carolina State University, Raleigh, North Carolina, May 23–25, 1990.

Shinnerl:1995:KIM

[Shi95] Joseph R. Shinnerl. KKT-based interior-point methods for numerical optimization. Thesis (ph.d.), University of California, San Diego, San Diego, CA, USA, 1995. xiii + 228 pp.

Shokin:1997:NRG

[Sho97] Y. I. Shokin. New resourcesparing grid methods for solving the problems of mathematical physics. In Boisvert [Boi97], pages 361–374. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Shu:1990:NMS

[Shu90] Chi-Wang Shu. Numerical methods for systems of conservation laws of mixed type using flux splitting. Institute for Computer Applications in Science and Engineering, NASA Langley Research Center, Hampton, VA, USA, 1990. ???? pp. Shipping list no.: 92-0600-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration], 1990. 1 microfiche.

Sieverding:1994:NMF

[Sie94] C. H. Sieverding. Numerical methods for flow calculation in turbomachines: May 16–20, 1994, volume 1994-06 of Lecture series / Von Karman Institute for Fluid Dynamics. Von Karman Institute for Fluid Dynamics, Rhode Saint Genese, Belgium, 1994. ISSN 0377-8312. various pp. LCCN TJ267 .N852 1994.

Siegelmann:1996:CNN

[Sie96] Hava T. Siegelmann. Computability with neural networks. In Renegar et al. [RSS96], pages 733–748. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Sinclair:1996:TFP

[Sin96]  R. Sinclair.  Tight floating

point error bounds for polynomial evaluation. In Alefeld and Herzberger [AH96], pages 234–240. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Sancaktar:1996:RSA

[SJ96] Erol Sancaktar and Shahriar Jahanian. Reliability, stress analysis, and failure prevention issues in fastening and joining, composite and smart structures, numerical and FEA methods, and risk minimization: presented at the 1996 ASME International Mechanical Engineering Congress and Exposition, November 17–22, 1996, Atlanta, Georgia, volume 92 of DE. American Society of Mechanical Engineers, United Engineering Center, 345 E. 47th St., New York, NY 10017, USA, 1996. ISBN 0-7918-1546-3. v + 133 pp. LCCN TA169 .R4473 1996.

Skeel:1993:ENC

[SK93] Robert D. Skeel and Jerry B.

Keiper. Elementary Numerical Computing with Mathematica. McGraw-Hill computer science series. McGraw-Hill, New York, NY, USA, 1993. ISBN 0-07-057820-6. xiv + 434 pp. LCCN QA297 .S545 1993. See [SK95].

Skeel:1995:ENC

[SK95] Robert D. Skeel and Jerry B.

Keiper. Elementary Numerical Computing with Mathematica. Kyoritsu Suppan, ??, Japan, 1995. ISBN 4-320-01488-X. 410 pp. LCCN ???? Japanese translation of [SK93].

Slawik:1996:PEE

[SK96] Lech Sławik and Andrzej Karafiat. A posteriori error estimates for operator equations. In Alefeld and Herzberger [AH96], pages 241–248. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455

Saigal:1990:SAO

[SM90] Sunil Saigal and Subrata Mukherjee, editors. Sensitivity analysis and optimization with numerical methods: presented at the Winter Annual Meeting of the American Society of Mechanical Engineers, Dallas, Texas, November 25–30, 1990, volume 115 of AMD. American Society of Mechanical Engineers, United Engineering Center, 345 E. 47th St., New York, NY 10017, USA, 1990. ISBN 0-7918-0595-6. LCCN TA658.2 .S47 1990.

Snyder:1991:NAS

[SM91] Willard M. Snyder and Richard H. McCuen. Numerical analysis with sliding polynomials. Lighthouse Publications, Mission Viejo, Calif., 1991. ISBN 0-914055-09-7. ix + 561 pp. LCCN QA281 .S57 1991.

Smelyakov:1996:IAS

[Sme96] S. V. Smelyakov. Interval approach for super-long-term forecasting of sunspot activity maxima. In Alefeld and Herzberger [AH96], pages 255–260. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Smith:1993:NMI

[Smi93] H. V. Smith. Numerical methods of integration. Studentlitteratur, Lund, Sweden, 1993. ISBN 0-86238-331-5, 91-44-38851-9. 147 pp. LCCN QA299.3.N845 1993.

Smith:1994:NMG

[Smi94] I. M. Smith. Numerical methods in geotechnical engineering: proceedings of the Third European Conference on Numerical Methods in Geotechnical Engineering: ECONMIG 94, Manchester, UK, 7–9 September 1994. A. A. Balkema, Brookfield, VT, USA, 1994. ISBN 90-5410-510-0. xi + 432 pp. LCCN TA703.5 .E78 1994. Conference sponsored by the International Society for Soil Mechanics and Foundation Engineering; organised by European Technical Committee on Numerical Methods in Geotechnical Engineering.

Snyder:1997:TFO

[Sny97] W. V. Snyder. Testing functions of one and two arguments. In Boisvert [Boi97], pages 155– 166. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Sol:1997:MIU

[SO97] H. Sol and C. W. J. Oomens.

Material identification using mixed numerical experimental methods: proceedings of the EU-ROMECH Colloquium held in Kerkrade, The Netherlands, 7– 9 April 1997. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1997. ISBN 0-7923-4779-X. ix + 240 pp. LCCN TA418.9.C6 E947 1997.

Sobol:1991:MCM

[Sob91] I. M. Sobol′. Die Monte-Carlo-Methode. (German) [The Monte Carlo Method]. Deutscher Verlag der Wissenschaften, Berlin, Germany, 1991. ISBN 3-326-00635-7. 108 pp. Translated from the fourth (1985) Russian edition by Klaus Lommatzsch and A. Matthes.

Sobol:1994:PMC

[Sob94] Ilya M. Sobol′. A primer for the Monte Carlo method. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1994. ISBN 0-8493-8673-X. xiv + 107 pp. LCCN QA298 .S6613 1994. URL

http://www.loc.gov/catdir/
enhancements/fy0738/93050716-
d.html.

Spath:1996:OSD

[Spä96] Helmuth Späth. Orthogonal squared distance fitting with parabolas. In Alefeld and Herzberger [AH96], pages 261– 269. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Shampine:1997:DOS

[SR97] L. F. Shampine and M. W. Reichelt. Developing ODE software in new computing environments. In Boisvert [Boi97], pages 224– 235. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Stewart:1990:MPT

[SS90] Gilbert Wright (Gilbert W.) Stewart and Ji-guang Sun. Matrix Perturbation Theory. Computer science and scientific computing. Academic Press, New York, NY, USA, 1990. ISBN 0-12-670230-6 (hardcover). xv + 365 pp. LCCN QA871 .S775 1990.

URL http://www.loc.gov/
catdir/description/els032/
90033378.htm;   http://www.
loc.gov/catdir/description/
els032/90033378.html;   http:
//www.loc.gov/catdir/toc/
els031/90033378.htm;   http:
//www.loc.gov/catdir/toc/
els031/90033378.html.

Sukharev:1992:MMT

[SS92] A. G. (Aleksei Grigor’evich)

Sukharev and A. G. Sukharev. Minimax models in the theory of numerical methods, volume 21 of Theory and decision library. Series B, Mathematical and statistical methods. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1992. ISBN 0-7923-1821-8. xiv + 256 pp. LCCN QA300 .S9613 1992. Translation of: Minimaksnye algoritmy v zadachakh chislennogo analiza. s.

Schulte:1996:CAR

[SS96a] Michael J. Schulte and Earl E.

Swartzlander, Jr. A coprocessor for accurate and reliable computing. In Alefeld and Herzberger [AH96], pages 217–223. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Singer:1996:FPE

[SS96b] Sanja Singer and Saša Singer.

Floating-point errors in complex Cholesky algorithms. In Alefeld and Herzberger [AH96], pages 249–254. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Sanz-Serna:1997:GI

[SS97a] J. M. Sanz-Serna. Geometric integration. In Duff and Watson [DW97], pages 121–144. ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

                                        //global.oup.com/academic/
                                        product/the-state-of-the-
                                        art-in-numerical-analysis-
                                        9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Schiesser:1997:CTP

[SS97b] W. E. Schiesser and C. A. Silebi. Computational transport phenomena: numerical methods for the solution of transport problems. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-55378-4 (hardbound), 0-521-55653-8 (paperback). x + 457 pp. LCCN QA929 .S35 1997.

Shanno:1997:IPM

[SS97c] David F. Shanno and Evangelia M. Simantiraki. Interior point methods for linear and nonlinear programming. In Duff and Watson [DW97], pages 339– 362. ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

                                        //global.oup.com/academic/

product/the-state-of-the-
art-in-numerical-analysis-
9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Schoenauer:1990:CDU

[SSES90] W. Schoenauer, A. Schreiner, and J. Ecke-Schueth. A critical discussion of the user interface of a PDE solver in the context of numerical expert systems. In Houstis et al. [HRV90], pages 353–360. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Sigmon:1998:MP

[ST98] Kermit Sigmon and The Math-Works Staff. MATLAB Primer. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, fifth edition, 1998. ISBN 0-8493-1305-8. 130 pp. LCCN QA297.S4787 1998. US$14.95.

Stewart:1990:IMS

[Ste90] G. W. (Pete) Stewart. An iterative method for solving linear inequalities. In Cox and Hammarling [CH90], pages 241– 247. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Stenger:1993:NMB

[Ste93] Frank Stenger. Numerical methods based on Sinc and analytic functions, volume 20 of Springer Series in Computational Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-94008-1 (New York), 3-540-94008-1 (Berlin). xv + 565 pp. LCCN QA372 .S82 1993.

Stewart:1996:ANA

[Ste96] G. W. (Gilbert W.) Stewart. Afternotes on numerical analysis: a series of lectures on elementary numerical analysis presented at the University of Maryland at College Park and recorded after the fact. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1996. ISBN 0-89871-362-5 (paperback). x + 200 pp. LCCN QA297 .S785 1996. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0726/95047768-
                                        d.html;   http://www.loc.
                                        gov/catdir/enhancements/fy0726/
                                        95047768-t.html.

Stewart:1998:AGG

[Ste98a] G. W. (Gilbert W.) Stewart. Afternotes goes to graduate school: lectures on advanced numerical analysis. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1998. ISBN 0-89871-404-4 (paperback). xii + 245 pp. LCCN QA297 .S784 1998. URL

                                        http://www.loc.gov/catdir/

enhancements/fy0726/97044449-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0726/
97044449-t.html. A series of

lectures on advanced numerical analysis presented at the University of Maryland at College Park and recorded after the fact.

Stewart:1998:MAV

[Ste98b] G. W. (Gilbert W.) Stewart.

Matrix algorithms. Volume 1. Basic decompositions. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1998. ISBN 0-89871-414-

1. xix + 458 pp. LCCN QA188 .S714 1998. Basic decompositions.

Schwetlick:1996:PFL

[STL96] Hubert Schwetlick, Gisela Timmermann, and Rralf Losche. Path following for large nonlinear equations by implicit block elimination based on recursive projections. In Renegar et al. [RSS96], pages 715–732. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Stoer:1993:NM

[Sto93] Josef Stoer. Numerische Mathematik. 1. (German) [Numerical mathematics. 1]. Springer-Lehrbuch. [Springer Textbook]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., sixth edition, 1993. ISBN 3-540-56213-3. xiv + 314 pp. Eine Einführung— unter Berücksichtigung von Vorlesungen von F. L. Bauer. [An introduction—based on the lectures of F. L. Bauer].

Stoer:1994:NM

[Sto94] Josef Stoer. Numerische Mathematik. 1. (German) [Numerical mathematics. 1]. Springer-Lehrbuch. [Springer Textbook]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., seventh edition, 1994. ISBN 3-540-57823-4. xii + 367 pp. Eine Einführung— unter Berücksichtigung von Vorlesungen von F. L. Bauer. [An introduction—based on the lectures of F. L. Bauer].

Stoyan:1996:EIS

[Sto96] Yu. G. Stoyan. Extended interval space and elementary mappings. In Alefeld and Herzberger [AH96], pages 270–279. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Strang:1996:WFB

[Str96] Gilbert Strang. Wavelets from filter banks. In Renegar et al. [RSS96], pages 765–806. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Stuart:1997:CSN

[Stu97] Andrew Stuart. Convergence and stability in the numerical approximation of dynamical systems. In Duff and Watson [DW97], pages 145–170. ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

                                        //global.oup.com/academic/

product/the-state-of-the-
art-in-numerical-analysis-
9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Strehmel:1995:NGD

[SW95] Karl Strehmel and Rüdiger Weiner. Numerik gewöhnlicher Differentialgleichungen. (German) [Numerical Solution of Ordinary Differential Equations]. Teubner Studienbücher Mathematik. [Teubner Mathematical Textbooks]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1995. ISBN 3-519-02097-1. 462 pp.

Sommese:1996:NAG

[SW96] Andrew J. Sommese and Charles W.

Wampler. Numerical algebraic geometry. In Renegar et al. [RSS96], pages 749–764. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Schwalbe:1997:VVD

[SW97] Dan Schwalbe and S. Wagon.

VisualDSolve: visualizing differential equations with Mathematica. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1997. ISBN 0-387-94721-3. xiv + 271 pp. LCCN QA 371.5 D37 S39

Szilard:1992:NTD

[Szi92] Ronaldo Henriques Szilard. Numerical transport and diffusion methods in radiative transfer. Thesis (ph.d.), University of California, Los Angeles, Los Angeles, CA, USA, 1992. xix + 252 pp.

LeTallec:1990:NAV

[Tal90] Patrick Le Tallec. Numerical analysis of viscoelastic problems, volume 15 of Recherches en mathématiques appliquées. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 2-225-82093-7 (Mason), 3-540-52450-9 (Springer-Verlag), 0-387-52450-9 (Springer-Verlag). ISSN 0298-3168. 136 pp. LCCN QA931 .L44 1990.

Tao:1994:TAN

[Tao94] Yue-Qun Tao. Theoretical analysis and numerical simulation of plasma behavior in a strongly sheared electric field. Thesis (ph.d.), University of California, Los Angeles, Los Angeles, CA, USA, 1994. xvi + 85 pp.

Trefethen:1997:NLA

[TB97] Lloyd N. (Lloyd Nicholas) Trefethen and David Bau III. Numerical Linear Algebra. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1997. ISBN 0-89871-361-7. xii + 361 pp. LCCN QA184 .T74 1997. URL

                                        http://www.loc.gov/catdir/

enhancements/fy0665/96052458-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0665/
96052458-t.html.

Toro:1998:NMW

[TC98] E. F. Toro and J. F. Clarke, editors. Numerical methods for wave propagation: selected contributions from the workshop held in Manchester, UK, containing the Harten memorial lecture, volume 47 of Fluid mechanics and its applications. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1998. ISBN 0-7923-5125-8. LCCN QA927 .N86 1998. Contributions from the International Workshop on Numerical Methods for Wave Propagation Phenomena, held in May 1995.

Tesar:1993:DWB

[Tes93]  Alexander Tesar.  Develop-

ment of wood bridges: numerical and experimental analysis of static and dynamic behaviour of wood bridges. Helsinki University of Technology, Espoo, Finland, 1993. 47 pp. LCCN PB95-149662 Microfiche. Performed by the Helsinki University of Technology, Dept. of Structural Engineering, Laboratory of Bridge Engineering. Publication no. 6.

Tornig:1991:NLP

[TGK91] Willi Törnig, Michael Gipser, and Bernhard Kaspar. Numerische Lösung von partiellen Differentialgleichungen der Technik. (German) [Numerical solution of partial differential equations of technology], volume 1 of Mathematische Methoden in der Technik [Mathematical Methods in Technology]. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, second edition, 1991. ISBN 3-519-12613-3. 183 pp. Differenzenverfahren, finite Elemente und die Behandlung großer Gleichungssysteme. [Difference methods, finite elements and the treatment of large systems of equations].

MathWorks:1997:SEM

[The97] The MathWorks, Inc. The Student Edition of Matlab: High-Performance Numeric Computation and Visualization Software: Version 5 User’s Guide. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN ???? ???? pp. LCCN ????

Thomas:1990:NCS

[Tho90] William A. Thomas. Nonconnah Creek sedimentation study: Analysis using a numerical modeling approach. National Technical Information Service, Washington, DC, USA, January 1990. 63 pp.

Thomas:1995:NPD

[Tho95] J. W. (James William) Thomas. Numerical partial differential equations: [Part 1]: finite difference methods, volume 22 of Texts in Applied Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-97999-9 (vol. 1). xx + 437 pp. LCCN QA377 .T495 1995. URL

http://www.loc.gov/catdir/
enhancements/fy0815/95017143-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0815/
95017143-t.html.

Thomas:1999:NPD

[Tho99] J. W. (James William) Thomas.

Numerical partial differential equations: Part 2: Conservation Laws and Elliptic Equations, volume 33 of Texts in Applied Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98346-5 (vol. 2). xxii + 556 pp. LCCN QA377 .T495 1995. URL

http://www.loc.gov/catdir/
enhancements/fy0815/95017143-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0815/
95017143-t.html. Conserva-

tion laws and elliptic equations.

Tikhonov:1990:CMR

[Tik90]  A. N. Tikhonov.  Chislen-

nye metody resheniia nekorrektnykh zadach. Nauka (Glav. red. fiziko-matematicheskoi litry), Moscow, USSR, 1990. ISBN 5-02-014135-6. 229 pp. LCCN QA377 .C48 1990. 2.40rub. Nauchnoe izdanie. Title on p. 228: Numerical methods for the solution of ill-posed problems. Summary and table of contents also in English.

Tikhonov:1995:NMS

[Tik95] A. N. Tikhonov. Numerical methods for the solution of illposed problems, volume 328 of Mathematics and its applications. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1995. ISBN 0-7923-3583-X (hardcover). ix + 253 pp. LCCN QA297 .C48313 1995.

Tang:1992:JF

[TJL92] Zhen Tang, Jian Ming Jin, and Zhi Jie Li. Jisuan fangfa. (Mandarin Chinese) [Calculation Method]. Gaodeng Xuexiao Shiyong Jiaocai. [Trial Educational Materials for Advanced Schools]. Gaodeng Jiaoyu Chubanshe (Higher Education Press), Beijing, People’s Republic of China, 1992. ISBN 7-04-003907-9. viii + 398 pp.

Todd:1962:SNA

[Tod62] John Todd, editor. Survey of Numerical Analysis. Mc-Graw-Hill, New York, NY, USA, 1962. xvi + 589 pp. LCCN QA297 .T63. URL https:/

                                        /babel.hathitrust.org/cgi/
                                        pt?id=mdp.39015000983901.

Todd:1978:BNM

[Tod78] John Todd. Basic Numerical Mathematics, volume 14, 22 of International series of numerical mathematics. Academic Press, New York, NY, USA, 1978. ISBN 0-12-692401-5 (vol. 1), 0-12-692402-3 (vol. 2). 253 (vol. 1, est.), 250 (vol. 2, est.) pp. LCCN QA297 .T58 1978.

Todd:1990:PEH

[Tod90] John Todd. The prehistory and early history of computation at the U.S. National Bureau of Standards. In Nash [Nas90b], pages 251–268. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Tong:1990:CSM

[Ton90] Sin Shing Tong. Coupling symbolic manipulation and numerical simulation for complex engineering designs. In Houstis et al. [HRV90], pages 241–256. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Toro:1997:RSN

[Tor97] E. F. Toro. Riemann solvers and numerical methods for fluid dynamics: a practical introduction. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-61676-4 (Berlin). xviii + 592 pp. LCCN QA911 .T66 1997.

Toro:1999:RSN

[Tor99] E. F. Toro. Riemann solvers and numerical methods for fluid dynamics: a practical introduction. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1999. ISBN 3-540-65966-8 (hardcover). xix + 624 pp. LCCN QA911 .T66 1999.

Terrano:1990:APP

[TPD90] A. E. Terrano, J. E. Peters, and S. M. Dunn. Automatic partitioning of PDE solvers for parallel computers. In Houstis et al. [HRV90], pages 75–86. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Tsuchiya:1991:NMR

[Tsu91] Nobushige Tsuchiya. Numerical methods for reaction diffusion equation on the torus. Thesis (m.a.), University of California, Santa Cruz, Santa Cruz, CA, USA, 1991. iv + 72 pp.

Turner:1990:GNA

[Tur90] P. R. (Peter R.) Turner. Guide to numerical analysis. CRC mathematical guides. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1990. ISBN 0-8493-7701-3. x + 208 pp. LCCN QA297 .T87 1990.

Turcotte:1998:CAM

[TW98a] Louis H. Turcotte and H. B. (Howard B.) Wilson. Computer applications in mechanics of materials using MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-749060-7. xi + 321 pp. LCCN TA405 .T787 1998.

Tveito:1998:IPD

[TW98b] Aslak Tveito and Ragnar Winther. Introduction to partial differential equations: a computational approach, volume 29 of Texts in Applied Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98327-9 (hardcover). xv + 392 pp. LCCN QA377 .T9 1998. A computational approach.

Tyrtyshnikov:1997:BIN

[Tyr97] E. E. (Evgenii Evgen’evich)

Tyrtyshnikov. A brief introduction to numerical analysis. Birkhäuser Boston Inc., Cambridge, MA, USA, 1997. ISBN 0-8176-3916-0 (hardcover), 3-7643-3916-0. xii + 202 pp. LCCN QA297 .T97 1997.

Ueberhuber:1997:NCM

[Ueb97] Christoph W. Ueberhuber. Numerical Computation: Methods, Software, and Analysis. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-62058-3 (vol. 1: softcover), 3-540-62057-5 (vol. 2: softcover), 3-642-59118-3 (e-book). xvi + 474 (vol. 1), xvi + 495 (vol. 2) pp. LCCN QA297 .U2413 1997. US$44.95 (vol. 1), US$49.95 (vol. 2).

Ullrich:1990:CAS

[Ull90a] Christian Ullrich, editor. Computer arithmetic and selfvalidating numerical methods, volume 7 of Notes and reports in mathematics in science and engineering. Academic Press, New York, NY, USA, 1990. ISBN 0-12-708245-X. LCCN QA76.9.C62 C663 1990. Proceedings of an international conference held Oct. 2–6, 1989, Basel.

Ullrich:1990:CCA

[Ull90b] Christian Ullrich, editor. Contributions to Computer Arithmetic and Self-Validating Numerical Methods. (Proceedings of SCAN 89, held in Basel, Oct. 2–6, 1989), volume 7 of IMACS annals on computing and applied mathematics. J. C. Baltzer AG, Scientific Publishing Company, Basel, Switzerland, 1990. ISBN ???? LCCN QA76.9.C62 C664 1990.

Umetani:1997:VDN

[Ume97] Y. Umetani. The visual diagnosis on numerical calculation of PDE problems and experiments. In Boisvert [Boi97], pages 333– 344. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

USGS:1998:AMN

[Uni98a] United States. Geological Survey. Analytical Methods, Numerical Modeling, and Monitoring Strategies for Evaluating the Effects of Ground-Water Withdrawals, etc. U.S. Geological Survey, Denver, CO, USA, 1998. Shipping List #: 98-0307-P Shipping List Date: 07/23/98.

USNASA:1998:CSN

[Uni98b] United States. National Aeronautics and Space Administration. Comparison of Several Numerical Methods for Simulation of Compressible Shear Layers. NASA, Washington, DC, USA, June 26, 1998. Shipping List #: 99-0723-M Shipping List Date: 08/02/99.

USNASA:1998:NAB

[Uni98c] United States. National Aeronautics and Space Administration. Numerical Analysis of Base Flowfield for a Four-Engine Clustered Nozzle Configuration. NASA, Washington, DC, USA, September 16, 1998. Shipping List #: 99-0055-M Shipping List Date: 11/23/98.

USNASA:1998:NPS

[Uni98d] United States. National Aeronautics and Space Administration. Numerical Propulsion System Simulation Software Standards and Procedures Specification. NASA, Washington, DC, USA, December 1, 1998. Shipping List #: 98-0733-M Shipping List Date: 06/08/98.

USNASA:1999:NSC

[Uni99] United States. National Aeronautics and Space Administration. Numerical Stability and Control Analysis Towards Falling-leaf Prediction Capabilities of Splitflow. NASA, Washington, DC, USA, January 20, 1999. Shipping List #: 99-0618-M Shipping List Date: 06/28/99.

VandeVelde:1994:CSC

[Van94] Eric F. Van de Velde. Concurrent scientific computing, volume 16 of Texts in Applied Mathematics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-94195-9. xix + 328 pp.

LCCN QA76.58 .V35 1994. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0815/93043289-
                                        d.html.

Vanderplaats:1996:NOM

[Van96] Garret N. Vanderplaats. Numerical optimization methods and applications: March 5–8, 1996, Engineering 856.9. University of California, Los Angeles, University Extension, Dept. of Engineering, Information Systems, and Technical Management, Short Course Program, Los Angeles, CA, USA, 1996. various pp. LCCN TAX 252. Cover title.

VanLoan:1997:ISC

[Van97] Charles F. Van Loan. Introduction to scientific computing: a matrix-vector approach using MATLAB. The MAT-LAB curriculum series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-125444-8. xi + 347 pp. LCCN QA76.9.M35V375 1997.

Varah:1990:IGF

[Var90a] James M. Varah. The influence of George Forsythe and his students. In Nash [Nas90b], pages 31–40. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Varah:1990:CPE

[Var90b] James M. Varah. On the conditioning of parameter estimation problems. In Cox and Hammarling [CH90], pages 187– 195. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00.

Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Varga:1990:RUM

[Var90c] Richard S. Varga. Reminiscences on the University of Michigan summer schools, the Gatlinburg Symposia, and Numerische Mathematik. In Nash [Nas90b], pages 206–210. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Varga:1990:SCM

[Var90d] Richard S. Varga. Scientific Computation on Mathematical Problems and Conjectures. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1990. ISBN 0-89871-257-2. vi + 122 pp.

Vassiliev:1996:TCR

[Vas96] Victor A. Vassiliev. Topological complexity of root-finding algorithms. In Renegar et al. [RSS96], pages 831–856. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Veldhuizen:1998:GFLb

[Vel98] Todd L. Veldhuizen. Grid Filters for Local Nonlinear Image Restoration. University of Waterloo, Waterloo, Ontario, Canada, 1998. URL http://

extreme.indiana.edu/~tveldhui/

papers/. M.A.Sc. Thesis.

Verheggen:1992:NMS

[Ver92] T. M. M. Verheggen, editor. Numerical methods for the simulation of multi-phase and complex flow, volume 398 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-540-55278-2 (Berlin), 0-387-55278-2 (New York). LCCN QA922 .N87 1990. Proceedings of a workshop held at Koninklijke/Shell-Laboratorium, Amsterdam, Amsterdam, the Netherlands, 30 May-1 June 1990.

vanderVorst:1997:YOS

[vG97] Henk A. van der Vorst and Gene H. Golub. 150 years old and still alive: Eigenproblems. In Duff and Watson [DW97], pages 93–120. ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

//global.oup.com/academic/
product/the-state-of-the-
art-in-numerical-analysis-
9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Vorozhtsov:1990:MLS

[VI90] E. V. (Evgenii Vasil’evich) Vorozhtsov and N. N. Ianenko. Methods for the localization of singularities in numerical solutions of gas dynamics problems. Springer series in computational physics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 0-387-50363-3. x + 406 pp. LCCN QA930 .V6613 1990. Translation of: Metody lokalizatsii osobennostei pri chislennom reshenii zadach gazodinamiki.

Vijayaraghavan:1995:ENP

[Vij95] D. Vijayaraghavan. An efficient numerical procedure for thermodydrodynamic [sic] analysis of cavitating bearings. National Technical Information Service, Washington, DC, USA, 1995. ???? pp. Shipping list no.: 96-0467-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1995] 1 microfiche.

Vinh:1991:INC

[Vin91] Hoang Vinh. An investigation of the numerical characteristics of finite-difference methods as applied to the time-domain maxwell’s equations. Thesis (m.s.) in mechanical engineering, University of California, Davis, Davis, CA, USA, 1991. 106 pp.

Vincent:1998:NMF

[Vin98] Alain Vincent. Numerical methods in fluid mechanics, volume 16 of CRM proceedings and lecture notes. American Mathematical Society, Providence, RI, USA, 1998. ISBN 0-8218-0813-3. ISSN 1065-8580. xiv + 199 pp. LCCN QC145.2 .N86 1998.

Vitasek:1994:ZTN

[Vit94] Emil Vitásek. Základy teorie numerických metod pro řešení diferenciálních rovnic. (Czech) [Basic theory of numerical methods for solving differential equations]. Academia [Publishing House of the Czech Academy of Sciences], Prague, Czech Republic, 1994. ISBN 80-200-0281-2. 411 pp.

Vandergrift:1995:NMA

[VJ95] Thomas L. Vandergrift and Charles V. Jude. Numerical modeling analysis of stress transfer modification concepts for deep longwall mines. U.S. Dept. of the Interior, Bureau of Mines, Pittsburgh, PA, USA, 1995. 43 pp. Shipping list no.: 96-0482-M.

Vreugdenhil:1993:NMA

[VK93] Cornelis Boudewijn Vreugdenhil and Barry Koren. Numerical methods for advection — diffusion problems, volume 45 of Notes on numerical fluid mechanics. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1993. ISBN 3-528-07645-3. ISSN 0179-9614. xii + 373 pp. LCCN QA901 .N86 1993. GB93-08615.

Vlchanov:1996:SMM

[VK96] Nikolaĭ Lyubomirov Vlchanov and Mikhail Mikhaĭlov Konstantinov. Svremenni matematicheski metodi za kompyutrni presmyataniya. Chast 1. (Bulgarian [Modern mathematical methods for computer calculations. Part 1]), volume 1 of Studii na BIAP. Matematicheski Nauki [Studies of the Bulgarian Institute of Analysis and Research. Mathematical Sciences]. Fondatsiya “Blgarski Institut za Analizi i Prouchvaniya”, Sofia, Bulgaria, 1996. ISBN 954-8949-01-6. 232 pp. Osnovi na kompyutrnite presmyataniya. Chisleno diferentsirane i integrirane. [Foundations of computer calculations. Numerical differentiation and integration].

vanLoon:1997:NMS

[vL97] M. van Loon. Numerical methods in smog prediction, volume 120 of CWI tract. CWI, Amsterdam, The Netherlands, 1997. ISBN 90-6196-473-3. iv + 149 pp. LCCN TD884.3 .L66 1997.

Volkov:1990:MN

[Vol90a]  E. A. Volkov.  Métodos

numéricos. (Spanish) [Numerical methods]. Editorial Mir, Moscow, Russia, 1990. ISBN 5-03-001551-5. 255 pp. Translated from the Russian by K. P. Medkov.

Volkov:1990:NM

[Vol90b] E. A. Volkov. Numerical methods. Mir Publishers and Hemisphere Publishing Corporation, Moscow, Russia and New York, NY, USA, English edition, 1990. ISBN 1-56032-011-7. 238 pp. LCCN QA297 .V5413 1990. Translated from the Russian by Leonid Levant.

Vreugdenhil:1994:NMS

[Vre94] Cornelis Boudewijn Vreugdenhil. Numerical methods for shallow-water flow, volume 13 of Water science and technology library. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1994. ISBN 0-7923-3164-8. ix + 261 pp. LCCN TC171 .V74 1994.

Vradis:1994:NMN

[VS94] George C. Vradis and Dennis A. Siginer, editors. Numerical methods for non-Newtonian fluid dynamics: presented at the 1994 ASME Fluids Engineering Division Summer Meeting, Lake Tahoe, Nevada, June 19– 23, 1994, volume 179 of FED. American Society of Mechanical Engineers, United Engineering Center, 345 E. 47th St., New York, NY 10017, USA, 1994. ISBN 0-7918-1362-2. LCCN QA929.5 .N85 1994.

Vouk:1997:QSS

[VS97] M. A. Vouk and M. P. Singh. Quality of service and scientific workflows. In Boisvert [Boi97], pages 77–89. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

vonSchwerin:1999:MSS

[vS99] Reinhold von Schwerin. Multibody system simulation: numerical methods, algorithms, and software, volume 7 of Lecture Notes in Computational Science and Engineering. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-65662-6 (softcover). ISSN 1439-7358. xx + 337 pp. LCCN QA845 .V66 1999.

Vulkov:1997:NAA

[VWY97] Lubin Vulkov, Jerzy Wasniewski, and Plamen Yalamov, editors. Numerical analysis and its applications: first international workshop, WNAA ’96, Rousse, Bulgaria, June 24–26, 1996: proceedings, number 1196 in Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-62598-4 (softcover). LCCN QA297 .W584 1996.

Vavasis:1996:RBL

[VY96] Stephen A. Vavasis and Yinyu Ye. On the relationship between layered least squares and affine scaling steps. In Renegar et al. [RSS96], pages 857–866. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Wakulicz:1990:NAM

[Wak90]  Andrzej Wakulicz.  Numer-

ical analysis and mathematical modelling, volume 24 of Banach Center publications. PWN-Polish Scientific Publishers, Warszawa, Poland, 1990. ISBN 83-01-09355-2. ISSN 0137-6934. 566 pp. LCCN QA297 .N842 1990. English and Russian.

Walker:1996:FFT

[Wal96a]  James S. Walker.  Fast

Fourier Transforms. Studies in Advanced Mathematics. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 1996. ISBN 0-8493-7163-5. xv + 447 pp. LCCN QA403 .W33 1996. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0646/96016852-

d.html. With 1 IBM-PC floppy disk (3.5 inch; HD).

Walter:1996:EMC

[Wal96b] Wolfgang Walter. Enclosure methods for capricious solutions of ordinary differential equations. In Renegar et al. [RSS96], pages 867–878. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Wang:1991:NSG

[Wan91] Hongtao Wang. Numerical software with guaranteed results. Thesis (m.s.) in computer science, University of California, Davis, Davis, CA, USA, 1991. 203 pp.

Watkins:1991:FMC

[Wat91] David S. Watkins. Fundamentals of Matrix Computations. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 1991. ISBN 0-471-61414-9. xiii + 449 pp. LCCN QA188 .W38 1991. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0608/90048895-

d.html;   http://www.loc.
gov/catdir/enhancements/fy0609/
90048895-b.html;   http://
www.loc.gov/catdir/toc/onix02/
90048895.html.

Watson:1992:NAO

[Wat92] K. K. Watson. Numerical analysis of one dimensional soil water systems. Number 179 in Research report. University of New South Wales, Water Research Laboratory, Australia, April 1992. ISBN 0-85824-403-9. 70 pp. LCCN G8942 X-1 no.179.

Watkins:1996:LAO

[Wat96] David S. Watkins. QR-like algorithms — An overview of convergence theory and practice. In Renegar et al. [RSS96], pages 879–894. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Watson:1997:AAE

[Wat97] G. A. Watson. Aspects of approximation with emphasis on the univariate case. In Duff and Watson [DW97], pages 261–

282. ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

//global.oup.com/academic/
product/the-state-of-the-
art-in-numerical-analysis-
9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Werschulz:1996:CPP

[Wer96] Arthur G. Werschulz. The complexity of the Poisson problem for spaces of bounded mixed derivatives. In Renegar et al. [RSS96], pages 895–914. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Wesseling:1990:PEG

[Wes90] Pieter Wesseling, editor. Proceedings of the Eighth GAMM-Conference on Numerical Methods in Fluid Mechanics, volume 29 of Notes on numerical fluid mechanics. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1990. ISBN 3-528-07629-1. ISSN 0179-9614. LCCN TA 357 G35 1989.

Watson:1996:NA

[WG96] G. A. Watson and D. F. Griffiths, editors. Numerical analysis 1995, volume 344 of Pitman research notes in mathematics series. Longman Scientific and Technical, Harlow, Essex, UK, 1996. ISBN 0-582-27633-0 (paperback). ISSN 0269-3674. LCCN QA297.N854 1996. Proceedings of the 16th Dundee Biennial Conference on Numerical Analysis, University of Dundee, June 27–30, 1995.

Wheeler:1990:PCU

[Whe90] David J. Wheeler. Programmed computing at the Universities of Cambridge and Illinois in the early fifties. In Nash [Nas90b], pages 269–279. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Whittington:1998:NMP

[Whi98] Stuart G. Whittington, editor. Numerical methods for polymeric systems, volume 102 of The IMA volumes in mathematics and its applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98557-3. LCCN QD381.9.M3 W48 1998 Bar. Based on the proceedings of a workshop held in May 1996 at the IMA.

Wieners:1996:NES

[Wie96] Christian Wieners. Numerical enclosures for solutions of the Navier–Stokes equation for small Reynolds numbers. In Alefeld and Herzberger [AH96], pages 280–286. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Watanabe:1994:MSP

[WNO94] T. (Tsutomu) Watanabe, Makoto Natori, and Tsutomu Oguni. Mathematical Software for the P.C. and Work Stations: a Collection of Fortran 77 Programs. North-Holland Publishing Co., Amsterdam, The Netherlands, June 1994. ISBN 0-444-82000-0. xiv + 387 pp. LCCN QA 76.73 F25 F6813 1994. US$178.50. URL http:/

/www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=

0444820000. Translation of:

FORTRAN 77 ni yoru suchi keisan sofutowea.

Webster:1990:SMS

[WO90] R. Webster and M. A. Oliver. Statistical methods in soil and land resource survey. Number UK£40.00 (US$80.00 U.S.) (hardcover) in Spatial information systems. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-823317-5 (hardcover), 0-19-823316-7 (paperback). 316 pp. LCCN S592.14 .W4 1990. UK£20.00 (US$40.00). Rev. ed. of: Quantitative and numerical methods in soil classification and survey. 1977.

Woodford:1992:SLN

[Woo92] Chris Woodford. Solving linear and nonlinear equations. Ellis Horwood Series: Mathematics and its Applications. Ellis Horwood, New York, NY, USA, 1992. ISBN 0-13-830415-7. 190 pp.

Wood:1993:INM

[Woo93] W. L. Wood. Introduction to numerical methods for water resources. Oxford science publications. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1993. ISBN 0-19-859690-1. xxii + 255 pp. LCCN TC163 .W66 1993. UK£35 (US$59.95).

Wozniakowski:1996:OIB

[Woz96] Henryk Wozniakowski. Overview of information-based complexity. In Renegar et al. [RSS96], pages 915–927. ISBN 0-8218-0530-4. ISSN 0075-8485. LCCN QA297 .A57 1995.

Weilmuenster:1991:NAS

[WSG91] K. James Weilmuenster, Robert E.

Smith, and Francis A. Greene. Numerical analysis and simulation of an assured crew return vehicle flow field. National Technical Information Service, Washington, DC, USA, 1991. ???? pp. Shipping list no.: 92-1251-M. Microfiche. [Washington, DC?: National Aeronautics and Space Administration], 1992. 1 microfiche.

Wang:1990:SBY

[WY90] De Ren Wang and Zhong Hua Yang. Shuzhi bijin yinlun. (Mandarin Chinese) []. Gaodeng Xuexiao Shiyong Jiaocai. [Trial Educational Materials for Advanced Schools]. Gaodeng Jiaoyu Chubanshe (Higher Education Press), Beijing, People’s Republic of China, 1990. ISBN 7-04-002244-3. vi + 387 pp.

Wu:1994:SAO

[WZ94] S. T. (Shi-Tsan) Wu and Minggang Zhou. Study for the analysis of the observations, and numerical data representing the planets as far-infrared calibration sources: final report. National Technical Information Service, Washington, DC, USA, 1994. ???? pp. Title from microfiche header. Shipping list no.: 95-0419-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1995] 1 microfiche.

Xu:1999:GED

[Xu99] Kun Xu. Gas evolution dynamics in Godunov-type schemes and analysis of numerical shock instability. National Technical Information Service, Washington, DC, USA, 1999. ???? pp. Shipping list no.: 99-0729-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1999] 1 microfiche.

Yamamoto:1996:SLM

[Yam96] T. Yamamoto. SOR-like methods for the simultaneous determination of polynomial zeros. In Alefeld and Herzberger [AH96], pages 287–302. ISBN 3-05-501696-3 (paperback). ISSN 0138-3019. LCCN QA297 .I455 1995.

Yang:1996:NSLa

[YCC96a] S. L. Yang, R. Chen, and M. C. Cline. Numerical simulation of a low emissions gas turbine combustor using KIVA-II. Part I, Fuel nozzle/rich-burn zone analysis. National Technical Information Service, Washington, DC, USA, 1996. ???? pp. Shipping list no.: 99-0069-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1997] 1 microfiche.

Yang:1996:NSLb

[YCC96b] S. L. Yang, R. Chen, and M. C. Cline. Numerical simulation of a low emissions gas turbine combustor using KIVA-II. Part II, Quick-quench/lean-combustion zones analysis. National Technical Information Service, Washington, DC, USA, 1996. ???? pp. Shipping list no.: 99-0069-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1997] 1 microfiche.

Kuo:1992:PSC

[yKaY92] Pen yu Kuo and Lung an Ying, editors. Proceedings of the Second Conference on Numerical Methods for Partial Differential Equations: Nankai University, Tianjin, China, 1–4 May, 1991. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1992. ISBN 981-02-0929-0. LCCN QA374 .C6535 1991.

Young:1990:HRI

[You90] David M. Young. A historical review of iterative methods. In Nash [Nas90b], pages 180–

194. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Zorkaletisev:1995:MNK

[ZAB95] V. I. (Valeriĭ Ivanovich) Zorkalëtísev, E. G. (Evgeniĭ Georgievich) Anëtísiferov, and V. P. (Valerëiían Pavlovich) Bulatov. Metod naimenshikh kvadratov: geometricheskie svoistva, alternativnye podkhody, prilozheniëiĭa. (Russian) [Least-squares method: geometric properties, alternative approaches, applications]. VO “Nauka”, Novosibirsk, Russia, 1995. ISBN 5-02-030676-2. 218 pp. LCCN QA275 .Z67 1995.

Zachary:1996:ISP

[Zac96] Joseph Zachary. Introduction to scientific programming: computational problem solving using Maple and C. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94630-6. xxiv + 380 pp. LCCN QA76.6 .Z32 1996.

Zachary:1998:ISP

[Zac98] Joseph L. Zachary. Introduction to scientific programming: computational problem solving using Mathematica and C. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1998. ISBN 0-387-98250-7 (hardcover). xxiv + 433 pp. LCCN QA76.6.Z324 1998.

Zeitzew:1995:NMN

[Zei95] Michael Alan Zeitzew. Numerical methods for nonlinear ordinary differential equations with different time-scales. Thesis (ph.d.), University of California, Los Angeles, Los Angeles, CA, USA, 1995. xi + 154 pp.

Zhou:1994:NEA

[Zho94] Jun Zhou. Numerical and experimental analysis of diesel and JP-5 transport in unsaturated soils. Thesis (ph.d.), University of California, Los Angeles, Los Angeles, CA, USA, 1994. xiii + 221 pp.

Zhou:1996:NAS

[Zho96] Hong Zhou. Numerical analysis of slender vortex motion. Thesis (ph.d.) in applied mathematics, University of California, Berkeley, Berkeley, CA, USA, May 1996.

Zhukov:1992:MFV

[Zhu92] A. I. Zhukov. Metod Fure v vychislitelnoi matematike. Nauka, Moscow, Russia, 1992. ISBN 5-02-014885-7. 176 pp.

Zorkaltsev:1995:MNK

[Zor95]  V. I. Zorkal′tsev.  Metod

naimenshikh kvadratov. (Russian) [Least-squares method]. VO “Nauka”, Novosibirsk, 1995. ISBN 5-02-030676-2. 220 pp. Geometricheskie svoistva, alternativnye podkhody, prilozheniya [Geometric properties, alternative approaches, applications].

Zwillinger:1992:HI

[Zwi92] Daniel Zwillinger. Handbook of Integration. A. K. Peters, Ltd., Wellesley, MA, USA, 1992. ISBN 0-86720-293-9, 1-4398-6584-1 (e-book). xv + 367 pp. LCCN QA299.3 .Z85 1992.